### v1

# Efficient Conditional Gradient Methods for Solving Stochastic Convex Bilevel Optimization Problems

Abstract
We propose efficient methods for solving stochastic convex bilevel optimization problems, where the goal is to minimize an outer stochastic objective function subject to the solution set of an inner stochastic optimization problem. Existing methods often rely on costly projection or linear optimization oracles over complex sets, which limits scalability. To overcome this, we propose an iteratively regularized conditional gradient framework that leverages efficient linear optimization oracles exclusively over the base feasible set. Our proposed methods employ a vanishing regularization sequence that progressively emphasizes the inner problem while biasing towards desirable minimal outer objective solutions. Under standard convexity assumptions, we establish non-asymptotic convergence rates of $O(t^{-({1}/{2}-p)})$ for the outer objective and $O(t^{-p})$ for the inner objective, where $p\in(0,1/2)$ controls the regularization decay, in the one-sample stochastic setting, and $O(t^{-(1-p)})$ and $O(t^{-p})$ in the finite-sum setting using a mini-batch scheme, where $p\in(0,1)$. Experimental results on over-parametrized regression and $\ell_{1}$-constrained logistics regression tasks demonstrate the practical advantages of our approach over existing methods, confirming our theoretical findings.
1 Introduction
We consider stochastic convex bilevel optimization problems, where the goal is to minimize an outer stochastic convex objective function subject to the solution set of an inner stochastic convex optimization problem. Formally, the problem is defined as
$F_{\operatorname{opt}}:=\min_{x\in{\mathbb{R}}^{d}}\left\{F(x):={\mathbb{E}}[f(x,\theta)]:x\in X_{\operatorname{opt}}\right\},\ \text{where }X_{\operatorname{opt}}:=\operatorname*{arg\,min}_{z\in X}G(z):={\mathbb{E}}[g(z,\xi)].$ (1)
Here $X\subset\mathbb{R}^{d}$ is a convex and compact set, and both $F$ and $G$ are smooth and convex, making the optimization problem (1) convex. This framework has broad applicability, including hyper-parameter optimization [15, 16, 38], meta-learning [16, 42], reinforcement learning [31] and game theory [50]. In particular, problem (1) arises in learning applications where $G$ is the prediction error of a model $x$ on data, and $F$ is an auxiliary objective (e.g., regularization, or a validation data set). The dependence on data is through the expectation operators and the variables $\theta,\xi$. The bilevel formulation provides a tuning-free alternative to the usual regularization approach where we optimize $\sigma F(x)+G(x)$ for a carefully tuned parameter $\sigma$.
Despite its wide applications in machine learning there are three primary challenges in solving (1). First, we do not have an explicit representation of the optimal set $X_{\operatorname{opt}}$ in general, which prevents us from using some common operations in optimization such as projection onto or linear optimization over $X_{\operatorname{opt}}$. An alternative is to reformulate (1) using its value function formulation
$\min_{x\in X}\quad F(x)\quad\text{s.t.}\quad G(x)\leq G_{\operatorname{opt}}:=\min_{z\in X}G(z).$ (2)
However, this leads to our second challenge. Problem (2) is inherently ill-conditioned: by definition of $G_{\operatorname{opt}}$, there exists no $x\in X$ such that $G(x)<G_{\operatorname{opt}}$, hence (2) does not satisfy Slater’s condition. As a result, the Lagrangian dual of (2) may not be solvable, complicating the use of standard primal-dual methods. In addition, we do not know $G_{\operatorname{opt}}$ a priori, so an alternative is to approximate it with $\bar{G}\leq G_{\operatorname{opt}}+\epsilon_{g}$, and consider the constraint $G(x)\leq\bar{G}$ instead. However, this approach does not solve the actual bilevel problem (1), and may still introduce numerical instability. The third challenge comes from the stochastic nature of the objectives. Since $F$ and $G$ are defined as expectations, their exact computation may be intractable when dealing with large-scale datasets. In single-level optimization, this is typically addressed through sampling-based methods that operate on mini-batches drawn from the distributions of $\theta$ and $\xi$ instead of the actual distributions themselves. However, such stochastic approximations introduce noise into the optimization process, necessitating new techniques to control the noise and ensure convergence.
1.1 Related works
Iterative regularization methods.
Simple bilevel optimization extends Tikhonov regularization [52] beyond $\ell_{2}$ penalties by considering a blended objective $\sigma F(x)+G(x)$, where $\sigma>0$ modulates the trade-off between inner and outer objectives. As $\sigma\to 0$, the problem converges to a bilevel form that prioritizes the inner objective. Friedlander and Tseng [17] established fundamental conditions ensuring solution existence for $\sigma>0$, forming the basis for iterative regularization, where $\sigma$ is gradually reduced. This idea was first implemented by Cabot [3] in the unconstrained case $X={\mathbb{R}}^{d}$ using a decreasing sequence $\sigma_{t}$, and later extended to general convex constraints by Dutta and Pandit [13]. While these methods use proximal steps, which can be computationally expensive, Solodov [49] proposed a more efficient gradient-based alternative. Helou and Simões [23] further generalized the framework to non-smooth settings via an $\epsilon$-subgradient method. These works show asymptotic convergence but lack convergence rates. Amini and Yousefian [1] analyzed an iterative regularized projected gradient method and established a convergence rate of $O(t^{-(1/2-b)})$ for the inner objective, assuming non-smooth $F$ and $G$, compact $X$, and strong convexity of $f$. Kaushik and Yousefian [29] later removed the strong convexity assumption and proved rates of $O(t^{-b})$ and $O(t^{-(1/2-b)})$ for the inner and outer objectives, respectively, for any $b\in(0,1/2)$. Malitsky [34] studied an accelerated variant based on Tseng’s method [53], achieving an $o(t^{-1})$ rate for the inner problem. More recently, Merchav et al. [36] proposed an accelerated proximal scheme with a $O(\log(t+1)/t)$ rate under smoothness assumption and $O(t^{-2})$ / $O(1/t^{2-\gamma})$ under a Hölderian error bound condition. Our work also follows an iterative regularization approach, but avoids expensive projections or proximal operations by leveraging linear optimization oracles.
Projection-free methods.
Several important applications have domains $X$ where projection-based operations are inefficient, yet linear optimization over $X$ is easier. To address the limitations of projection-based methods, recent work has focused on projection-free bilevel optimization via linear optimization oracles. Jiang et al. [28] approximated $X_{\operatorname{opt}}$ by replacing the constraint $G(x)\leq G_{\operatorname{opt}}$ with a linear approximation, which removed the need for Slater’s condition, and applying the conditional gradient method, refining the approximation at each step. They established $O(t^{-1})$ convergence rates for both objectives under smoothness and compactness assumptions. However, their method requires a pre-specified tolerance parameter $\epsilon_{g}$, and only guarantees $(\epsilon_{g}/2)$-infeasibility. Doron and Shtern [11] proposed an alternative approach based on sublevel sets of the outer function $F$. Their method performs conditional gradient updates over sets of the form $X\times\{x:F(x)\leq\alpha\}$, with a surrogate $\hat{G}_{t}$ that is updated iteratively. They achieved rates of $O(t^{-1})$ for the inner and $O(t^{-1/2})$ for the outer objective under composite structure in $G$ and an error bound on $F$. While linear optimization over $X$ may be efficient, this is not always true for sets like $X\cap H$, where $H$ is a halfspace, or $\{F(x)\leq\alpha\}$ unless $F$ has a special structure. To address this, Giang-Tran et al. [19] introduced a conditional gradient-based iterative regularization scheme that only requires linear optimization over $X$, albeit with slower rates: $O(t^{-p})$ and $O(t^{-(1-p)})$ for the inner and outer problems, respectively, for any $p\in(0,1)$. We extend this framework to the stochastic setting, maintaining the same oracle-based reliance on linear optimization over $X$.
Stochastic methods.
Stochastic bilevel optimization remains less explored. Jalilzadeh et al. [26] developed an iterative regularization-based stochastic extragradient algorithm, requiring $O(1/\epsilon_{f}^{4})$ and $O(1/\epsilon_{g}^{4})$ stochastic gradient queries to achieve $\epsilon$-optimality for both objectives. Cao et al. [4] extended the projection-free framework of Jiang et al. [28] to the stochastic setting. Their methods achieve $\tilde{O}(t^{-1})$ convergence rate when the noise distributions have finite support, and $O(t^{-1/2})$ convergence rates under sub-Gaussian noise. However, their method requires linear optimization over intersections of $X$ with halfspaces. In contrast, our work relaxes this requirement by relying solely on linear optimization over the base set $X$.
Other methods and bilevel problem classes.
Several alternative approaches exist for simple bilevel optimization, including sequential averaging schemes [44, 46, 35], sublevel set-based methods [2, 5, 58, 20], accelerated algorithms [45, 54, 6, 5] and primal-dual strategies [47]. These methods are less closely related to our framework and do not address the stochastic setting. More general bilevel problems have seen considerable attention, particularly in Stackelberg game formulations where the upper and lower levels model leader-follower dynamics [18, 7, 30, 27, 9, 32, 56, 33]. Recent extensions include contextual bilevel optimization with exogenous variables [55, 24, 51] and problems with functional constraints at the lower level [29, 40]. In contrast, our focus is restricted to the stochastic convex formulation of simple bilevel problems.
1.2 Contributions
The key contributions of this paper are summarized below.
  • $\diamond$
    One-sample stochastic setting: We develop a projection-free method for stochastic bilevel optimization that achieves high-probability convergence rates of $O(t^{-(1/2-p)}\sqrt{\log(dt^{2}/\delta)})$ for the outer-level problem and $O(t^{-p}\sqrt{\log(d/\delta)})$ for the inner-level problem, with probability at least $1-\delta$ for any $p\in(0,1/2)$. Compared to the state-of-the-art approach by Cao et al. [4, Algorithm 1], which achieves the $O(t^{-1/2}\sqrt{\log(Td/\delta)})$ rate for both the outer and inner problems, our method offers several key advantages. First, it eliminates the need for optimization over intersections of the base feasible set with halfspaces, requiring only linear optimization over the original constraint set. Second, it removes the dependency on their pre-specified tolerance parameter $\epsilon_{g}$, and the associated computationally expensive initialization procedure to find an $\epsilon_{g}$-optimal starting point. Thus, our algorithm enjoys anytime guarantees.
  • $\diamond$
    Finite-sum setting: For problems with $n$ component functions, we establish convergence rates of $O(t^{-(1-p)}\sqrt{\log(t^{2}/\delta)})$ (outer) and $O(t^{-p}\sqrt{\log(1/\delta)})$ (inner) with probability $1-\delta$, for any $p\in(0,1)$. Although Cao et al. [4, Algorithm 2] achieves improved $O(t^{-1}\log(T)\sqrt{\log(T/\delta)})$ rates for both objectives, it again relies on more computationally demanding halfspace-intersection oracles, requires an additional initialization step, and depends on a fixed stepsize.
  • $\diamond$
    Numerical Validation: We demonstrate the practical efficiency of our methods on over-parametrized regression and $\ell_{1}$-constrained logistic regression tasks, showing significant speedups over existing projection-based and projection-free approaches. The experiments validate both the convergence rates and computational advantages of our framework.
Proof of these results are provided in Appendix A. Additional results are presented in Appendix B.
2 Algorithms and convergence guarantees
In this section, we present two iterative regularization approaches for solving problem (1) in both stochastic and finite-sum settings. The core idea is to introduce a decreasing regularization sequence $\{\sigma_{t}\}_{t\geq 1}$, and at each iteration $t$, perform a gradient-based update on the composite objective $\sigma_{t}F+G$. As $\sigma_{t}$ decreases, the algorithm gradually shifts focus from $F$ to $G$, thereby steering $x_{t}$ toward points in $X_{\operatorname{opt}}$. At the same time the $\sigma_{t}F$ term encourages $x_{t}$ towards a point in $X_{\operatorname{opt}}$ that minimizes $F$. Giang-Tran et al. [19] applied this idea to design a projection-free algorithm for the deterministic setting, establishing convergence rates of $O(t^{-p})$ for the upper level and $O(t^{-(1-p)})$ for the lower level when using $\sigma_{t}=\varsigma(t+1)^{-p}$ with $p\in(0,1)$.
In the stochastic setting Eq. 1, however, computing exact gradients of the upper and lower objective functions is computationally expensive. To address this challenge, we rely on sample-based gradient estimators. Specifically, given mini-batch sample sets ${\cal S}_{f}:=\{\theta_{1},\ldots,\theta_{|{\cal S}_{f}|}\}$ and ${\cal S}_{g}:=\{\xi_{1},\ldots,\xi_{|{\cal S}_{g}|}\}$ drawn independently at each iteration, we define the empirical gradients as
$\nabla f_{{\cal S}_{f}}(x):=\frac{1}{|{\cal S}_{f}|}\sum_{\theta\in{\cal S}_{f}}\nabla f(x,\theta)\quad\&\quad\nabla g_{{\cal S}_{g}}(x):=\frac{1}{|{\cal S}_{g}|}\sum_{\xi\in{\cal S}_{g}}\nabla g(x,\xi),$
which provide unbiased estimates of $\nabla F(x)$ and $\nabla G(x)$, respectively. A naïve approach is to use these sample gradients directly, that is, $\widehat{\nabla F}(x)=\nabla f_{\mathcal{S}_{f}}(x)$ and $\widehat{\nabla G}(x)=\nabla g_{\mathcal{S}_{g}}(x)$. However, these estimators maintain constant variance over optimization time, which prevents achieving optimal convergence rates. To address this limitation, we implement two distinct variance-reduced gradient estimation schemes tailored to different problem settings.
  • $\diamond$
    One-sample stochastic setting: When we access single samples of $\theta$ and $\xi$ per iteration, we employ the STOchastic Recursive Momentum (STORM) estimator [8], which progressively reduces variance through momentum-based updates under sub-Gaussian noise assumptions. The resulting method, called the Iteratively Regularized Stochastic Conditional Gradient (IR-SCG) algorithm, is summarized in Algorithm 1. We note that the use of the STORM estimator in projection-free methods was first introduced in [60].
  • $\diamond$
    Finite-sum setting: When the expectations in (1) are finite sums of size $n$, we use the the Stochastic Path-Integrated Differential EstimatoR (SPIDER[37, 14] with $\sqrt{n}$-sized mini-batches, achieving variance reduction through periodic gradient recomputation. The resulting method, called the Iteratively Regularized Finite-Sum Conditional Gradient (IR-FSCG) algorithm, is summarized in Algorithm 2. Here, for any $n\in{\mathbb{N}}$, $[n]:=\{1,\dots,n\}$. We note that the SPIDER estimator in projection-free methods has been previously used in [48, 57, 59, 22].
While a mini-batch scheme can be applied even when expectations are not finite sums, we emphasize that this approach only improves the constants in the convergence rates and enhances algorithmic stability without affecting the asymptotic rate. Additionally, if the upper- and lower-level problems have different numbers of component functions ($n_{f}\neq n_{g}$), Algorithm 2 can be easily modified by introducing $S_{f}$ and $q_{f}$ for the upper level, and $S_{g}$ and $q_{g}$ for the lower level.
Data: Parameters $\{\alpha_{t}\}_{t\geq 0}\subseteq[0,1]$,$\{\sigma_{t}\}_{t\geq 0}\subseteq\mathbb{R}_{++}$.
Result: sequences $\{x_{t},z_{t}\}_{t\geq 1}$.
Initialize $x_{0}\in X$;
for $t=0,1,2,\ldots$ do
       if $t=0$ then
            Compute
$\widehat{\nabla F}_{t}:=\nabla_{x}f(x_{t},\theta_{t}),\quad\widehat{\nabla G}_{t}:=\nabla_{x}g(x_{t},\xi_{t})$
      else
            Compute
$\widehat{\nabla F}_{t}:=(1-\alpha_{t})\widehat{\nabla F}_{t-1}+\nabla_{x}f(x_{t},\theta_{t})-(1-\alpha_{t})\nabla_{x}f(x_{t-1},\theta_{t})$
$\widehat{\nabla G}_{t}:=(1-\alpha_{t})\widehat{\nabla G}_{t-1}+\nabla_{x}g(x_{t},\xi_{t})-(1-\alpha_{t})\nabla_{x}g(x_{t-1},\xi_{t})$
      Compute
$v_{t}$ $\in\operatorname*{arg\,min}_{v\in X}\left\{(\sigma_{t}\widehat{\nabla F}_{t}+\widehat{\nabla G}_{t})^{\top}v\right\}$
$x_{t+1}$ $:=x_{t}+\alpha_{t}(v_{t}-x_{t})$
$S_{t+1}$ $:=(t+2)(t+1)\sigma_{t+1}+\sum_{i\in[t+1]}(i+1)i(\sigma_{i-1}-\sigma_{i})$
$z_{t+1}$ $:=\frac{(t+2)(t+1)\sigma_{t+1}x_{t+1}+\sum_{i\in[t+1]}(i+1)i(\sigma_{i-1}-\sigma_{i})x_{i}}{S_{t+1}}$
Algorithm 1 Iteratively Regularized Stochastic Conditional Gradient (IR-SCG) Algorithm.
Data: Parameters $\{\alpha_{t}\}_{t\geq 0}\subseteq[0,1]$, $\{\sigma_{t}\}_{t\geq 0}\subseteq\mathbb{R}_{++}$, $S,q\in[n]$.
Result: sequences $\{x_{t},z_{t}\}_{t\geq 1}$.
Initialize $x_{0}\in X$;
for $t=0,1,2,\ldots$ do
       if $t=0\mod q$ then
            Compute
$\widehat{\nabla F}_{t}:=\nabla F(x_{t}),\quad\widehat{\nabla G}_{t}:=\nabla G(x_{t}).$
      else
            Draw $S$ new i.i.d. samples ${\cal S}_{f}=\{\theta_{1},\ldots,\theta_{S}\}$, ${\cal S}_{g}=\{\xi_{1},\ldots,\xi_{S}\}$;
             Compute
$\widehat{\nabla F}_{t}:=\widehat{\nabla F}_{t-1}+\nabla f_{\mathcal{S}_{f}}(x_{t})-\nabla f_{\mathcal{S}_{f}}(x_{t-1})$
$\widehat{\nabla G}_{t}:=\widehat{\nabla G}_{t-1}+\nabla g_{\mathcal{S}_{g}}(x_{t})-\nabla g_{\mathcal{S}_{g}}(x_{t-1}).$
      Compute
$v_{t}$ $\in\operatorname*{arg\,min}_{v\in X}\left\{(\sigma_{t}\widehat{\nabla F}_{t}+\widehat{\nabla G}_{t})^{\top}v\right\}$
$x_{t+1}$ $:=x_{t}+\alpha_{t}(v_{t}-x_{t})$
if $t\geq q$ then
            Compute
$S_{t+1}$ $:=(t+2)(t+1)\sigma_{t+1}+\sum_{i\in[t+1]\setminus[q]}(i+1)i(\sigma_{i-1}-\sigma_{i})$
$z_{t+1}$ $:=\frac{(t+2)(t+1)\sigma_{t+1}x_{t+1}+\sum_{i\in[t+1]\setminus[q]}(i+1)i(\sigma_{i-1}-\sigma_{i})x_{i}}{S_{t+1}}.$
Algorithm 2 Iteratively Regularized Finite-Sum Conditional Gradient (IR-FSCG) Algorithm.
We emphasize that, unlike [4], both Algorithms 1 and 2 are anytime methods that do not require careful initialization step, preset stepsizes, or a priori choice of the number of total iterations $T$ based on target accuracies. This anytime property enables us to establish asymptotic convergence guarantees to the optimal solution, as formally stated in Theorems 3 and 4 in Appendix B. To establish convergence rates for these algorithms, we impose the following standard assumptions on the bilevel problem (1). In the following, all norms refer to the Euclidean norm.
Assumption 1.
The following hold.
  1. (a)
    $X\subseteq\mathbb{R}^{d}$ is convex and compact with diameter $D<\infty$, i.e., $\|x-y\|\leq D$ for any $x,y\in X$.
  2. (b)
    Functions $F,G$ are convex over $X$ and continuously differentiable on an open neighborhood of $X$.
  3. (c)
    For any $\theta$, $f(\cdot,\theta)$ is $L_{f}$-smooth on an open neighborhood of $X$, i.e., it is continuously differentiable and its derivative is $L_{f}$-Lipschitz:
    $\|\nabla_{x}f(x,\theta)-\nabla_{x}f(y,\theta)\|\leq L_{f}\|x-y\|,$
    for any $x,y\in X$.
  4. (d)
    For any $\xi$, $g(\cdot,\xi)$ is $L_{g}$-smooth on an open neighborhood of $X$.
  5. (e)
    For any $x\in X$, the stochastic gradients noise is sub-Gaussian, i.e., there exists some $\sigma_{f},\sigma_{g}>0$ such that
    ${\mathbb{E}}\left[\exp\left(\|\nabla_{x}f(x,\theta)-\nabla F(x)\|^{2}/\sigma_{f}^{2}\right)\right]$ $\leq\exp(1),$
    ${\mathbb{E}}\left[\exp\left(\|\nabla_{x}g(x,\xi)-\nabla G(x)\|^{2}/\sigma_{g}^{2}\right)\right]$ $\leq\exp(1).$
Our convergence rate guarantees are detailed in Theorems 2 and 1.
Theorem 1 (One-sample stochastic setting).
Let $\{z_{t}\}_{t\geq 1}$ be the iterates generated by Algorithm 1 with $\alpha_{t}=2/(t+2)$ for any $t\geq 0$, and regularization parameters $\{\sigma_{t}\}_{t\geq 0}$ given by $\sigma_{t}:=\varsigma(t+1)^{-p}$ for some chosen $\varsigma>0$, $p\in(0,1/2)$. Under 1, for any $t\geq 1$, with probability at least $1-\delta$, we have
$F(z_{t})-F_{\operatorname{opt}}\leq O\left(t^{-(1/2-p)}\sqrt{\log\left(\frac{dt^{2}}{\delta}\right)}\right)\quad\&\quad G(z_{t})-G_{\operatorname{opt}}$ $\leq O\left(t^{-p}\sqrt{\log\left(\frac{d}{\delta}\right)}\right).$
Theorem 2 (Finite sum setting).
Suppose the expectations defining $F$ and $G$ in Eq. 1 are from uniform distributions over finite sets of size $[n]$. Let $\{z_{t}\}_{t\geq 1}$ be the iterates generated by Algorithm 2 with $\alpha_{t}=\log(q)/q$ for $0\leq t<q$, $\alpha_{t}=2/(t+2)$ for any $t\geq q$, and $\sigma_{t}:=\varsigma(\max\{t,q\}+1)^{-p}$ for some chosen $\varsigma>0$, $p\in(0,1)$ and $S=q$. Under 1, with probability at least $1-\delta$, for any $t\geq q$, it holds that
$F(z_{t})-F_{\operatorname{opt}}$ $\leq O\left(\max\left\{1,\frac{q\log(q)}{t}\right\}t^{-(1-p)}\sqrt{\log\left(\frac{t^{2}}{\delta}\right)}\right),$
$G(z_{t})-G_{\operatorname{opt}}$ $\leq O\left(t^{-p}\sqrt{\log\left(\frac{1}{\delta}\right)}\right).$
We present a proof sketch here; complete proofs are provided in Appendix A. Both results share a common structure, with key differences highlighted below.
Proof sketch of Theorems 1 and 2.
We employ the notation $\Phi_{t}(x):=\sigma_{t}F(x)+G(x)$. Let $\widehat{\nabla\Phi}_{t}=\sigma_{t}\widehat{\nabla F}_{t}+\widehat{\nabla G}_{t}$ be the gradient estimate at step $t$. Using standard analysis for conditional gradient methods with the stepsize $\alpha_{t}=2/(t+2)$, Lemmas 2 and 1 in Appendix A establish the inequality
$t(t+1)\left(\Phi_{t}(x_{t})-\Phi_{t}(x_{\operatorname{opt}})\right)$ $+\textstyle\sum_{i\in[t]\setminus[k]}(i+1)i(\sigma_{i-1}-\sigma_{i})(F(x_{i})-F_{\operatorname{opt}})$
$\textstyle\leq(k+1)k\left(\Phi_{k}(x_{k})-\Phi_{k}(x_{\operatorname{opt}})\right)+2(\sigma_{0}L_{f}+L_{g})D^{2}(t-k)$ (3)
$\quad\textstyle+\sum_{i\in[t]\setminus[k]}4iD\|\widehat{\nabla\Phi}_{i-1}-\nabla\Phi_{i-1}(x_{i-1})\|$
for every $t\geq k$. Proposition 1 in Appendix A uses (3) to provide bounds for both the upper- and lower-level problems. Note that in the proof of Algorithm 1, we set $k=0$, while in in the proof of Algorithm 2, we set $k=q$. Nevertheless, the underlying idea remains essentially the same.
The key is to bound the last term $\sum_{i\in[t]}4iD\|\widehat{\nabla\Phi}_{i-1}-\nabla\Phi_{i-1}(x_{i-1})\|$, which is the error term from using noisy gradient estimates. To establish this bound with high probability, we leverage the concentration properties of the STORM estimator employed in Algorithm 1. Thanks to [4, Lemma 4.1], we obtain the following key probabilistic guarantee. For any iteration $i\in[t]$ and confidence level $\eta_{i}\in(0,1)$, with probability at least $1-\eta_{i}$, we (jointly) have
$\left\|\widehat{\nabla F}_{i}-\nabla F(x_{i})\right\|$ $\leq O\left(\frac{\sqrt{\log(4d/\eta_{i})}}{(i+1)^{1/2}}\right)\quad\&\quad\left\|\widehat{\nabla G}_{i}-\nabla G(x_{i})\right\|\leq O\left(\frac{\sqrt{\log(4d/\eta_{i})}}{(i+1)^{1/2}}\right),$
where the notation $O(\cdot)$ hides problem-specific constants $L_{f},L_{g},\sigma_{f},\sigma_{g},D$. Proposition 2 combines these inequalities for all $i\in[t]$ with $\eta_{i}=\delta/(i(i+1))$ and applies a union bound to conclude that with probability at least $1-\delta$, for any $t\geq 1$, we have
$\sum_{i\in[t]}i\left\|\widehat{\nabla\Phi}_{i-1}-\nabla\Phi_{i-1}(x_{i-1})\right\|\leq O\left((t+1)^{3/2}\sqrt{\log\left({8dt^{2}}/{\delta}\right)}\right).$
Substituting this bound into (3), and using the convexity of $F$ as well as the feasibility of $x_{t}\in X$, implying the inequality $G(x_{t})-G_{\operatorname{opt}}\geq 0$, we obtain
$F(z_{t})-F_{\operatorname{opt}}\leq O\left(\frac{\sqrt{\log(8dt^{2}/\delta)}}{(t+1)^{1/2-p}}\right).$
This in turn yields the bounds on $G(z_{t})-G_{\operatorname{opt}}$. Details are presented in Lemmas 6, 4 and 5.
Algorithm 2 uses the SPIDER estimator, resulting in slightly different probabilistic bounds that depend on the frequency $q$ in which we compute the full gradient and the mini-batch size $S$. Thanks to [41, Theorem 3.5], we show in Lemma 7 that with probability at least $1-\eta_{i}$, we (jointly) have
$\left\|\widehat{\nabla F}_{i}\!-\!\nabla F(x_{i})\right\|\leq O\left(\frac{\log(q)}{\sqrt{qS}}\sqrt{\log\left(\frac{8}{\eta_{i}}\right)}\right)~{}~{}\&~{}~{}\left\|\widehat{\nabla G}_{i}\!-\!\nabla G(x_{i})\right\|\leq O\left(\frac{\log(q)}{\sqrt{qS}}\sqrt{\log\left(\frac{8}{\eta_{i}}\right)}\right)$
when $1\leq i\leq q$, and
$\left\|\widehat{\nabla F}_{i}-\nabla F(x_{i})\right\|$ $\leq O\left(\sqrt{\frac{(i-s_{i})}{S(i+1)(s_{i}+1)}}\sqrt{\log\left(\frac{8}{\eta_{i}}\right)}\right)$
$\left\|\widehat{\nabla G}_{i}-\nabla G(x_{i})\right\|$ $\leq O\left(\sqrt{\frac{(i-s_{i})}{S(i+1)(s_{i}+1)}}\sqrt{\log\left(\frac{8}{\eta_{i}}\right)}\right),$
when $i>q$, where $s_{i}:=q\lfloor i/q\rfloor$ is the closest multiple of $q$ below $i$. Since the numerator of the fraction term $i-s_{i}$ is always bounded by $q$, the same aggregation trick as in Proposition 3, except ignoring the first $q$ terms (which are handled in Lemma 8), leads to an improved dependence. Specifically, we can show that, with probability $1-\delta$, we have
$\sum_{i\in[t]\setminus[q]}i\left\|\widehat{\nabla\Phi}_{i-1}-\nabla\Phi_{i-1}(x_{i-1})\right\|\leq O\left((8\sqrt{2}t-4q)\sqrt{\log\left({16t^{2}}/{\delta}\right)}\right).$
This improvement carries through to get the improved dependence on $t$ in Theorem 2; full details are given in Lemmas 8, 9, 10 and 11. ∎
Remark 1 (Sample complexity).
To attain an $\epsilon$-level optimality gap for both the outer- and inner-level problems, Algorithm 1 with $p=1/4$ requires $O(1/\epsilon^{4})$ stochastic oracle calls under the assumptions of Theorem 1. Likewise, under the assumptions of Theorem 2 and with $S=q=\lfloor\sqrt{n}\rfloor$ and $p=1/2$, Algorithm 2 achieves the same accuracy with $O(\sqrt{n}/\epsilon^{2})$ stochastic oracle calls.
3 Numerical results
We showcase the performance of our proposed algorithms in an over-parametrized regression and an $\ell_{1}$-constrained logistic regression problems. Each experiment is repeated $10$ times, and we report the average performance within a $4$-minute execution time limit. For IR-SCG, presented in Algorithm 1, we set $\sigma_{t}=\varsigma(t+1)^{-1/4}$, where $\varsigma=10$ in the first experiment and $\varsigma=0.01$ in the second experiment, along with $\alpha_{t}=2/(t+2)$. For IR-FSCG, presented in Algorithm 2, we set $S=q=\lfloor\sqrt{n}\rfloor$, $\sigma_{t}=\varsigma(\max\{t,q\}+1)^{-1/2}$, where $\varsigma=10$ in the first experiment and $\varsigma=0.01$ in the second experiment, along with $\alpha_{t}=2/(t+1)$ for every $t\geq q$ and $\alpha_{t}=\log(q)/q$ for every $t<q$. In both experiments, the linear optimization oracle admits an analytical solution; see [25, Section 4.1].
For performance comparison, we implement four algorithms: SBCGI [4, Algorithm 1] with stepsize $\gamma_{t}=0.01/(t+1)$, SBCGF [4, Algorithm 2] with constant stepsize $\gamma_{t}=10^{-5}$ (both using $K_{t}=10^{-4}/\sqrt{t+1},S=q=\lfloor\sqrt{n}\rfloor$), aR-IP-SeG [26] with long ($\gamma_{t}=10^{-2}/(t+1)^{3/4}$) and short stepsizes ($\gamma_{t}=10^{-7}/(t+1)^{3/4}$, $\rho_{t}=10^{3}(t+1)^{1/4}$, $r=0.5$), and the stochastic variant of the dynamic barrier gradient descent (SDBGD) [20] with long ($\gamma_{t}=10^{-2}$) and short ($\gamma_{t}=10^{-6}$) stepsizes. We initialized all algorithms with randomized starting points. For SBCGI and SBCGF, we generated the required initial point $x^{\prime}_{0}$ by running the SPIDER-FW algorithm [57, Algorithm 2] on the inner-level problem in (6), using a stepsize of $\gamma_{t}=0.1/(t+1)$. The initialization process terminated either after $10^{5}$ stochastic oracle queries or when the computation time exceeded 100 seconds, with the resulting point serving as $x_{0}$ for both SBCGI and SBCGF. The implementation of these algorithms rely on certain linear optimization or projection oracles. For SBCGI and SBCGF, a linear optimization oracle over the $\ell_{1}$-norm ball intersecting with a half-space is required, which we employ CVXPY [10] to solve the problems, similar to the implementation in [4, Appendix F.1]. To compute the projection onto the feasible set required for aR-IP-SeG, we used [12, Algorithm 1]. To approximate the outer- and inner-level optimal values $F_{\operatorname{opt}},G_{\operatorname{opt}}$, we again employ CVXPY to solve the inner-level problem and the reformulation (2) of the bilevel problem. All implementation details follow those in the corresponding papers.
3.1 Over-parameterized regression
We study the following bilevel optimization problem
$\begin{array}[]{cl}\displaystyle\min_{x\in{\mathbb{R}}^{d}}&\displaystyle F(x):=\frac{1}{|\mathcal{D}_{\text{val}}|}\|A_{\text{val}}x-b_{\text{val}}\|_{2}^{2}\\ \text{s.t.}&\displaystyle x\in\operatorname*{arg\,min}_{\|z\|_{1}\leq\delta}G(z):=\frac{1}{|\mathcal{D}_{\text{tr}}|}\|A_{\text{tr}}z-b_{\text{tr}}\|_{2}^{2},\end{array}$ (6)
where the goal is to minimize the validation loss by choosing among optimal solutions of the training loss constrained by the $\ell_{1}$-norm ball. We use the same training and validation datasets, $(A_{\text{tr}},b_{\text{tr}})$ and $(A_{\text{val}},b_{\text{val}})$, from the Wikipedia Math Essential dataset [43], as in [4]. This dataset consists of $n=1068$ samples and $d=730$ features.
Figure 1 shows the convergence behavior of different algorithms for the inner-level (left) and outer-level (right) optimality gaps in terms of execution time. Our proposed stochastic methods outperform existing approaches. In particular, IR-FSCG achieves high-accuracy solutions for the outer-level objective while maintaining strong inner-level feasibility, surpassing its mini-batch counterpart, SBCGI. Similarly, IR-SCG outperforms its single-sample counterpart, SBCGI. We also observe that both SBCGI and SBCGF exhibit degradation in their inner-level optimality gaps over time. This behavior aligns with the theoretical guarantees in [4], where the inner-level convergence can only be bounded by $\epsilon_{g}/2$ without monotonic improvement guarantees. Furthermore, neither aR-IP-SeG nor SDBGD makes significant progress in optimizing the outer-level objective function. Finally, the table below reports the average number of iterations and stochastic oracle calls over the 4-minute time limit for each algorithm. Our projection-free method completes more iterations within the fixed time budget, benefiting from a simpler linear minimization oracle compared to the projection-free approach in [4]. We also observe that SDBGD and IR-SCG perform a similar number of iterations, as the $\ell_{1}$-norm projection oracle has comparable computational cost to the linear minimization oracle.
Method Average # of iterations Average # of oracle calls
IR-SCG $1182617.0$ $2365234.0$
IR-FSCG $369952.5$ $28240425.0$
SBCGI $6726.9$ $7428681.9$
SBCGF $6713.4$ $7935076.2$
aR-IP-SeG (short step) $807700.2$ $3230800.8$
aR-IP-SeG (long step) $799725.0$ $3199300.0$
SDBGD (short step) $1250392.6$ $2500782.2$
SDBGD (long step) $1243720.1$ $2487440.2$
Figure 1: The inner-level optimality gap (left) and the outer-level absolute optimality gap (right) over time for different algorithms on the over-parametrized regression problem Eq. 6.
3.2 Minimal norm logistic regression
Pedregosa et al. [39] studied the following $\ell_{1}$-constrained logistic regression problem
$\operatorname*{arg\,min}_{\|x\|_{1}\leq\beta}\frac{1}{n}\sum_{i\in[n]}\psi(a_{i}^{\top}x,b_{i})+\frac{\lambda}{2}\|x\|_{2}^{2},$ (7)
where $\psi$ is the logistic loss function. In this section we consider the bilevel variant of problem (7), in which we seek the minimal $\ell_{2}$-norm solution to $\ell_{1}$-constrained logistic regression problem, that is, we aim to solve the following bilevel optimization problem
$\min_{x\in{\mathbf{R}}^{d}}$ $\quad F(x):=\frac{1}{2}\|x\|_{2}^{2}$ (8)
s.t. $\quad x\in\operatorname*{arg\,min}_{\|z\|_{1}\leq\beta}G(z):=\frac{1}{n}\sum_{i\in[n]}\psi(a_{i}^{\top}z,b_{i}).$
We note that here we treat the outer-level as deterministic. Therefore, the randomness in the implemented algorithms is inherited from the solely the inner-level problem. We evaluated the method on the Madelon dataset [21], which consists of a data matrix $A\in\mathbb{R}^{n\times d}$ with $n=2000$ samples and $d=500$ features, along with an output vector $b\in\mathbb{R}^{n}$. We set $\beta=20$ for all experiments.
Figure 2 illustrates the convergence behavior of various algorithms in terms of execution time, showing the optimality gaps for the inner-level (left) and outer-level (right) problems. The table below reports the average number of iterations and stochastic oracle calls over the 4-minute time limit for each algorithm. The overall performance trends are consistent with the previous experiment. However, all baseline stochastic methods fail to make meaningful progress within the fixed time limit. In contrast, our proposed algorithms, in particular IR-SCG, demonstrate strong performance on both the lower- and upper-level problems.
Method Average # of iterations Average # of oracle calls
IR-SCG $1726604.0$ $1726604.0$
IR-FSCG $225263.8$ $19926327.2$
SBCGI $6207.8$ $8852217.8$
SBCGF $5986.5$ $9376410.0$
aR-IP-SeG (short step) $1218831.0$ $2437662.0$
aR-IP-SeG (long step) $973268.7$ $1946537.4$
SDBGD (short step) $1623034.9$ $1623034.9$
SDBGD (long step) $1618114.5$ $1618114.5$
Figure 2: The inner-level optimality gap (left) and the outer-level absolute optimality gap (right) over time for different algorithms on the minimal norm logistic regression problem Eq. 8.
Appendix A Convergence proofs
This section provides the proofs of Theorems 1 and 2. We employ the notation $\Phi_{t}(x):=\sigma_{t}F(x)+G(x)$ and $\widehat{\nabla\Phi}_{t}:=\sigma_{t}\widehat{\nabla F}_{t}+\widehat{\nabla G}_{t}$ to denote the gradient estimate at step $t$. All norms refer to the Euclidean norm.
A.1 Preparatory lemmas
Lemma 1.
Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm 1 or Algorithm 2 with some given stepsizes $\{\alpha_{t}\}_{t\geq 0}$. If 1 holds, then for every $t\geq k$, we have
$\Phi_{t}(x_{t+1})-\Phi_{t}(x_{\operatorname{opt}})\!\leq\!(1-\alpha_{t})(\Phi_{t}(x_{t})-\Phi_{t}(x_{\operatorname{opt}}))+\frac{(L_{f}\sigma_{t}\!+\!L_{g})D^{2}\alpha_{t}^{2}}{2}+2\alpha_{t}D\left\|\nabla\Phi_{t}(x_{t})\!-\!\widehat{\nabla\Phi}_{t}\right\|.$
Proof of Lemma 1.
Recall that
$x_{t+1}=x_{t}+\alpha_{t}(v_{t}-x_{t}),\quad\text{where}\quad v_{t}\in\operatorname*{arg\,min}_{v\in X}v^{\top}\widehat{\nabla\Phi}_{t}.$
As both functions $F$ and $G$ are assumed to be convex and smooth, it is easy to show that $\Phi$ is convex and $\sigma_{t}L_{f}+L_{g}$-smooth. By smoothness of $\Phi_{t}$ and the assumption that $\operatorname{diam}(X)=D$, one can show that
$\Phi_{t}(x_{t+1})$ $\leq\Phi_{t}(x_{t})+\alpha_{t}\nabla\Phi_{t}(x_{t})^{\top}(v_{t}-x_{t})+\frac{(\sigma_{t}L_{f}+L_{g})\alpha_{t}^{2}}{2}\|v_{t}-x_{t}\|^{2}$
$\leq\Phi_{t}(x_{t})+\alpha_{t}\nabla\Phi_{t}(x_{t})^{\top}(v_{t}-x_{t})+\frac{(\sigma_{t}L_{f}+L_{g})D^{2}\alpha_{t}^{2}}{2}.$
Note that for any $w_{t}\in X$ and any gradient estimator $\widehat{\nabla\Phi}_{t}$, we have
$\nabla$ $\Phi_{t}(x_{t})^{\top}(v_{t}-x_{t})$
$=(\nabla\Phi_{t}(x_{t})-\widehat{\nabla\Phi}_{t})^{\top}(v_{t}-x_{t})+\widehat{\nabla\Phi}_{t}^{\top}(v_{t}-x_{t})$
$\leq(\nabla\Phi_{t}(x_{t})-\widehat{\nabla\Phi}_{t})^{\top}(v_{t}-x_{t})+\widehat{\nabla\Phi}_{t}^{\top}(w_{t}-x_{t})$
$=(\nabla\Phi_{t}(x_{t})-\widehat{\nabla\Phi}_{t})^{\top}(v_{t}-x_{t})+(\widehat{\nabla\Phi}_{t}-\nabla\Phi_{t}(x_{t}))^{\top}(w_{t}-x_{t})+\nabla\Phi_{t}(x_{t})^{\top}(w_{t}-x_{t})$
$\leq 2D\|\widehat{\nabla\Phi}_{t}-\nabla\Phi_{t}(x_{t})\|+\nabla\Phi_{t}(x_{t})^{\top}(w_{t}-x_{t}),$
where the first inequality follows from the definition of $v_{t}$, and the second inequality uses the Cauchy-Schwarz inequality, and the assumption that $\operatorname{diam}(X)=D$. Thus, by smoothness of $\Phi_{t}$, for any $w_{t}\in X$ and any gradient estimator $\widehat{\nabla\Phi}_{t}$, we have
$\Phi_{t}(x_{t+1})\!\leq\!\Phi_{t}(x_{t})\!+\!\alpha_{t}\nabla\Phi_{t}(x_{t})^{\top}(w_{t}\!-\!x_{t})\!+\!2D\alpha_{t}\|\widehat{\nabla\Phi}_{t}\!-\!\nabla\Phi_{t}(x_{t})\|+\frac{(\sigma_{t}L_{f}\!+\!L_{g})D^{2}\alpha_{t}^{2}}{2}.$ (9)
Let $x_{\operatorname{opt}}\in X_{\operatorname{opt}}$, and set $w_{t}=x_{\operatorname{opt}}$. Thanks to the convexity of $\Phi_{t}$, one can easily show that $\nabla\Phi_{t}(x_{t})^{\top}(x_{\operatorname{opt}}-x_{t})\leq\Phi_{t}(x_{\operatorname{opt}})-\Phi_{t}(x_{t})$. Thus, we obtain
$\Phi_{t}(x_{t+1})-\Phi_{t}(x_{\operatorname{opt}})$
$\leq\Phi_{t}(x_{t})-\Phi_{t}(x_{\operatorname{opt}})+\alpha_{t}\left(\Phi_{t}(x_{\operatorname{opt}})-\Phi_{t}(x_{t})+2D\|\widehat{\nabla\Phi}_{t}-\nabla\Phi_{t}(x_{t})\|\right)+\frac{(\sigma_{t}L_{f}+L_{g})D^{2}\alpha_{t}^{2}}{2}$
$=(1-\alpha_{t})\left(\Phi_{t}(x_{\operatorname{opt}})-\Phi_{t}(x_{t})\right)+2\alpha_{t}D\|\widehat{\nabla\Phi}_{t}-\nabla\Phi_{t}(x_{t})\|+\frac{(\sigma_{t}L_{f}+L_{g})D^{2}\alpha_{t}^{2}}{2}.$
The proof concludes by using the definition of the stepsizes $\alpha_{t}$. ∎
Lemma 2.
Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm 1 or Algorithm 2 with stepsizes $\{\alpha_{t}\}_{t\geq 0}$ given by $\alpha_{t}=2/(t+2)$ for any $t\geq k$. If 1 holds, then for each $t\geq k$, we have
$t(t+1)\left(\Phi_{t}(x_{t})-\Phi_{t}(x_{\operatorname{opt}})\right)$ $+\sum_{i\in[t]\setminus[k]}(i+1)i(\sigma_{i-1}-\sigma_{i})(F(x_{i})-F_{\operatorname{opt}})$ (10)
$\leq(k+1)k\left(\Phi_{k}(x_{k})-\Phi_{k}(x_{\operatorname{opt}})\right)+2(\sigma_{0}L_{f}+L_{g})D^{2}(t-k)$
$\quad+\sum_{i\in[t]\setminus[k]}4iD\|\widehat{\nabla\Phi}_{i-1}-\nabla\Phi_{i-1}(x_{i-1})\|.$
Proof of Lemma 2.
By definition of $\Phi_{t}$, we have
$\Phi_{t}(x_{t+1})\!-\!\Phi_{t}(x_{\operatorname{opt}})$ $=\sigma_{t+1}(F(x_{t+1})-F_{\operatorname{opt}})\!+\!G(x_{t+1})\!-\!G_{\operatorname{opt}}\!+\!(\sigma_{t}-\sigma_{t+1})(F(x_{t+1})-F_{\operatorname{opt}})$
$=\Phi_{t+1}(x_{t+1})-\Phi_{t+1}(x_{\operatorname{opt}})+(\sigma_{t}-\sigma_{t+1})(F(x_{t+1})-F_{\operatorname{opt}}).$
To simplify the derivations, we introduce the shorthands $h_{t}:=\Phi_{t}(x_{t})-\Phi_{t}(x_{\operatorname{opt}})$ and
$\eta_{t}:=\frac{2(L_{f}\sigma_{t}+L_{g})D^{2}}{(t+2)^{2}}+\frac{4D}{t+2}\|\widehat{\nabla\Phi}_{t}-\nabla\Phi_{t}(x_{t})\|-(\sigma_{t}-\sigma_{t+1})(F(x_{t+1})-F_{\operatorname{opt}}).$
Applying Lemma 1 with stepsize $\alpha_{t}=2/(t+2)$ for all $t\geq k$ gives the recursion $h_{t+1}\leq\frac{t}{t+2}h_{t}+\eta_{t}$ for any $t\geq k$. This can be re-expressed as
$(t+1)th_{t}\leq t(t-1)h_{t-1}+t(t+1)\eta_{t-1}$
for any $t\geq k+1$. Thus, for $t\geq k+1$, it holds that
$(t+1)th_{t}-(k+1)kh_{k}=\sum_{i\in[t]\setminus[k]}\left((i+1)ih_{i}-(i-1)ih_{i-1}\right)\leq\sum_{i\in[t]\setminus[k]}(i+1)i\eta_{i-1}.$
By definition of $h_{t}$ and $\eta_{t}$, and using the inequalities $i/(i+1)<1$ and $\sigma_{i-1}\leq\sigma_{0}$ for $i\in[t]\setminus[k]$, we arrive at
$(t+1)t$ $\left(\Phi_{t}(x_{t})-\Phi_{t}(x_{\operatorname{opt}})\right)-(k+1)k\left(\Phi_{k}(x_{k})-\Phi_{k}(x_{\operatorname{opt}})\right)$
$\textstyle\leq 2(\sigma_{0}L_{f}+L_{g})D^{2}(t-k)-\sum_{i\in[t]\setminus[k]}(i+1)i(\sigma_{i-1}-\sigma_{i})(F(x_{i})-F_{\operatorname{opt}})$
$\quad\textstyle+\sum_{i\in[t]\setminus[k]}4iD\|\widehat{\nabla\Phi}_{i-1}-\nabla\Phi_{i-1}(x_{i-1})\|,$
which implies (10). This completes the proof. ∎
Proposition 1.
Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm 1 or Algorithm 2 with stepsizes $\{\alpha_{t}\}_{t\geq 0}$ given by $\alpha_{t}=2/(t+2)$ for any $t\geq k$. If 1 holds, then for each $t\geq k$, we have
[…]
and
$(t+1)t\sigma_{t}(F(x_{t})-F_{\operatorname{opt}})$ $+\sum_{i\in[t]\setminus[k]}(i+1)i(\sigma_{i-1}-\sigma_{i})(F(x_{i})-F_{\operatorname{opt}})$
$\leq(k+1)k\left(\Phi_{k}(x_{k})-\Phi_{k}(x_{\operatorname{opt}})\right)+2(\sigma_{0}L_{f}+L_{g})D^{2}(t-k)$ (12)
$\quad+\sum_{i\in[t]\setminus[k]}4iD\|\widehat{\nabla\Phi}_{i-1}-\nabla\Phi_{i-1}(x_{i-1})\|.$
Proof of Proposition 1.
By definition of the function $\Phi_{t}$, the first term in Eq. 10 satisfies
$t(t\!+\!1)\left(\Phi_{t}(x_{t})\!-\!\Phi_{t}(x_{\operatorname{opt}})\right)\!=\!t(t\!+\!1)\left(G(x_{t})\!-\!G(x_{\operatorname{opt}})\right)\!+\!t(t\!+\!1)\sigma_{t}\left(F(x_{t})\!-\!F(x_{\operatorname{opt}})\right).$ (13)
By construction, $x_{0},v_{t}\in X$ Hence, $x_{t+1}\in X$ as it is a convex combination of $x_{t}$ and $v_{t}$, and $X$ is assumed to be convex. By a simple induction argument, one can easily show that the sequence $\{x_{t}\}_{t\geq 0}\in X$. Therefore, we have $\min_{x\in X}F(x)-F_{\operatorname{opt}}\leq F(x_{t})-F_{\operatorname{opt}}$ for all $t\geq 0$. Using this bound together with the identity in Eq. 13, we can further lower bound the left-hand side of inequality (10), and arrive at the bound (11). This completes the proof of the first claim.
For the second claim, since $\{x_{t}\}_{t\geq 0}\in X$ is feasible in the lower-level problem, it follows directly that $G(x_{t})\geq G_{\operatorname{opt}}$. Combining this inequality with (13) and  (10), we obtain Eq. 12, which concludes the proof. ∎
A.2 Proof of Theorem 1
The proof proceeds by applying Proposition 1 with $k=0$. In this case, the main random component in Eq. 11 and Eq. 12 is the term $\sum_{i\in[t]}4iD\|\widehat{\nabla\Phi}_{i-1}-\nabla\Phi_{i-1}(x_{i-1})\|$. We first derive a probabilistic upper bound for this term, then carefully set the parameters in Algorithm 1 to establish the convergence rate for both the lower- and upper-level problems.
We begin by bounding the gradient estimator used in Algorithm 1.
Lemma 3.
Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm 1 with stepsize $\alpha_{t}=2/(t+2)$ for any $t\geq 0$. If 1 holds, then for any $t\geq 1$, given $\delta\in(0,1)$, with probability at least $1-\delta$, for some absolute constant $c$, it (jointly) holds that
$\|\widehat{\nabla F}_{t}-\nabla F(x_{t})\|$ $\leq c\sqrt{2}\left(4L_{f}D+3\sigma_{f}\right)\frac{\sqrt{\log(4d/\delta)}}{(t+1)^{1/2}}$ (14)
$\|\widehat{\nabla G}_{t}-\nabla G(x_{t})\|$ $\leq c\sqrt{2}\left(4L_{g}D+3\sigma_{g}\right)\frac{\sqrt{\log(4d/\delta)}}{(t+1)^{1/2}}.$
Proof of Lemma 3.
The proof closely follows the argument in [4, Lemma 4.1]. In particular, the bounds in [4, Lemma B.1] still hold under the modified stepsize $\alpha_{t}=2/(t+2)$, instead of $1/(t+1)$. The constant inside the logarithmic term changes from 6 to 4, as we now apply a union bound over two events rather than three. Additionally, the constants of the upper bounds is doubled due to use of the modified stepsize. We omit the details for brevity. ∎
Proposition 2.
Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm 1 with $\alpha_{t}=2/(t+2)$ for any $t\geq 0$. If 1 holds and the parameters $\{\sigma_{t}\}_{t\geq 0}$ are non-increasing and positive, then given $\delta\in(0,1)$, with probability at least $1-\delta$, for some absolute constant $c$ and any $t\geq 1$, we have
$\sum_{i\in[t]}i\|\widehat{\nabla\Phi}_{i-1}-\nabla\Phi_{i-1}(x_{i-1})\|\leq c\left(4(\sigma_{0}L_{f}+L_{g})D+3(\sigma_{0}\sigma_{f}+\sigma_{g})\right)(t+1)^{3/2}\sqrt{\log\left({8dt^{2}}/{\delta}\right)}.$
Proof of Proposition 2.
Given $i\geq 1$ and $\delta\in(0,1)$, by Lemma 3, with probability at least $1-\delta/i(i+1)$, we have
$i\|\widehat{\nabla\Phi}_{i-1}-\nabla\Phi_{i-1}(x_{i-1})\|$ $\leq i(\sigma_{i}\|\widehat{\nabla F}_{i-1}-\nabla F(x_{i-1})\|+\|\widehat{\nabla G}_{i-1}-\nabla G(x_{i-1})\|)$
$\leq i(\sigma_{0}\|\widehat{\nabla F}_{i-1}-\nabla F(x_{i-1})\|+\|\widehat{\nabla G}_{i-1}-\nabla G(x_{i-1})\|)$
$\leq c\sqrt{2}\left(4(\sigma_{0}L_{f}+L_{g})D+3(\sigma_{0}\sigma_{f}+\sigma_{g})\right)\sqrt{i\log\left({4di(i+1)}/{\delta}\right)}$
Thus, using the union bound, with probability at least $1-\delta\sum_{i\geq 1}1/i(i+1)=1-\delta$, for any $i\in[t]$, we have
$\sum_{i\in[t]}i\|\widehat{\nabla\Phi}_{i-1}-\nabla\Phi_{i-1}(x_{i-1})\|$ $\leq\sum_{i\in[t]}c\sqrt{2}\left(4(\sigma_{0}L_{f}+L_{g})D+3(\sigma_{0}\sigma_{f}+\sigma_{g})\right)\sqrt{i\log\left({8dt^{2}}/{\delta}\right)}$
$\leq c\left(4(\sigma_{0}L_{f}+L_{g})D+3(\sigma_{0}\sigma_{f}+\sigma_{g})\right)(t+1)^{3/2}\sqrt{\log\left({8dt^{2}}/{\delta}\right)},$
where the second inequality follows from bounding $\sum_{i\in[t]}\sqrt{i}\leq\tfrac{2}{3}(t+1)^{3/2}$ via the Riemann sum approximation, and observing that $2\sqrt{2}\leq 3$. ∎
To establish the desired convergence results for Algorithm 1, we need to impose certain conditions on the regularization parameters $\{\sigma_{t}\}_{t\geq 0}$, stated below.
Condition 1.
The regularization parameters $\{\sigma_{t}\}_{t\geq 0}$ are non-increasing, positive, and converge to 00.
Condition 2.
There exists $L\in\mathbb{R}$ such that
$\lim_{t\to\infty}t\left(\frac{\sigma_{t}}{\sigma_{t+1}}-1\right)=L.$
Condition 3.
$(t+1)\sigma_{t+1}^{2}>t\sigma_{t}^{2}$ for any $t\geq 0$, and $\log(t)=o(t\sigma_{t}^{2})$ as $t\to\infty$.
The following parameters will appear in our analysis and convergence rates
$C$ $:=\sup_{t\geq 1}\left\{\left(F_{\operatorname{opt}}\!-\!\min_{x\in X}F(x)\right)\left(1\!+\!\frac{\sum_{i\in[t]}(i\!+\!1)i(\sigma_{i-1}\!-\!\sigma_{i})}{(t+1)t\sigma_{t}}\right)\!+\!\frac{2(\sigma_{0}L_{f}\!+\!L_{g})D^{2}}{(t+1)\sigma_{t}}\right\},$ (15a)
$\bar{C}_{\delta}$ $:=\sup_{t\geq 1}\left\{\frac{8c\left(4(\sigma_{0}L_{f}+L_{g})D^{2}+3(\sigma_{0}\sigma_{f}+\sigma_{g})D\right)}{(t+1)^{1/2}\sigma_{t}}\sqrt{\log\left({8dt^{2}}/{\delta}\right)}\right\},$ (15b)
$C_{\delta}$ $:=\bar{C}_{\delta}+C,$ (15c)
$V$ $:=\sup_{t\geq 1}\left\{\frac{\sum_{i\in[t]}(i+1)i(\sigma_{i-1}-\sigma_{i})\sigma_{i}}{(t+1)t\sigma_{t}^{2}}\right\}.$ (15d)
While it is not obvious a priori, Lemma 16 in Section B.4 guarantees that these quantities remain finite for any parameter choice satisfying 1 to 3. In particular, 1 to 3 hold when $\sigma_{t}=\varsigma(t+1)^{-p}$ for any $p\in(0,1)$, which is what is used in Theorems 1 and 2, and Lemma 6 provides explicit bounds for this case. We now analyze the sequence $\{x_{t}\}_{t\geq 0}$ in Algorithm 1 for the inner-level problem.
Lemma 4.
Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm 1 with $\alpha_{t}=2/(t+2)$ for any $t\geq 0$, and let $C_{\delta}$ be defined as in (15a). If 1 and 1 to 3 hold, with probability at least $1-\delta$, it holds that
$G(x_{t})-G_{\operatorname{opt}}\leq C_{\delta}\sigma_{t},$ (16)
for any $t\geq 0$.
Proof of Lemma 4.
Combining (11) from Proposition 1 with the probabilistic bound in Proposition 2, and after some straightforward calculations, with probability at least $1-\delta$, we have
$G(x_{t})-G_{\operatorname{opt}}$ $\leq\frac{8c\left(4(\sigma_{0}L_{f}+L_{g})D^{2}+3(\sigma_{0}\sigma_{f}+\sigma_{g})D\right)}{(t+1)^{1/2}}\sqrt{\log\left({8dt^{2}}/{\delta}\right)}$
$\quad+\frac{2(\sigma_{0}L_{f}+L_{g})D^{2}}{(t+1)}\!+\!\left(F_{\operatorname{opt}}\!-\!\min_{x\in X}F(x)\right)\left(\sigma_{t}\!+\!\frac{\sum_{i\in[t]}(i+1)i(\sigma_{i-1}\!-\!\sigma_{i})}{(t+1)t}\right).$
By definition of $C_{\delta}$, the right-hand side is at most $C_{\delta}\sigma_{t}$ for any $t\geq 1$. This completes the proof. ∎
We next analyze the convergence of the sequence $\{z_{t}\}_{t\geq 0}$ in Algorithm 1 for both outer- and inner-level problems.
Lemma 5.
Let $\{z_{t}\}_{t\geq 1}$ denote the iterates generated by Algorithm 1 with $\alpha_{t}=2/(t+2)$ for any $t\geq 0$, and let $C_{\delta},V$ be defined as in Eq. 15. If 1 and 1 to 3 hold, then with probability at least $1-\delta$, it (jointly) holds that
$F(z_{t})-F_{\operatorname{opt}}$ $\leq\frac{2(\sigma_{0}L_{f}+L_{g})D^{2}}{(t+1)\sigma_{t}}+\frac{4c\left(4(\sigma_{0}L_{f}+L_{g})D+3(\sigma_{0}\sigma_{f}+\sigma_{g})\right)}{(t+1)^{1/2}\sigma_{t}}\sqrt{\log\left({8dt^{2}}/{\delta}\right)},$
$G(z_{t})-G_{\operatorname{opt}}$ $\leq C_{\delta}(1+V)\sigma_{t},$
for any $t\geq 1$.
Proof of Lemma 5.
Recall from Algorithm 1 that we have defined
$\begin{cases}S_{t+1}:=(t+2)(t+1)\sigma_{t+1}+\sum_{i\in[t+1]}(i+1)i(\sigma_{i-1}-\sigma_{i})\\[8.61108pt] z_{t+1}:=\frac{1}{S_{t+1}}\left((t+2)(t+1)\sigma_{t+1}x_{t+1}+\sum_{i\in[t+1]}(i+1)i(\sigma_{i-1}-\sigma_{i})x_{i}\right),\end{cases}$
thus $z_{t}$ is simply a convex combination of $x_{0},\dots,x_{t}$ for every $t\geq 1$. Therefore, as $F$ is convex, we can apply Jensen’s inequality to the left-hand side of the inequality (12) with $k=0$, and after some tedious calculation, we arrive at
$F(z_{t})-F_{\operatorname{opt}}$ $\leq\frac{2(\sigma_{0}L_{f}+L_{g})D^{2}t}{S_{t}}+\sum_{i\in[t]}\frac{4iD\|\widehat{\nabla\Phi}_{i-1}-\nabla\Phi_{i-1}(x_{i-1})\|}{S_{t}}.$
Using the inequality $S_{t}\geq(t+1)t\sigma_{t}$, which holds thanks to 1, and applying Proposition 2, with probability at least $1-\delta$, we have
$F(z_{t})-F_{\operatorname{opt}}$ $\leq\frac{2(\sigma_{0}L_{f}+L_{g})D^{2}}{(t+1)\sigma_{t}}+\frac{8c\left(4(\sigma_{0}L_{f}+L_{g})D+3(\sigma_{0}\sigma_{f}+\sigma_{g})\right)}{(t+1)^{1/2}\sigma_{t}}\sqrt{\log\left({8dt^{2}}/{\delta}\right)}.$
This completes the proof of the first claim.
For the second claim, we follow the same procedure. In particular, applying Jensen’s inequality with respect to the convex function $G$ to the left-hand side of (16) and using the inequality $S_{t}\geq(t+1)t\sigma_{t}$, with probability at least $1-\delta$, we have
$G(z_{t})-G_{\operatorname{opt}}\leq\frac{C_{\delta}}{t(t+1)\sigma_{t}}\left((t+1)t\sigma_{t}^{2}+\sum_{i\in[t]}(i+1)i(\sigma_{i-1}-\sigma_{i})\sigma_{i}\right).$
The proof concludes by using the definition of $V$. ∎
Lemma 5 establishes a convergence result in terms of the regularization parameters $\{\sigma_{t}\}_{t\geq 0}$. The next lemma specifies an update rule for these parameters and provides bounds on the quantities introduced in Eq. 15.
Lemma 6.
Consider the sequence $\sigma_{t}:=\varsigma(t+1)^{-p}$ for $t\geq 0$ and the quantities defined in Eq. 15.
  • (i)
    If $p\in(0,1)$, then $\{\sigma_{t}\}_{t\geq 0}$ satisfies 1 and 2 with $L=p$. Furthermore,
    $C\leq(1+2p)\left(F_{\operatorname{opt}}-\min_{x\in X}F(x)\right)+\frac{2(\varsigma L_{f}+L_{g})D^{2}}{\varsigma},\quad V\leq\frac{2p}{\min\{1,2(1-p)\}}.$
  • (ii)
    If $p\in(0,1/2)$, then $\{\sigma_{t}\}_{t\geq 0}$ satisfies 3, and we have
    $\bar{C}_{\delta}\leq 4c\left(8\left(L_{f}+\frac{L_{g}}{\varsigma}\right)D^{2}+3\left(\sigma_{f}+\frac{\sigma_{g}}{\varsigma}\right)D\right)\sqrt{\frac{2}{(1-2p)e}+\log\left(\frac{8d}{\delta}\right)}.$
Proof of Lemma 6.
As for assertion (i), it is trivial to see that the sequence $\{\sigma_{t}\}_{t\geq 0}$ satisfies Conditions 1. To validate 2, observe that
$\lim_{t\to\infty}t\left(\frac{\sigma_{t}}{\sigma_{t+1}}-1\right)=\lim_{t\to\infty}t\left(\left(1+\frac{1}{t+1}\right)^{p}-1\right)=\lim_{t\to\infty}\frac{t}{t+1}\cdot\lim_{\delta\to 0}\frac{(1+\delta)^{p}-1}{\delta}=p,$
where the last equality holds as the derivative of $y^{p}$ at $y=1$ equals $p$. We next establish the bound for $C$. For $t>1$, it follows from the mean value theorem that there exists $b_{t}\in(t,t+1)$ such that
$t^{-p}-(t+1)^{-p}=-pb_{t}^{-(p+1)}(t-t-1)=pb_{t}^{-(p+1)}\leq pt^{-(p+1)}.$
Hence, we observe that
$\textstyle\sum_{i\in[t]}(i+1)i(\sigma_{i-1}-\sigma_{i})\leq 2\varsigma p\sum_{i\in[t]}i^{1-p}\leq{2\varsigma p}t^{2-p},$
where the first inequality is due to $t+1\leq 2t$ and the second inequality holds because $t^{1-p}$ is an increasing function in $t$. This implies
$\frac{\sum_{i\in[t]}(i+1)i(\sigma_{i-1}-\sigma_{i})}{(t+1)t\sigma_{t}}\leq\frac{2pt^{2-p}}{t(t+1)^{1-p}}\leq 2p.$
Since $\min_{t\geq 0}(t+1)\sigma_{t}=\varsigma$, we obtain
$C\leq(1+2p)\left(F_{\operatorname{opt}}-\min_{x\in X}F(x)\right)+\frac{2(\varsigma L_{f}+L_{g})D^{2}}{\varsigma}.$
We next establish the bound for $V$. Using similar arguments, we observe that
$\textstyle\sum_{i\in[t]}(i+1)i(\sigma_{i-1}-\sigma_{i})\sigma_{i}\leq 2\varsigma^{2}p\sum_{i\in[t]}i^{1-2p}.$
If $1-2p\geq 0$, then $t^{1-2p}$ is an increasing function in $t$ and thus
$\sum_{i\in[t]}(i+1)i(\sigma_{i-1}-\sigma_{i})\sigma_{i}\leq 2p\varsigma^{2}t^{2(1-p)}.$
Dividing both sides by $t(t+1)\sigma_{t}^{2}$, we deduce that $V\leq 2p$. When $01-2p<$ 0, we have
$\textstyle\sum_{i\in[t]}(i+1)i(\sigma_{i-1}-\sigma_{i})\sigma_{i}\leq 2p\varsigma^{2}\left(1+\frac{1}{2-2p}\left(t^{2-2p}-1\right)\right)\leq\varsigma^{2}\frac{p}{1-p}t^{2(1-p)},$
where the first inequality is due to the Riemann sum approximation $\sum_{i\in[t]}i^{1-2p}\!\leq\!1\!+\!\int_{i=1}^{t}i^{1-2p}\mathrm{d}i$. Dividing both sides by $t(t+1)\sigma_{t}^{2}$, we deduce that $V\leq{p}/({1-p})$, thus the claim follows.
As for assertion (ii), it is trivial to see that 3 holds. We thus focus on bounding $\bar{C}_{\delta}$. Using the fact that $\sup_{t\geq 1}{\log(t)}/{t}={1}/{e},$ for any $t\geq 1$, we may conclude that
$\frac{8c\left(4(\sigma_{0}L_{f}+L_{g})D^{2}+3(\sigma_{0}\sigma_{f}+\sigma_{g})D\right)}{(t+1)^{1/2}\sigma_{t}}\sqrt{\log\left({8dt^{2}}/{\delta}\right)}$
$=8c\left(4\left(L_{f}+\frac{L_{g}}{\varsigma}\right)D^{2}+3\left(\sigma_{f}+\frac{\sigma_{g}}{\varsigma}\right)D\right)\sqrt{\frac{1}{(t+1)^{1-2p}}\left(\frac{2}{1-2p}\log\left(t^{1-2p}\right)+\log\left(\frac{8d}{\delta}\right)\right)}$
$\leq 8c\left(4\left(L_{f}+\frac{L_{g}}{\varsigma}\right)D^{2}+3\left(\sigma_{f}+\frac{\sigma_{g}}{\varsigma}\right)D\right)\sqrt{\frac{2}{(1-2p)e}+\log\left(\frac{8d}{\delta}\right)}.$
The claim then follows from the definition of $\bar{C}_{\delta}$. ∎
Proof of Theorem 1.
This is an immediate consequence of Lemmas 5 and 6. ∎
A.3 Proof of Theorem 2
The proof proceeds by applying Proposition 1 with $k=q$. Unlike the proof of Theorem 3, here we have two main random components in Eq. 11 and Eq. 12: $(q+1)q\left(\Phi_{q}(x_{q})-\Phi_{q}(x_{\operatorname{opt}})\right)$ and $\sum_{i\in[t]\setminus[q]}4iD\|\widehat{\nabla\Phi}_{i-1}-\nabla\Phi_{i-1}(x_{i-1})\|$. We first derive probabilistic upper bounds for these terms, then carefully set the parameters in Algorithm 2 to establish the convergence rate for both lower- and upper-level problems. In the following we use the notation
$F(x)=\sum_{i\in[n]}f_{i}(x),\quad G(x)=\sum_{i\in[n]}g_{i}(x).$
For simplicity, we assume that $F$ and $G$ have the same size support, though our arguments are easily extended to the more general case. Also, given ${\cal S}\subseteq[n]$, we write
$f_{{\cal S}}(x):=\frac{1}{|{\cal S}|}\sum_{i\in{\cal S}}f_{i}(x),\quad g_{{\cal S}}(x):=\frac{1}{|{\cal S}|}\sum_{i\in{\cal S}}f_{i}(x).$
We begin with a probabilistic bound on the gradient estimator used in Algorithm 2.
Lemma 7.
Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm 2 with $\alpha_{t}=\log(q)/q$ for any $0\leq t<q$ and $\alpha_{t}=2/(t+2)$ for any $t\geq q$. If 1 holds, then for any $t\geq 0$, given $\delta\in(0,1)$, with probability at least $1-\delta$, it (jointly) holds that
$\left\|\widehat{\nabla F}_{t}-\nabla F(x_{t})\right\|$ $\leq 8L_{f}D\frac{\log(q)}{\sqrt{qS}}\sqrt{\log\left(\frac{8}{\delta}\right)},$ (17)
$\left\|\widehat{\nabla G}_{t}-\nabla G(x_{t})\right\|$ $\leq 8L_{g}D\frac{\log(q)}{\sqrt{qS}}\sqrt{\log\left(\frac{8}{\delta}\right)},$
provided that $0\leq t<q$, and
$\left\|\widehat{\nabla F}_{t}-\nabla F(x_{t})\right\|$ $\leq 8L_{f}D\sqrt{\frac{(t-s_{t})}{S(t+1)(s_{t}+1)}}\sqrt{\log\left(\frac{8}{\delta}\right)}$ (18)
$\left\|\widehat{\nabla G}_{t}-\nabla G(x_{t})\right\|$ $\leq 8L_{g}D\sqrt{\frac{(t-s_{t})}{S(t+1)(s_{t}+1)}}\sqrt{\log\left(\frac{8}{\delta}\right)},$
where $s_{t}:=q\lfloor t/q\rfloor$, provided that $t>q$.
Proof of Lemma 7.
If $t=s_{t}$, we have $\widehat{\nabla F}_{t}=\nabla F(x_{t})$. Otherwise, let ${\cal S}_{t}\subset[n]$ be the index set of size $S$ chosen at iteration $t$, and for $i\in{\cal S}_{t}$ define
$\epsilon_{t,i}:=\frac{1}{S}\left(\nabla f_{i}(x_{t})-\nabla f_{i}(x_{t-1})-\nabla F(x_{t})+\nabla F(x_{t-1})\right).$
From the update rule for $x_{t}$, we have $\|x_{t}-x_{t-1}\|=\alpha_{t-1}\|v_{t-1}-x_{t-1}\|\leq\alpha_{t-1}D$ for any $t\geq 1$. As a result,
$\|\epsilon_{t,i}\|$ $\leq\frac{1}{S}\left(\|\nabla f_{i}(x_{t})-\nabla f_{i}(x_{t-1})\|+\|\nabla F(x_{t})-\nabla F(x_{t-1})\|\right)\leq\frac{2L_{f}D}{S}\alpha_{t-1},$
for any $t\in\{s_{t}+1,\dots,s_{t}+q\}$ and $i\in[S_{t}]$. For any $t$ such that $t\neq s_{t}$, we have from the definition of $\widehat{\nabla F}_{t}$ that
$\widehat{\nabla F}_{t}-\nabla F(x_{t})=\widehat{\nabla F}_{t-1}-\nabla F(x_{t-1})+\sum_{i\in[S]}\epsilon_{t,i}.$
Thus, we deduce that
$\left\|\widehat{\nabla F}_{t}-\nabla F(x_{t})\right\|=\left\|\sum_{s=s_{t}+1}^{t}\sum_{i\in{\cal S}_{s}}\epsilon_{s,i}\right\|.$
Since $\sum_{s=s_{t}+1}^{t}\sum_{i\in{\cal S}_{s}}\epsilon_{s,i}$ is a martingale with bounded increments, we apply a concentration inequality [41, Theorem 3.5] to get
$P\left(\left\|\widehat{\nabla F}_{t}-\nabla F(x_{t})\right\|\geq\lambda\right)$ $\leq 4\exp\left(-\frac{\lambda^{2}}{4\sum_{s=s_{t}+1}^{t}\frac{4L_{f}^{2}D^{2}}{S}\alpha_{s-1}^{2}}\right).$
If $1\leq t\leq q$, we have
$\sum_{s=s_{t}+1}^{t}\frac{4L_{f}^{2}D^{2}}{S}\alpha_{s-1}^{2}\leq\frac{4L_{f}^{2}D^{2}q}{S}\left(\frac{\log(q)}{q}\right)^{2}=4L_{f}^{2}D^{2}\frac{\log^{2}(q)}{qS},$
and thus
$P\left(\left\|\widehat{\nabla F}_{t}-\nabla F(x_{t})\right\|\geq\lambda\right)\leq 4\exp\left(-\frac{\lambda^{2}qS}{16L_{f}^{2}D^{2}\log^{2}(q)}\right).$
If $t>q$, we observe that
$\sum_{s=s_{t}+1}^{t}\alpha_{s-1}^{2}=4\sum_{s=s_{t}+1}^{t}\frac{1}{(s+1)^{2}}\leq 4\int_{s_{t}}^{t}\frac{du}{(u+1)^{2}}=\frac{4(t-s_{t})}{(t+1)(s_{t}+1)}.$
Then, we have
$P\left(\left\|\widehat{\nabla F}_{t}-\nabla F(x_{t})\right\|\geq\lambda\right)\leq 4\exp\left(-\frac{\lambda^{2}S(t+1)(s_{t}+1)}{64L_{f}^{2}D^{2}(t-s_{t})}\right).$
Arguing similarly for $\widehat{\nabla G}_{t}$, we have
$P\left(\left\|\widehat{\nabla G}_{t}-\nabla G(x_{t})\right\|\geq\lambda\right)\leq 4\exp\left(-\frac{\lambda^{2}qS}{16L_{g}^{2}D^{2}\log^{2}(q)}\right),$
if $1\leq t\leq q$ and
$P\left(\left\|\widehat{\nabla G}_{t}-\nabla G(x_{t})\right\|\geq\lambda\right)\leq 4\exp\left(-\frac{\lambda^{2}S(t+1)(s_{t}+1)}{64L_{g}^{2}D^{2}(t-s_{t})}\right),$
if $t>q$. Given $\delta\in(0,1)$ and $1\leq t\leq q$, setting the right hand side $=\delta/2$ and solving for $\lambda$, we have
$P\left(\left\|\widehat{\nabla F}_{t}-\nabla F(x_{t})\right\|\geq 8L_{f}D\frac{\log(q)}{\sqrt{qS}}\sqrt{\log\left(\frac{8}{\delta}\right)}\right)\leq\frac{\delta}{2}$
and
$P\left(\left\|\widehat{\nabla G}_{t}-\nabla G(x_{t})\right\|\geq 8L_{g}D\frac{\log(q)}{\sqrt{qS}}\sqrt{\log\left(\frac{8}{\delta}\right)}\right)\leq\frac{\delta}{2}.$
Applying union bound, we deduce that with probability at least $1-\delta$, (17) holds. Arguing similarly for $t>q$, we also deduce that (18) holds with probability at least $1-\delta$. ∎
To establish convergence results for Algorithm 2, we replace 3 with the following.
Condition 4.
$\sigma_{t}=\sigma_{0}$ for $0\leq t\leq q$, $(t+1)\sigma_{t+1}>t\sigma_{t}$ for any $t\geq 0$, and $\log(t)=o(t\sigma_{t})$ as $t\to\infty$.
Note that $\sigma_{t}$ is fixed during the initialization phase, and $\sigma_{t}^{2}$ in 3 is replaced by $\sigma_{t}^{2}$ in 4. Unlike the proof of Theorem 1, where we set $k=0$, the proof of Theorem 2 requires setting $k=q$. We therefore begin by analyzing the sequence $\{x_{t}\}_{t\geq 0}$ generated by Algorithm 2 over the first $q$ iterations.
Lemma 8.
Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm 2 with $\alpha_{t}=\log(q)/q$ for any $0\leq t<q$ and $S=q$. If 1 and 4 hold, we have then given $\delta\in(0,1)$, with probability at least $1-\delta\sum_{0\leq i<q}1/(i+1)(i+2)$, for any $t\geq q$, it holds that
$\Phi_{q}(x_{q})-\Phi_{0}(x_{\operatorname{opt}})$ $\leq\left(1-\frac{\log(q)}{q}\right)^{q}\left(\Phi_{0}(x_{0})-\Phi_{0}(x_{\operatorname{opt}})\right)$
$\quad+(L_{f}\sigma_{0}+L_{g})D^{2}\left(\frac{1}{2}+16\sqrt{\log\left(\frac{16t^{2}}{\delta}\right)}\right)\frac{\log(q)}{q}.$
Proof of Lemma 8.
By Lemma 1, for any $0\leq i<q$, we have
[…]
By 4, we have $\alpha_{0}=\dots=\alpha_{q-1}$ and
$\sigma_{0}=\dots=\sigma_{q}\quad\implies\Phi_{0}=\dots=\Phi_{q},$
where the implication follows from the definition of function $\Phi_{t}$. Thus, we may conclude that
$\Phi_{q}(x_{q})-\Phi_{q}(x_{\operatorname{opt}})$
$\leq(1-\alpha_{0})(\Phi_{0}(x_{q-1})-\Phi_{0}(x_{\operatorname{opt}}))+\frac{(L_{f}\sigma_{0}+L_{g})D^{2}\alpha_{0}^{2}}{2}+2D\alpha_{0}\|\nabla\Phi_{q-1}(x_{q-1})-\widehat{\nabla\Phi}_{q-1}\|$
$\leq(1-\alpha_{0})^{q}\left(\Phi_{0}(x_{0})-\Phi_{0}(x_{\operatorname{opt}})\right)+\sum_{i=0}^{q-1}\frac{(L_{f}\sigma_{0}+L_{g})D^{2}\alpha_{0}^{2}}{2}(1-\alpha_{0})^{i}$
$\quad+\sum_{i=0}^{q-1}2D\alpha_{0}\|\nabla\Phi_{i}(x_{i})-\widehat{\nabla\Phi}_{i}\|(1-\alpha_{0})^{q-1-i},$
where both inequalities are direct consequences of 4 and (19). Note that
$\sum_{i=0}^{q-1}(1-\alpha_{0})^{q-1-i}=\sum_{i=0}^{q-1}(1-\alpha_{0})^{i}\leq\sum_{i=0}^{\infty}(1-\alpha_{0})^{i}=\frac{1}{\alpha_{0}},$
where the last equality follows from convergence of geometric series. We thus arrive at
[…]
Note that by the definition of $\Phi_{i}$ and the triangle inequality, we have
$\|\nabla\Phi_{i}(x_{i})-\widehat{\nabla\Phi}_{i}\|\leq\sigma_{0}\|\widehat{\nabla F}_{i}-\nabla F(x_{i})\|+\|\widehat{\nabla G}_{i}-\nabla G(x_{i})\|.$
Hence, applying Lemma 7 along with a union bound argument implies that, for any $\delta\in(0,1)$, with probability at least $1-\delta\sum_{0\leq i<q}1/(i+1)(i+2)$, we have
$\sum_{i=0}^{q-1}\|\nabla\Phi_{0}(x_{i})-\widehat{\nabla\Phi}_{i}\|\leq\sum_{i=0}^{q-1}8D(\sigma_{0}L_{f}+L_{g})\frac{\log(q)}{\sqrt{qS}}\sqrt{\log\left(\frac{8(i+1)(i+2)}{\delta}\right)}.$
Plugging the above probabilistic bound into Eq. 20, noting that $t\geq q$ and $S=q$, and using the definition of $\alpha_{0}$ complete the proof. ∎
We next analyze the sequence $\{x_{t}\}_{t\geq 0}$ generated by Algorithm 2 after the first $q$ iterations.
Proposition 3.
Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm 2 with $\alpha_{t}=\log(q)/q$ for any $0\leq t<q$, $\alpha_{t}=2/(t+2)$ for any $t\geq q$ and $S=q$. If 1 holds and the parameters $\{\sigma_{t}\}_{t\geq 0}$ are non-increasing and positive, then given $\delta\in(0,1)$, with probability at least $1-\delta\sum_{i\in[t]\setminus[q]}1/i(i+1)$, for any $t>q$, it holds that
$\sum_{i\in[t]\setminus[q]}i\|\widehat{\nabla\Phi}_{i-1}-\nabla\Phi_{i-1}(x_{i-1})\|\leq\left(8\sqrt{2}t-4q\right)(\sigma_{0}L_{f}+L_{g})D\sqrt{\log\left(\frac{16t^{2}}{\delta}\right)}.$ (21)
Proof of Proposition 3.
For any $i\geq 1$, we have
$\|\widehat{\nabla\Phi}_{i-1}-\nabla\Phi_{i-1}(x_{i-1})\|$ $\leq(\sigma_{i-1}\|\widehat{\nabla F}_{i-1}-\nabla F(x_{i-1})\|+\|\widehat{\nabla G}_{i-1}-\nabla G(x_{i-1})\|)$
$\leq(\sigma_{0}\|\widehat{\nabla F}_{i-1}-\nabla F(x_{i-1})\|+\|\widehat{\nabla G}_{i-1}-\nabla G(x_{i-1})\|).$ (22)
Given $\delta\in(0,1)$ and $q<i\leq 2q$, by Lemma 7 and the inequality Eq. 22, with probability at least $1-\delta/i(i+1)$, we have
$i\|\widehat{\nabla\Phi}_{i-1}-\nabla\Phi_{i-1}(x_{i-1})\|$ $\leq 8(\sigma_{0}L_{f}+L_{g})Di\sqrt{\frac{i-1-q}{qi(q+1)}}\sqrt{\log\left(\frac{8i(i+1)}{\delta}\right)}$
$\leq 8(\sigma_{0}L_{f}+L_{g})\frac{Di}{q}\sqrt{\log\left(\frac{8i(i+1)}{\delta}\right)}.$
Hence, if $q<t\leq 2q$, with probability at least $1-\delta\sum_{q<i\leq t}1/i(i+1)$, we have
$\sum_{q<i\leq t}i\|\widehat{\nabla\Phi}_{i-1}-\nabla\Phi_{i-1}(x_{i-1})\|$ $\leq\sum_{q<i\leq t}8(\sigma_{0}L_{f}+L_{g})\frac{Di}{q}\sqrt{\log\left(\frac{16t^{2}}{\delta}\right)}$
$\leq(\sigma_{0}L_{f}+L_{g})D(8\sqrt{2}t-4q)\sqrt{\log\left(\frac{16t^{2}}{\delta}\right)},$
where the second inequality holds as $\frac{1}{q}\sum_{q<i\leq t}i\leq\sqrt{2}t-q/2$ for all $q<t\leq 2q$. Similarly, given $\delta\in(0,1)$ and $i>2q$, by Lemma 7 and the inequality Eq. 22, with probability at least $1-\delta/i(i+1)$, we have
$i\|\widehat{\nabla\Phi}_{i-1}-\nabla\Phi_{i-1}(x_{i-1})\|$ $\leq\frac{8(\sigma_{0}L_{f}+L_{g})Di}{\sqrt{i(q\lfloor(i-1)/q\rfloor+1)}}\sqrt{\log\left(\frac{8i(i+1)}{\delta}\right)}$
$\leq 8(\sigma_{0}L_{f}+L_{g})D\sqrt{\frac{i}{i-q}}\sqrt{\log\left(\frac{8i(i+1)}{\delta}\right)}$
$\leq 8\sqrt{2}(\sigma_{0}L_{f}+L_{g})D\sqrt{\log\left(\frac{8i(i+1)}{\delta}\right)},$
where the second inequality follows since $q\lfloor(i-1)/q\rfloor+1\geq i-q$, and the third inequality follows since $i/(i-q)<2$ when $i>2q$. Thus, if $t>2q$, with probability at least $1-\delta\sum_{2q<i\leq t}1/i(i+1)=1-\delta$, we have
$\sum_{2q<i\leq t}i\|\widehat{\nabla\Phi}_{i-1}-\nabla\Phi_{i-1}(x_{i-1})\|$ $\leq\sum_{2q<i\leq t}8\sqrt{2}(\sigma_{0}L_{f}+L_{g})D\sqrt{\log\left(\frac{16t^{2}}{\delta}\right)}$
$=8\sqrt{2}(\sigma_{0}L_{f}+L_{g})D(t-2q)\sqrt{\log\left(\frac{16t^{2}}{\delta}\right)}.$
For $t>2q$, with probability at least $1-\delta\sum_{q<i\leq t}1/i(i+1)$, the sum over all $q<i\leq t$ is thus
$\sum_{q<i\leq t}i\|\widehat{\nabla\Phi}_{i-1}-\nabla\Phi_{i-1}(x_{i-1})\|$
$\leq(\sigma_{0}L_{f}+L_{g})D(8\sqrt{2}(2q)-4q)\sqrt{\log\left(\frac{16t^{2}}{\delta}\right)}+8\sqrt{2}(\sigma_{0}L_{f}+L_{g})D(t-2q)\sqrt{\log\left(\frac{16t^{2}}{\delta}\right)}$
$=\left(8\sqrt{2}t-(16\sqrt{2}-16\sqrt{2}+4)q\right)(\sigma_{0}L_{f}+L_{g})D\sqrt{\log\left(\frac{16t^{2}}{\delta}\right)}$
$=\left(8\sqrt{2}t-4q\right)(\sigma_{0}L_{f}+L_{g})D\sqrt{\log\left(\frac{16t^{2}}{\delta}\right)}.$
Therefore, regardless of whether $q<t\leq 2q$ or $t>2q$, (21) holds. ∎
The following parameters will appear in our analysis and convergence rates
$C_{q}$ $:=\sup_{t\geq q}\left\{\begin{array}[]{l}\displaystyle\left(F_{\operatorname{opt}}-\min_{x\in X}F(x)\right)\left(1+\frac{\sum_{i\in[t]}(i+1)i(\sigma_{i-1}-\sigma_{i})}{(t+1)t\sigma_{t}}\right)\\[12.91663pt] \displaystyle+\frac{2(\sigma_{0}L_{f}+L_{g})D^{2}(t-q)}{(t+1)t\sigma_{t}}\end{array}\right\},$ (23c)
$\bar{C}_{\delta,q}$ $:=\sup_{t\geq q}\left\{\begin{array}[]{l}\frac{(q+1)\left(\max\left\{\Phi_{0}(x_{0})-\Phi_{0}(x_{\operatorname{opt}}),0\right\}+(L_{f}\sigma_{0}+L_{g})D^{2}\left(\frac{1}{2}+16\sqrt{\log\left(\frac{16t^{2}}{\delta}\right)}\right)\log(q)\right)}{(t+1)t\sigma_{t}}\\[8.61108pt] +\frac{\left(8\sqrt{2}t-4q\right)(\sigma_{0}L_{f}+L_{g})D^{2}\sqrt{\log\left(\frac{16t^{2}}{\delta}\right)}}{(t+1)t\sigma_{t}}\end{array}\right\},$ (23f)
$C_{\delta,q}$ $:=\bar{C}_{\delta,q}+C_{q},$ (23g)
$V_{q}$ $:=\sup_{t\geq q}\left\{\frac{\sum_{i\in[t]\setminus[q]}(i+1)i(\sigma_{i-1}-\sigma_{i})\sigma_{i}}{(t+1)t\sigma_{t}^{2}}\right\}.$ (23h)
Lemma 16 in Section B.4 guarantees that these quantities remain finite for any parameter choice satisfying 1 to 2 and 4. In particular, these conditions hold when $\sigma_{t}=\varsigma(t+1)^{p}$ for any $p\in(0,1)$, which is what is used in Theorem 2. We now analyze the sequence $\{x_{t}\}_{t\geq 0}$ in Algorithm 2 for the inner-level problem.
Lemma 9.
Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm 2 with $\alpha_{t}=\log(q)/q$ for $0\leq t<q$ $\alpha_{t}=2/(t+2)$ for any $t\geq q$, $S=q$, and let $C_{\delta,q}$ be defined as in (23g). If 1 and 1, 2 and 4 hold, then with probability at least $1-\delta$, it holds that
$G(x_{t})-G_{\operatorname{opt}}\leq C_{\delta,q}\sigma_{t},$ (24)
for any $t\geq q$.
Proof of Lemma 9.
Combining (11) from Proposition 1 with the probabilistic bounds in Lemma 8 and Proposition 3, and after some straightforward calculations, with probability at least $1-\delta$, we have
$G(x_{t})-G_{\operatorname{opt}}$
$\leq\frac{(q+1)\left(\max\left\{\Phi_{0}(x_{0})-\Phi_{0}(x_{\operatorname{opt}}),0\right\}+(L_{f}\sigma_{0}+L_{g})D^{2}\left(\frac{1}{2}+16\sqrt{\log\left(\frac{16t^{2}}{\delta}\right)}\right)\log(q)\right)}{(t+1)t}$
$\quad+\frac{2(\sigma_{0}L_{f}+L_{g})D^{2}(t-q)+\left(8\sqrt{2}t-4q\right)(\sigma_{0}L_{f}+L_{g})D^{2}\sqrt{\log\left(\frac{16t^{2}}{\delta}\right)}}{(t+1)t}$
$\quad+\left(F_{\operatorname{opt}}\!-\!\min_{x\in X}F(x)\right)\left(\sigma_{t}\!+\!\frac{\sum_{i\in[t]}(i+1)i(\sigma_{i-1}\!-\!\sigma_{i})}{(t+1)t}\right).$
Here, the right-hand side is at most $C_{\delta,q}$, and therefore, $G(x_{t})-G_{\operatorname{opt}}\leq C_{\delta,q}\sigma_{t}$ for any $t\geq q$. ∎
We next analyze the convergence of the sequence $\{z_{t}\}_{t\geq 0}$ in Algorithm 2 for both outer- and inner-level problems.
Lemma 10.
Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm 2 with $\alpha_{t}=\log(q)/q$ for $0\leq t<q$ $\alpha_{t}=2/(t+2)$ for any $t\geq q$, $S=q$, and let $C_{\delta,q}$ and $V_{q}$ defined as in Eq. 23. If 1 and 1, 2 and 4 hold, then with probability at least $1-\delta$, it (jointly) holds that
$F(z_{t})-F_{\operatorname{opt}}$
$\leq\frac{(q+1)\left(\max\left\{\Phi_{0}(x_{0})-\Phi_{0}(x_{\operatorname{opt}}),0\right\}+(L_{f}\sigma_{0}+L_{g})D^{2}\left(\frac{1}{2}+16\sqrt{\log\left(\frac{16t^{2}}{\delta}\right)}\right)\log(q)\right)}{(t+1)t\sigma_{t}}$
$\quad+\frac{2(\sigma_{0}L_{f}+L_{g})D^{2}(t-q)+\left(8\sqrt{2}t-4q\right)(\sigma_{0}L_{f}+L_{g})D^{2}\sqrt{\log\left(\frac{16t^{2}}{\delta}\right)}}{(t+1)t\sigma_{t}},$
$G(z_{t})-G_{\operatorname{opt}}\leq C_{\delta,q}(1+V_{q})\sigma_{t},$
for any $t\geq q$.
Proof of Lemma 10.
Recall from Algorithm 1 that we have defined
$\begin{cases}S_{t+1}:=(t+2)(t+1)\sigma_{t+1}+\sum_{i\in[t+1]\backslash[q]}(i+1)i(\sigma_{i-1}-\sigma_{i})\\[8.61108pt] z_{t+1}:=\frac{1}{S_{t+1}}\left((t+2)(t+1)\sigma_{t+1}x_{t+1}+\sum_{i\in[t+1]\backslash[q]}(i+1)i(\sigma_{i-1}-\sigma_{i})x_{i}\right),\end{cases}$
thus $z_{t}$ is simply a convex combination of $x_{0},\dots,x_{t}$ for every $t\geq q$. Therefore, as $F$ is convex, we can apply Jensen’s inequality to the left-hand side of the inequality (12) with $k=q$, and after some tedious calculation, for any $t\geq q$, we arrive at
$F(z_{t})-F_{\operatorname{opt}}$ $\leq\frac{2(\sigma_{0}L_{f}+L_{g})D^{2}t}{S_{t}}+\frac{(q+1)q\left(\Phi_{q}(x_{q})-\Phi_{q}(x_{\operatorname{opt}})\right)}{S_{t}}$
$\quad+\sum_{i\in[t]}\frac{4iD\|\widehat{\nabla\Phi}_{i-1}-\nabla\Phi_{i-1}(x_{i-1})\|}{S_{t}}.$
Using the inequality $S_{t}\geq(t+1)t\sigma_{t}$, which holds thanks to 1, and applying Lemma 8 and Proposition 2, with probability at least $1-\delta$, we have
$F(z_{t})-F_{\operatorname{opt}}$
$\leq\frac{(q+1)\left(\max\left\{\Phi_{0}(x_{0})-\Phi_{0}(x_{\operatorname{opt}}),0\right\}+(L_{f}\sigma_{0}+L_{g})D^{2}\left(\frac{1}{2}+16\sqrt{\log\left(\frac{16t^{2}}{\delta}\right)}\right)\log(q)\right)}{(t+1)t\sigma_{t}}$
$\quad+\frac{2(\sigma_{0}L_{f}+L_{g})D^{2}(t-q)+\left(8\sqrt{2}t-4q\right)(\sigma_{0}L_{f}+L_{g})D^{2}\sqrt{\log\left(\frac{16t^{2}}{\delta}\right)}}{(t+1)t\sigma_{t}}$
for every $t\geq q$. This completes the proof of the first claim.
For the second claim, we follow the same procedure. In particular, applying Jensen’s inequality with respect to the convex function $G$ to the left-hand side of (24) and using the inequality $S_{t}\geq(t+1)t\sigma_{t}$, with probability at least $1-\delta$, we have
$G(z_{t})-G_{\operatorname{opt}}\leq\frac{C_{\delta,q}}{t(t+1)\sigma_{t}}\left((t+1)t\sigma_{t}^{2}+\sum_{i\in[t]/[q]}(i+1)i(\sigma_{i-1}-\sigma_{i})\sigma_{i}\right).$
The proof concludes by using the definition of $V_{q}$. ∎
Lemma 11.
Consider the sequence $\sigma_{t}:=\varsigma(t+1)^{-p}$ for $t\geq 0$ and the quantities defined in Eq. 23. If $p\in(0,1)$, then the sequence $\{\sigma_{t}\}_{t\geq 0}$ satisfies 1, 2 and 4 with $L=p$. Furthermore,
$C_{q}\leq(1+2p)\left(F_{\operatorname{opt}}-\min_{x\in X}F(x)\right)+\frac{2(\varsigma L_{f}+L_{g})D^{2}}{\varsigma},\quad V_{q}\leq\frac{2p}{\min\{1,2(1-p)\}}$
and
$\bar{C}_{\delta,q}$ $\leq\max_{x\in X}F(x)-f(x_{0})+\frac{\max_{x\in X}g(x)-g(x_{0})}{\varsigma}$
$\quad+\frac{2}{1-p}\left(L_{f}+\frac{L_{g}}{\varsigma}\right)D^{2}\left(\frac{1}{2}+16\sqrt{\frac{2}{e(1-p)}+\log\left(\frac{16}{\delta}\right)}\right)$
$\quad+8\sqrt{2}\left(L_{f}+\frac{L_{g}}{\varsigma}\right)D^{2}\sqrt{\frac{1}{e(1-p)}+\log\left(\frac{16}{\delta}\right)}.$
Proof of Lemma 11.
We observe that $C_{q},V_{q}$ can be bounded from above by $C,V$ defined in (15) under the conditions in Lemma 6. Therefore, we focus on $\bar{C}_{\delta,q}$ for the remainder of the proof. Using the basic inequality $\log(x)/x\leq 1/e$ for any $x>0$, observe that for any $t\geq q$
$\frac{(q+1)\left(\max\left\{\Phi_{0}(x_{0})-\Phi_{0}(x_{\operatorname{opt}}),0\right\}+(L_{f}\sigma_{0}+L_{g})D^{2}\left(\frac{1}{2}+16\sqrt{\log\left(\frac{16t^{2}}{\delta}\right)}\right)\log(q)\right)}{(t+1)t\sigma_{t}}$
$\leq\frac{\max\left\{\Phi_{0}(x_{0})-\Phi_{0}(x_{\operatorname{opt}}),0\right\}+(L_{f}\varsigma+L_{g})D^{2}\left(\frac{1}{2}+16\sqrt{\log\left(\frac{16t^{2}}{\delta}\right)}\right)\log(t)}{\varsigma t^{1-p}}.$
We have
$\Phi_{0}(x_{0})-\Phi_{0}(x_{\operatorname{opt}})=\sigma_{0}\left(f(x_{0})-F_{\operatorname{opt}}\right)+g(x_{0})-G_{\operatorname{opt}}\leq\varsigma\left(\max_{x\in X}F(x)-F_{\operatorname{opt}}\right)+\max_{x\in X}g(x)-G_{\operatorname{opt}},$
which implies
$\frac{\max\left\{\Phi_{0}(x_{0})-\Phi_{0}(x_{\operatorname{opt}}),0\right\}}{\varsigma t^{1-p}}\leq\varsigma\left(\max_{x\in X}F(x)-F_{\operatorname{opt}}\right)+\max_{x\in X}g(x)-G_{\operatorname{opt}}.$
We also have
$\frac{\log(t)}{t^{(1-p)/2}}=\frac{2}{1-p}\frac{\log(t^{(1-p)/2})}{t^{(1-p)/2}}\leq\frac{2}{e(1-p)},$
and
$\frac{1}{t^{1-p}}\log\left(\frac{16t^{2}}{\delta}\right)=\frac{1}{t^{1-p}}\log\left(\frac{16}{\delta}\right)+\frac{2}{1-p}\frac{\log(t^{1-p})}{t^{1-p}}\leq\log\left(\frac{16}{\delta}\right)+\frac{2}{e(1-p)}.$
Therefore,
$\frac{(L_{f}\varsigma+L_{g})D^{2}\left(\frac{1}{2}+16\sqrt{\log\left(\frac{16t^{2}}{\delta}\right)}\right)\log(t)}{\varsigma t^{1-p}}$
$\leq\frac{2}{1-p}\left(L_{f}+\frac{L_{g}}{\varsigma}\right)D^{2}\left(\frac{1}{2}+16\sqrt{\frac{2}{e(1-p)}+\log\left(\frac{16}{\delta}\right)}\right)$
Similarly, we have
$\frac{\left(8\sqrt{2}t-4q\right)(\sigma_{0}L_{f}+L_{g})D^{2}\sqrt{\log\left(\frac{16t^{2}}{\delta}\right)}}{(t+1)t\sigma_{t}}\leq 8\sqrt{2}\left(L_{f}+\frac{L_{g}}{\varsigma}\right)D^{2}\sqrt{\frac{1}{e(1-p)}+\log\left(\frac{16}{\delta}\right)}$
This concludes the proof. ∎
Proof of Theorem 2.
This is an immediate consequence of Lemmas 10 and 11. ∎
Appendix B Additional Results
B.1 Asymptotic convergence
We show that, even when $\sigma_{t}$ is not chosen as $\varsigma(t+1)^{-p}$ as in Theorems 1 and 2, we can still get almost sure convergence of our iterates for both Algorithms 1 and 2 (though without rates), as long as the regularization sequence satisfies the conditions in Lemma 4 or Lemma 9 respectively.
Theorem 3.
Let $\{x_{t}\}_{t\geq 0}$ be the iterates generated by either Algorithm 1 under the same conditions as Lemma 4. Then with probability $1$, we have
$\lim_{t\to\infty}F(x_{t})=F_{\operatorname{opt}},\quad\lim_{t\to\infty}G(x_{t})=G_{\operatorname{opt}}.$
Proof of Theorem 3.
Lemma 4 implies that $\lim_{t\to\infty}G(x_{t})=G_{\operatorname{opt}}$ with probability at least $1-\delta$. As $\delta$ can be arbitrarily small, we may conclude that $\lim_{t\to\infty}G(x_{t})=G_{\operatorname{opt}}$, almost surely. This implies that any limit point of $\{x_{t}\}_{t\geq 0}$ is in $X_{\operatorname{opt}}$ almost surely. Since $F$ is convex, hence lower semi-continuous over $X$, and by definition of $F_{\operatorname{opt}}$, we have $\liminf_{t\to\infty}F(x_{t})\geq F_{\operatorname{opt}}$ almost surely.
Besides, by combining Propositions 1 and 2, we deduce that for any $\delta\in(0,1)$, with probability at least $1-\delta$, we have, for all $t\geq 1$,
$F(x_{t})-F_{\operatorname{opt}}$ $\leq\frac{4c\left(4(\sigma_{0}L_{f}+L_{g})D^{2}+3(\sigma_{0}\sigma_{f}+\sigma_{g})D\right)}{(t+1)^{1/2}\sigma_{t}}\sqrt{\log\left(\frac{8dt^{2}}{\delta}\right)}$
$\quad+\frac{2(\sigma_{0}L_{f}+L_{g})D^{2}}{(t+1)\sigma_{t}}+\frac{\sum_{i\in[t]}(i+1)i(\sigma_{i-1}-\sigma_{i})(F_{\operatorname{opt}}-F(x_{i}))}{t(t+1)\sigma_{t}}$
$\leq\frac{4c\left(4(\sigma_{0}L_{f}+L_{g})D^{2}+3(\sigma_{0}\sigma_{f}+\sigma_{g})D\right)}{(t+1)^{1/2}\sigma_{t}}\sqrt{\log\left(\frac{8dt^{2}}{\delta}\right)}$
$\quad+\frac{2(\sigma_{0}L_{f}+L_{g})D^{2}}{(t+1)\sigma_{t}}+\frac{\sum_{i\in[t]}(i+1)i(\sigma_{i-1}-\sigma_{i})\max\{F_{\operatorname{opt}}-F(x_{i}),0\}}{t(t+1)\sigma_{t}},$
where the second inequality holds due to 1. We claim that all terms on the right hand side converge to 00 as $t\to\infty$. This is obvious except for the last term.
Recall that almost surely $\liminf_{t\to\infty}F(x_{t})\geq F_{\operatorname{opt}}$. Therefore we have $\limsup_{t\to\infty}(F_{\operatorname{opt}}-F(x_{t}))\leq 0$, hence $\lim_{t\to\infty}\max\{F_{\operatorname{opt}}-F(x_{t}),0\}=0$ almost surely. Applying the Stolz–Cesàro theorem and (33) leads to
$\lim_{t\to\infty}\frac{\sum_{i\in[t]}(i+1)i(\sigma_{i-1}-\sigma_{i})\max\{F_{\operatorname{opt}}-F(x_{i}),0\}}{t(t+1)\sigma_{t}}=0,$
almost surely. Then each term on the right hand side of the inequality converges to 00 as $t\to\infty$. It follows that with probability at least $1-\delta$,
$\limsup_{t\to\infty}(F(x_{t})-F_{\operatorname{opt}})$ $\leq 0.$
Thus, since this holds for any $\delta\in(0,1)$, it holds almost surely. In conclusion, we have shown that $\limsup_{t\to\infty}(F(x_{t})-F_{\operatorname{opt}})\leq 0$ and $\liminf_{t\to\infty}(F(x_{t})-F_{\operatorname{opt}})\geq 0$, which implies that $\lim_{t\to\infty}F(x_{t})=F_{\operatorname{opt}}$ almost surely. ∎
Theorem 4.
Let $\{x_{t}\}_{t\geq 0}$ be the iterates generated by Algorithm 2 under the same conditions as Lemma 9. Then with probability $1$, we have
$\lim_{t\to\infty}F(x_{t})=F_{\operatorname{opt}},\quad\lim_{t\to\infty}G(x_{t})=G_{\operatorname{opt}}.$
Proof of Theorem 4.
Lemma 9 implies that $\lim_{t\to\infty}G(x_{t})=G_{\operatorname{opt}}$ with probability at least $1-\delta$. As $\delta$ can be arbitrarily small, we conclude that $\lim_{t\to\infty}G(x_{t})=G_{\operatorname{opt}}$, almost surely. This implies that any limit point of $\{x_{t}\}_{t\geq 0}$ is in $X_{\operatorname{opt}}$ almost surely. Since $F$ is convex, hence lower semi-continuous over $X$, and by definition of $F_{\operatorname{opt}}$, we have $\liminf_{t\to\infty}F(x_{t})\geq F_{\operatorname{opt}}$ almost surely.
Besides, by combining Proposition 1, Lemma 8 and Proposition 3, with probability at least $1-\delta$, we have
$F(x_{t})-F_{\operatorname{opt}}$
$\leq\frac{(q+1)\left(\max\left\{\Phi_{0}(x_{0})-\Phi_{0}(x_{\operatorname{opt}}),0\right\}+(L_{f}\sigma_{0}+L_{g})D^{2}\left(\frac{1}{2}+16\sqrt{\log\left(\frac{16t^{2}}{\delta}\right)}\right)\log(q)\right)}{(t+1)t\sigma_{t}}$
$\quad+\frac{2(\sigma_{0}L_{f}+L_{g})D^{2}(t-q)+\left(8\sqrt{2}t-4q\right)(\sigma_{0}L_{f}+L_{g})D^{2}\sqrt{\log\left(\frac{16t^{2}}{\delta}\right)}}{(t+1)t\sigma_{t}}$
$\quad+\frac{\sum_{i\in[t]}(i+1)i(\sigma_{i-1}-\sigma_{i})\max\{F_{\operatorname{opt}}-F(x_{i}),0\}}{t(t+1)\sigma_{t}}.$
Similar to the proof of Theorem 3, all terms on the right-hand side converge to 00 as $t\to\infty$, implying that $\limsup_{t\to\infty}(F(x_{t})-F_{\operatorname{opt}})\leq 0$. Combined with $\liminf_{t\to\infty}(F(x_{t})-F_{\operatorname{opt}})\geq 0$, this yields $\lim_{t\to\infty}F(x_{t})=F_{\operatorname{opt}}$ almost surely, concluding the proof. ∎
B.2 Nonconvex setting
Define the auxiliary functions
$\mathcal{F}(x):=\max_{v\in X_{\operatorname{opt}}}\nabla F(x)^{\top}(x-v)\quad\&\quad\mathcal{G}(x):=G(x)-G_{\operatorname{opt}}$
for all $x\in X$. Recall that if $x\in X_{\operatorname{opt}}$ then $\mathcal{F}(x)=0$ is the necessary and sufficient optimality condition in convex case. Thus, $\mathcal{F}$ serves as a generalization of the stationarity condition $\nabla F(x)=0$ in the unconstrained setting. We start by establishing a bound, akin to Lemma 2, that holds even if $F$ is nonconvex.
Lemma 12.
Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm 1 or Algorithm 2 with $\{\alpha_{t},\sigma_{t}\}_{t\geq 0}$ given by $\alpha_{t}=(t+1)^{-\omega}$ and $\sigma_{t}=\varsigma(t+1)^{-p}$ for $t\geq k$, with $0<p\leq\omega$. If 1 holds with the exception that $F$ may be nonconvex, then for every $t\geq k$, we have
$(t+1)^{\omega}\mathcal{G}(x_{t+1})-k^{\omega}\mathcal{G}(x_{k})$ $\leq 2\varsigma\sup_{z\in X}|F(z)|(t+1)^{\omega-p}+\sigma_{k}F(x_{k})-\sum_{i=k}^{t}\sigma_{i}\mathcal{F}(x_{i})$
$\quad+\sum_{i=k}^{t}\left(2D\left\|\widehat{\nabla\Phi}_{i}-\nabla\Phi_{i}(x_{i})\right\|+\frac{(\sigma_{i}L_{f}+L_{g})D^{2}}{2(i+1)^{\omega}}\right).$
Proof of Lemma 12.
Using the definition of $\Phi_{t}=\sigma_{t}F+G$, for any $t\geq k$, we have
$\sigma_{t}F(x_{t+1})+G(x_{t+1})$ $\leq\sigma_{t}F(x_{t})+G(x_{t})+\alpha_{t}\min_{v\in X}(\sigma_{t}\nabla F(x_{t})+\nabla G(x_{t}))^{\top}(v-x_{t})$
$\quad+2D\alpha_{t}\|\widehat{\nabla\Phi}_{t}-\nabla\Phi_{t}(x_{t})\|+\frac{(\sigma_{t}L_{f}+L_{g})D^{2}\alpha_{t}^{2}}{2}$
where the inequality follows directly from (9) and by minimizing over $w_{t}\in X$. We emphasize that (9) does not involve any convexity assumptions on $\Phi_{t}$ and relies solely on its smoothness. Furthermore, notice that
$\min_{v\in X}~{}(\sigma\nabla F(x)+\nabla G(x))^{\top}(v-x)$ $\leq\min_{v\in X_{\operatorname{opt}}}~{}(\sigma\nabla F(x)+\nabla G(x))^{\top}(v-x)$
$\leq\min_{v\in X_{\operatorname{opt}}}~{}\left\{\sigma\nabla F(x)^{\top}(v-x)+G(x)-G(v)\right\}$
$=-\sigma\mathcal{F}(x)-\mathcal{G}(x),$
where the last equality follows from the definition of $\mathcal{F}$ and $\mathcal{G}$. Using the above bound, we may thus conclude that
$\sigma_{t}F(x_{t+1})+G(x_{t+1})$ $\leq\sigma_{t}F(x_{t})+G(x_{t})-\alpha_{t}(\sigma_{t}\mathcal{F}(x_{t})+\mathcal{G}(x_{t}))$
$\quad+2D\alpha_{t}\|\widehat{\nabla\Phi}_{t}-\nabla\Phi_{t}(x_{t})\|+\frac{(\sigma_{t}L_{f}+L_{g})D^{2}\alpha_{t}^{2}}{2}.$
A simple re-arrangement and using the definition of $\mathcal{G}$ then yields the bound
[…]
Using the definition of $\alpha_{t}$, we arrive at
$(t+1)^{\omega}\mathcal{G}(x_{t+1})$ $\leq t^{\omega}\mathcal{G}(x_{t})+(t+1)^{\omega}\sigma_{t}(F(x_{t})-F(x_{t+1}))-\sigma_{t}\mathcal{F}(x_{t})$
$\quad+2D\|\widehat{\nabla\Phi}_{t}-\nabla\Phi_{t}(x_{t})\|+\frac{(\sigma_{t}L_{f}+L_{g})D^{2}}{2(t+1)^{\omega}}.$
Unwinding this recursion back to $i=k$, we obtain
$(t+1)^{\omega}\mathcal{G}(x_{t+1})$ $\leq k^{\omega}\mathcal{G}(x_{k})+\sum_{i=k}^{t}\left((i+1)^{\omega}\sigma_{i}(F(x_{i})-F(x_{i+1}))\right)-\sum_{i=k}^{t}\sigma_{i}\mathcal{F}(x_{i})$
$\quad+\sum_{i=k}^{t}\left(2D\|\widehat{\nabla\Phi}_{i}-\nabla\Phi_{i}(x_{i})\|+\frac{(\sigma_{i}L_{f}+L_{g})D^{2}}{2(i+1)^{\omega}}\right).$
Since $p\leq\omega$, we have $(i+1)^{\omega}\sigma_{i}-i^{\omega}\sigma_{i-1}\geq 0$, therefore
$\sum_{i=k}^{t}(i+1)^{\omega}\sigma_{i}(F(x_{i})-F(x_{i+1}))$
$=\sum_{i=k}^{t}\Big{(}F(x_{i})((i+1)^{\omega}\sigma_{i}-i^{\omega}\sigma_{i-1})\Big{)}+\sigma_{k}F(x_{k})-(t+1)^{\omega}\sigma_{t}F(x_{t+1})$
$\leq\sup_{z\in X}|F(z)|\sum_{i=k}^{t}\Big{(}((i+1)^{\omega}\sigma_{i}-i^{\omega}\sigma_{i-1})\Big{)}+\sigma_{k}F(x_{k})+(t+1)^{\omega}\sigma_{t}\sup_{z\in X}|F(z)|$
$=2\varsigma\sup_{z\in X}|F(z)|(t+1)^{\omega-p}+\sigma_{k}F(x_{k}).$
This completes the proof. ∎
B.2.1 One-sample stochastic setting
In this section we establish the convergence of Algorithm 1 in the nonconvex setting.
Theorem 5.
Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm 1 with stepsizes $\alpha_{t}=(t\!+\!1)^{-\omega}$ and regularization parameters $\sigma_{t}=\varsigma(t\!+\!1)^{-p}$ for any $t\!\geq\!0$, with $p\!=\!\frac{2}{7}$ and $\omega\!=\!\frac{6}{7}$. If Assumption 1 holds with the exception that $F$ may be nonconvex, then with probability at least $1-\delta$, for every $t\geq 0$, it (jointly) holds that
$\frac{\sum_{i=0}^{t}\sigma_{i}\mathcal{F}(x_{i})}{\sum_{i=0}^{t}\sigma_{i}}\leq O\left(t^{-\frac{1}{7}}\log\left(\frac{dt^{2}}{\delta}\right)\right)\quad\&\quad\mathcal{G}(x_{t})\leq O\left(t^{-\frac{1}{7}}\log\left(\frac{dt^{2}}{\delta}\right)\right).$
Proof of Theorem 5.
Following the proof of [4, Lemma 4.1], for any $i\geq 0$, one can easily show that, with probability at least $1-\eta_{i}$, we have
$\|\widehat{\nabla\Phi}_{i}-\nabla\Phi_{i}(x_{i})\|\leq c\sqrt{2}\left(2(\sigma_{i}L_{f}+L_{g})D+\frac{3^{\omega}}{3^{\omega}-1}(\sigma_{i}\sigma_{f}+\sigma_{g})\right)(i+1)^{-\omega/2}\sqrt{\log\left(\frac{4d}{\eta_{i}}\right)}.$
Setting $\eta_{i}=\frac{\delta}{(i+1)(i+2)}$, applying a union bound, and using the fact that $\sigma_{i}\leq\sigma_{0}=\varsigma$ we have that with probability at least $1-\delta\sum_{i=0}^{t}\frac{1}{(i+1)(i+2)}\geq 1-\delta$, the following holds:
$\sum_{i=0}^{t}\left(2D\|\widehat{\nabla\Phi}_{i}-\nabla\Phi_{i}(x_{i})\|+\frac{(\sigma_{i}L_{f}+L_{g})D^{2}}{2(i+1)^{\omega}}\right)$
$\leq\sum_{i=0}^{t}\left(\frac{2\sqrt{2}c(2(\varsigma L_{f}+L_{g})D+\frac{3^{\omega}}{3^{\omega}-1}(\sigma_{i}\sigma_{f}+\sigma_{g}))D}{(i+1)^{\omega/2}}\sqrt{\log\left(\frac{4d(t+2)^{2}}{\delta}\right)}+\frac{(\varsigma L_{f}+L_{g})}{2(i+1)^{\omega}}D^{2}\right)$
$\leq\frac{2\sqrt{2}c(2(\varsigma L_{f}+L_{g})D+\frac{3^{\omega}}{3^{\omega}-1}(\varsigma\sigma_{f}+\sigma_{g}))D}{1-\omega/2}(t+1)^{1-\omega/2}\sqrt{\log\left(\frac{4d(t+2)^{2}}{\delta}\right)}$
$\quad+\frac{(\varsigma L_{f}+L_{g})D^{2}}{1-\omega}(t+1)^{1-\omega}.$
Applying Lemma 12 with $k=0$, with probability at least $1-\delta$, we thus have
[…]
Note that the function $\mathcal{F}$ is defined with respect to $X_{\operatorname{opt}}$, not $X$. Therefore, $\mathcal{F}(x_{i})$ can take both positive and negative values since $x_{i}\in X$. However, we can derive the following bound
$-\sum_{i=0}^{t}\sigma_{i}\mathcal{F}(x_{i})\leq\sup_{z\in X}|\mathcal{F}(z)|\sum_{i=0}^{t}\sigma_{i}\leq\varsigma\sup_{z\in X}|\mathcal{F}(z)|\frac{(t+1)^{1-p}}{1-p},$
which further implies that
$\mathcal{G}(x_{t})$ $\leq\varsigma\sup_{z\in X}|F(z)|\left(2t^{-p}+\frac{t^{1-p-w}}{1-p}\right)+\varsigma F(x_{0})t^{-\omega}+\frac{(\varsigma L_{f}+L_{g})D^{2}}{1-\omega}t^{1-2\omega}$
$\quad+\frac{2c\sqrt{2}(2(\varsigma L_{f}+L_{g})D+\frac{3^{\omega}}{3^{\omega}-1}(\varsigma\sigma_{f}+\sigma_{g}))D}{1-\omega/2}t^{1-3\omega/2}\sqrt{\log\left(\frac{4d(t+q)^{2}}{\delta}\right)}.$
We next focus on the upper-level problem. Since $\mathcal{G}(x_{t+1})\geq 0$, using Eq. 26, we have
$\sum_{i=0}^{t}\sigma_{i}\mathcal{F}(x_{i})$ $\leq 2\varsigma\sup_{z\in X}|F(z)|(t+1)^{\omega-p}+\varsigma F(x_{0})+\frac{(\varsigma L_{f}+L_{g})D^{2}}{1-\omega}(t+1)^{1-\omega}$
$+\frac{2c\sqrt{2}(2(\varsigma L_{f}+L_{g})D+\frac{3^{\omega}}{3^{\omega}-1}(\varsigma\sigma_{f}+\sigma_{g}))D}{1-\omega/2}(t+1)^{1-\omega/2}\sqrt{\log\left(\frac{4d(t+2)^{2}}{\delta}\right)}.$
Dividing both sides by $\sum_{i=0}^{t}\sigma_{i}$, and furthermore exploiting the bounds
$\sum_{i=0}^{t}\sigma_{i}\geq\varsigma\frac{(t+2)^{1-p}-1}{1-p}\geq\varsigma\frac{(t+1)^{1-p}}{2(1-p)},$
we arrive at
$\frac{\sum_{i=0}^{t}\sigma_{i}\mathcal{F}(x_{i})}{\sum_{i=0}^{t}\sigma_{i}}$ $\leq 4(1-p)\sup_{z\in X}|F(z)|(t+1)^{\omega-1}+2(1-p)F(x_{0})(t+1)^{p-1}$
$\quad+\frac{2(1-p)(\varsigma L_{f}+L_{g})D^{2}}{\varsigma(1-\omega)}(t+1)^{p-\omega}$
$+\tfrac{4c(1-p)\sqrt{2}(2(\varsigma L_{f}+L_{g})D+\frac{3^{\omega}}{3^{\omega}-1}(\varsigma\sigma_{f}+\sigma_{g}))D}{\varsigma(1-\omega/2)}(t+1)^{p-\omega/2}\sqrt{\log\left(\frac{4d(t+2)^{2}}{\delta}\right)}.$
Both bounds involving $\mathcal{G}$ and $\mathcal{F}$ hold with probability $\geq 1-\delta$. We want the right hand side of both bounds $\to 0$. In order to guarantee this, we choose $p,\omega$ to minimize the slowest rate in terms of $t$:
$\min_{p,w:0<p\leq\omega}\max\left\{-p,1-p-\omega,-\omega,1-2\omega,1-3\omega/2,\omega-1,p-1,p-\omega,p-\omega/2\right\}=-\frac{1}{7}$
which is realized by setting $p=2/7$, $\omega=6/7$ as required. ∎
B.2.2 Finite-sum setting
In this section we establish the convergence of Algorithm 2 in the nonconvex setting. To that end, we introduce the sequence $\{\beta_{t}\}_{t\geq 0}$, defined as
$\beta_{t}=\begin{cases}\displaystyle\frac{\sigma_{0}(q+1)^{\omega}\log(q+1)}{q+1}\left(1-\frac{\log(q+1)}{q+1}\right)^{q-t}&0\leq t\leq q,\\ \sigma_{t}&t>q.\end{cases}$ (27)
Theorem 6.
Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm 2 with stepsizes $\alpha_{t}=\log(q+1)/(q+1)$ for any $0\leq t\leq q$, $\alpha_{t}=1/(t+1)^{\omega}$ for any $t>q$ and $S=q$, and regularization parameters $\sigma_{t}=\varsigma(\max\{t,q+1\}+1)^{-p}$ with $p=\frac{1}{2}$ and $w=\frac{3}{4}$. If 1 holds with the exception that $F$ may be nonconvex and the sequence $\{\beta_{t}\}_{t\geq 0}$ is defined as in (27), then with probability at least $1-\delta$, for every $t\geq q$, it (jointly) holds that
$\frac{\sum_{i=0}^{t}\beta_{i}\mathcal{F}(x_{i})}{\sum_{i=0}^{t}\beta_{i}}\leq O\left(t^{-\frac{1}{4}}\log\left(\frac{dt^{2}}{\delta}\right)\right)\quad\&\quad\mathcal{G}(x_{t})\leq O\left(t^{-\frac{1}{4}}\log\left(\frac{dt^{2}}{\delta}\right)\right).$
The proof relies on several intermediate results. We begin with the following probabilistic bound.
Lemma 13.
Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm 2 with $\alpha_{t}\!=\!\alpha$ for any $0\leq t\leq q$ and $\alpha_{t}=(t+1)^{-\omega}$ with $\omega>\frac{1}{2}$ for any $t>q$. If 1 holds with the exception that $F$ may be nonconvex, then for any $t\geq 0$, given $\delta\in(0,1)$, with probability at least $1-\delta$, it (jointly) holds that
$\left\|\widehat{\nabla F}_{t}-\nabla F(x_{t})\right\|$ $\leq 8L_{f}D\sqrt{\frac{\alpha^{2}q}{S}}\sqrt{\log\left(\frac{8}{\delta}\right)}$ (28)
$\left\|\widehat{\nabla G}_{t}-\nabla G(x_{t})\right\|$ $\leq 8L_{g}D\sqrt{\frac{\alpha^{2}q}{S}}\sqrt{\log\left(\frac{8}{\delta}\right)},$
provided that $0\leq t<q$, and
$\left\|\widehat{\nabla F}_{t}-\nabla F(x_{t})\right\|$ $\leq 8L_{f}D\sqrt{\frac{(t^{1-2w}-s_{t}^{1-2w})}{S(1-2w)}}\sqrt{\log\left(\frac{8}{\delta}\right)}$ (29)
$\left\|\widehat{\nabla G}_{t}-\nabla G(x_{t})\right\|$ $\leq 8L_{g}D\sqrt{\frac{(t^{1-2w}-s_{t}^{1-2w})}{S(1-2w)}}\sqrt{\log\left(\frac{8}{\delta}\right)},$
where $s_{t}:=q\lfloor t/q\rfloor$, provided that $t>q$.
Proof of Lemma 13.
Recall from the proof of Lemma 7 that, employing the concentration inequality in [41, Theorem 3.5] yields
$P\left(\left\|\widehat{\nabla F}_{t}-\nabla F(x_{t})\right\|\geq\lambda\right)$ $\leq 4\exp\left(-\frac{\lambda^{2}}{4\sum_{s=s_{t}+1}^{t}\frac{4L_{f}^{2}D^{2}}{S}\alpha_{s-1}^{2}}\right).$
If $1\leq t\leq q$, one can easily show that
$P\left(\left\|\widehat{\nabla F}_{t}-\nabla F(x_{t})\right\|\geq\lambda\right)\leq 4\exp\left(-\frac{\lambda^{2}S}{16L_{f}^{2}D^{2}q\alpha^{2}}\right).$
If $t>q$, we observe that
$\sum_{s=s_{t}+1}^{t}\alpha_{s-1}^{2}=4\sum_{s=s_{t}+1}^{t}\frac{1}{s^{2w}}\leq 4\int_{s_{t}}^{t}\frac{du}{s^{2w}}=\frac{4}{2w-1}(s_{t}^{1-2w}-t^{1-2w}).$
Thus, we have
$P\left(\left\|\widehat{\nabla F}_{t}-\nabla F(x_{t})\right\|\geq\lambda\right)\leq 4\exp\left(-\frac{\lambda^{2}S(1-2w)}{64L_{f}^{2}D^{2}(t^{1-2w}-s_{t}^{1-2w})}\right).$
Arguing similarly for $\widehat{\nabla G}_{t}$, we have
$P\left(\left\|\widehat{\nabla G}_{t}-\nabla G(x_{t})\right\|\geq\lambda\right)\leq 4\exp\left(-\frac{\lambda^{2}S}{16L_{g}^{2}D^{2}q\alpha^{2}}\right),$
if $0\leq t\leq q$ and
$P\left(\left\|\widehat{\nabla G}_{t}-\nabla G(x_{t})\right\|\geq\lambda\right)\leq 4\exp\left(-\frac{\lambda^{2}S(1-2w)}{64L_{g}^{2}D^{2}(t^{1-2w}-s_{t}^{1-2w})}\right),$
if $t>q$. Given $\delta\in(0,1)$ and $1\leq t\leq q$, setting the right hand side $=\delta/2$ and solving for $\lambda$ yields
$P\left(\left\|\widehat{\nabla F}_{t}-\nabla F(x_{t})\right\|\geq 8L_{f}D\sqrt{\frac{\alpha^{2}q}{S}}\sqrt{\log\left(\frac{8}{\delta}\right)}\right)\leq\frac{\delta}{2}$
and
$P\left(\left\|\widehat{\nabla G}_{t}-\nabla G(x_{t})\right\|\geq 8L_{g}D\sqrt{\frac{\alpha^{2}q}{S}}\sqrt{\log\left(\frac{8}{\delta}\right)}\right)\leq\frac{\delta}{2}.$
Applying union bound, we deduce that with probability at least $1-\delta$, (17) holds. Arguing similarly for $t>q$, we also deduce that (18) holds with probability at least $1-\delta$. ∎
We next analyze the sequence $\{x_{t}\}_{t\geq 0}$ generated by Algorithm 2 over the first $q$ iterations.
Lemma 14.
Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm 2 with $\alpha_{t}=\alpha$ for any $0\leq t\leq q$ and $S=q$. If 1 holds with the exception that $F$ may be nonconvex and 4 is satisfied, then given $\delta\in(0,1)$, with probability at least $1-\delta\sum_{0\leq i\leq q}1/(i+1)(i+2)$, for any $10\leq t\leq q+$ 1, it holds that
$\mathcal{G}(x_{t})$ $\leq(1-\alpha)^{t}\mathcal{G}(x_{0})+2\sigma_{0}\sup_{z\in X}|F(z)|-\sigma_{0}\sum_{i=0}^{t-1}(1-\alpha)^{t-i}\alpha\mathcal{F}(x_{i})$
$\quad+\frac{(\sigma_{0}L_{f}+L_{g})D^{2}\alpha}{2}\left(1+16\sqrt{\log\left(\frac{8(t+2)^{2}}{\delta}\right)}\right).$
Proof of Lemma 14.
Applying the bound (25) and after a straightforward re-arrangement, under 4, for any $0\leq i\leq t$, we obtain
$\mathcal{G}(x_{i+1})$ $\leq(1-\alpha)\mathcal{G}(x_{i})+\sigma_{0}(F(x_{i})-\alpha\mathcal{F}(x_{i})-F(x_{i+1}))$
$\quad+2D\alpha\|\nabla\Phi_{0}(x_{i})-\widehat{\nabla\Phi}_{i}\|+\frac{(L_{f}\sigma_{0}+L_{g})D^{2}\alpha^{2}}{2}.$
Given $\delta\in(0,1)$, we apply Lemma 13 to deduce that with probability at least $1-\delta/(i+1)(i+2)$, we have
$\mathcal{G}(x_{i+1})$ $\leq(1-\alpha)\mathcal{G}(x_{i})+\sigma_{0}(F(x_{i})-\alpha\mathcal{F}(x_{i})-F(x_{i+1}))$
$\quad+8(\sigma_{0}L_{f}+L_{g})D^{2}\alpha^{2}\sqrt{\log\left(\frac{8(i+1)(i+2)}{\delta}\right)}+\frac{(L_{f}\sigma_{0}+L_{g})D^{2}\alpha^{2}}{2}.$
Applying the union bound, for any $0\leq t\leq q$, with probability at least $1-\delta\sum_{0\leq i\leq t}1/(i+1)(i+2)$, we obtain
$\mathcal{G}(x_{t+1})$ $\!\leq\!(1-\alpha)^{t+1}\mathcal{G}(x_{0})+\sum_{i=0}^{t}\sigma_{0}(F(x_{i})-\alpha\mathcal{F}(x_{i})-F(x_{i+1}))(1-\alpha)^{t-i}$
$\!+\!8(\sigma_{0}L_{f}\!+\!L_{g})D^{2}\alpha^{2}\sum_{i=0}^{t}(1-\alpha)^{t-i}\sqrt{\log\left(\tfrac{8(i+1)(i+2)}{\delta}\right)}\!+\!\tfrac{(L_{f}\sigma_{0}\!+\!L_{g})D^{2}\alpha^{2}}{2}\sum_{i=0}^{t}(1-\alpha)^{t-i}.$
Using the geometric series bound $\sum_{i=0}^{t}(1-\alpha)^{t-i}\leq 1/\alpha$ and observing that
$\sum_{i=0}^{t}(F(x_{i})-F(x_{i+1}))(1-\alpha)^{t-i}$
$=$ $\sum_{i\in[t]}F(x_{i})\left((1-\alpha)^{t-i}-(1-\alpha)^{t+1-i}\right)+F(x_{0})(1-\alpha)^{t}-F(x_{t+1})$
$=$ $\sum_{i\in[t]}\alpha F(x_{i})(1-\alpha)^{t-i}+F(x_{0})(1-\alpha)^{t}-F(x_{t+1})$
$\leq$ $\sup_{z\in X}|F(z)|\textstyle\left(\sum_{i\in[t]}\alpha(1-\alpha)^{t-i}+(1-\alpha)^{t}+1\right)\leq 2\sup_{z\in X}|F(z)|$
conclude the proof. ∎
We finally analyze the sequence $\{x_{t}\}_{t\geq 0}$ generated by Algorithm 2 after the first $q$ iterations.
Proposition 4.
Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm 2 with $\alpha_{t}=\alpha$ for any $0\leq t<q$, $\alpha_{t}=(t+1)^{-w}$ with $\omega>\frac{1}{2}$ for any $t\geq q$ and $S=q$. If 1 holds with the exception that $F$ may be nonconvex and the parameters $\{\sigma_{t}\}_{t\geq 0}$ are non-increasing and positive, then given $\delta\in(0,1)$, with probability at least $1-\delta\sum_{i\in[t]\setminus[q]}\frac{1}{(i+1)(i+2)}$, for any $t>q$, it holds that
$\sum_{i\in[t]\setminus[q]}\left\|\widehat{\nabla\Phi}_{i}-\nabla\Phi_{i}(x_{i})\right\|\leq\frac{8(\sigma_{0}L_{f}+L_{g})D(t-q)^{1-w}}{1-w}\sqrt{\log\left(\frac{8(t+2)^{2}}{\delta}\right)}.$ (30)
Proof of Proposition 4.
For any $i\geq 0$, we have
$\|\widehat{\nabla\Phi}_{i}-\nabla\Phi_{i}(x_{i})\|$ $\leq(\sigma_{i}\|\widehat{\nabla F}_{i}-\nabla F(x_{i})\|+\|\widehat{\nabla G}_{i}-\nabla G(x_{i})\|)$
$\leq(\sigma_{0}\|\widehat{\nabla F}_{i}-\nabla F(x_{i})\|+\|\widehat{\nabla G}_{i}-\nabla G(x_{i})\|).$ (31)
Moreover, for any $i>q$, by the mean value theorem there exists $c_{i}\in[s_{i},i]$ such that
$\frac{(i^{1-2w}-s_{i}^{1-2w})}{(1-2w)}=c_{i}^{-2w}(i-s_{i})\leq\frac{q}{(i-q)^{2w}}.$
Given $\delta\in(0,1)$ and $i>q$, by Lemma 13 and the inequality Eq. 31, with probability at least $1-\delta/(i+1)(i+2)$, we have
$\|\widehat{\nabla\Phi}_{i}-\nabla\Phi_{i}(x_{i})\|$ $\leq\frac{8(\sigma_{0}L_{f}+L_{g})D}{(i-q)^{w}}\sqrt{\log\left(\frac{8(i+1)(i+2)}{\delta}\right)}.$
Hence, if $t>q$, with probability at least $1-\delta\sum_{q<i\leq t}1/(i+1)(i+2)$, we have
$\sum_{q<i\leq t}\|\widehat{\nabla\Phi}_{i}-\nabla\Phi_{i}(x_{i})\|$ $\leq\sum_{q<i\leq t}\frac{8(\sigma_{0}L_{f}+L_{g})D}{(i-q)^{w}}\sqrt{\log\left(\frac{8(i+1)(i+2)}{\delta}\right)}$
$\leq\frac{8(\sigma_{0}L_{f}+L_{g})D(t-q)^{1-w}}{1-w}\sqrt{\log\left(\frac{8(t+2)^{2}}{\delta}\right)}$
where the last inequality follows from the Riemann sum approximation. ∎
We now instantiate the regularization parameter and analyze Algorithm 2.
Lemma 15.
Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm 2 with $\alpha_{t}=\alpha:=\log(q+1)/(q+1)$ for any $0\leq t\leq q$, $\alpha_{t}=1/(t+1)^{\omega}$ for any $t>q$ and $S=q$. If 1 holds with the exception that $F$ may be nonconvex and 4 is satisfied, then given $\delta\in(0,1)$, with probability at least $1-\delta\sum_{0\leq i<t-1}1/(i+1)(i+2)$, for any $t>q+1$, it holds that
$t^{\omega}\mathcal{G}(x_{t})$ $\leq\frac{\mathcal{G}(x_{0})}{(q+1)^{1-\omega}}+\left(2\varsigma(q+1)^{\omega-p}+\varsigma(q+1)^{-p}\right)\sup_{z\in X}|F(z)|$
$\quad+\frac{(\sigma_{0}L_{f}+L_{g})D^{2}\log(q+1)}{2(q+1)^{1-w}}\left(1+16\sqrt{\log\left(\frac{8(q+2)^{2}}{\delta}\right)}\right)$
$\quad+\frac{8(\varsigma L_{f}+L_{g})D^{2}(t-1-q)^{1-w}}{1-w}\sqrt{\log\left(\frac{8(t+1)^{2}}{\delta}\right)}+\frac{(\varsigma L_{f}+L_{g})D^{2}}{1-\omega}t^{1-\omega}$
$\quad+2\varsigma\sup_{z\in X}|F(z)|t^{\omega-p}-\sigma_{0}(q+1)^{\omega}\sum_{i=0}^{q}\alpha\mathcal{F}(x_{i})(1-\alpha)^{q-i}-\sum_{i=q+1}^{t-1}\sigma_{i}\mathcal{F}(x_{i}).$
Proof of Lemma 15.
By Lemma 12, for any $t\geq q+1$, we have
$(t+1)^{\omega}\mathcal{G}(x_{t+1})$ $\leq(q+1)^{\omega}\mathcal{G}(x_{q+1})+2\varsigma\sup_{z\in X}|F(z)|(t+1)^{\omega-p}+\sigma_{q+1}F(x_{q+1})$
$\quad-\sum_{i=q+1}^{t}\sigma_{i}\mathcal{F}(x_{i})+\sum_{i=q+1}^{t}\left(2\|\widehat{\nabla\Phi}_{i}-\nabla\Phi_{i}(x_{i})\|D+\frac{(\sigma_{i}L_{f}+L_{g})D^{2}}{2(i+1)^{\omega}}\right).$
Applying Lemma 14, with probability at least $1-\delta\sum_{0\leq i\leq q}1/(i+1)(i+2)$, we have
$\mathcal{G}(x_{q+1})$ $\leq\frac{1}{q+1}\mathcal{G}(x_{0})+2\sigma_{0}\sup_{z\in X}|F(z)|-\sigma_{0}\sum_{i=0}^{q}\alpha\mathcal{F}(x_{i})(1-\alpha)^{q-i}$
$\quad+\frac{(\sigma_{0}L_{f}+L_{g})D^{2}\log(q+1)}{2(q+1)}\left(1+16\sqrt{\log\left(\frac{8(q+2)^{2}}{\delta}\right)}\right).$
Combining these two bounds with Proposition 4 then yields
$(t+1)^{\omega}\mathcal{G}(x_{t+1})$ $\leq\frac{1}{(q+1)^{1-\omega}}\mathcal{G}(x_{0})+2\varsigma(q+1)^{\omega-p}\sup_{z\in X}|F(z)|$
$\quad+\frac{(\sigma_{0}L_{f}+L_{g})D^{2}\log(q+1)}{2(q+1)^{1-w}}\left(1+16\sqrt{\log\left(\frac{8(q+2)^{2}}{\delta}\right)}\right)$
$\quad+\frac{8(\sigma_{0}L_{f}+L_{g})D(t-q)^{1-w}}{1-w}\sqrt{\log\left(\frac{8(t+2)^{2}}{\delta}\right)}$
$\quad+2\varsigma\sup_{z\in X}|F(z)|(t+1)^{\omega-p}+\frac{(\varsigma L_{f}+L_{g})D^{2}}{1-\omega}(t+1)^{1-\omega}+\sigma_{q+1}F(x_{q+1})$
$\quad-\sigma_{0}(q+1)^{\omega}\sum_{i=0}^{q}\alpha\mathcal{F}(x_{i})(1-\alpha)^{q-i}-\sum_{i=q+1}^{t}\sigma_{i}\mathcal{F}(x_{i}).$
The proof concludes using the simple inequality $\sigma_{q+1}F(x_{q+1})\leq\varsigma(q+2)^{-p}\sup_{z\in X}|F(z)|$. ∎
We are now well-equipped to prove Theorem 6.
Proof of Theorem 6.
We first focus on the terms involving $\mathcal{F}(x_{i})$ in Lemma 15. Note that the function $\mathcal{F}$ is defined with respect to $X_{\operatorname{opt}}$, not $X$. Therefore, $\mathcal{F}(x_{i})$ can take both positive and negative values since $x_{i}\in X$. However, we can derive the following bound
$\textstyle-\sigma_{0}(q+1)^{\omega}\sum_{i=0}^{q}\alpha\mathcal{F}(x_{i})(1-\alpha)^{q-i}-\sum_{i=q+1}^{t-1}\sigma_{i}\mathcal{F}(x_{i})$
$\leq$ $\sup_{z\in X}|\mathcal{F}(z)|\textstyle\left(\sigma_{0}(q+1)^{\omega}\sum_{i=0}^{q}\alpha(1-\alpha)^{q-i}+\sum_{i=q+1}^{t-1}\sigma_{i}\right)$
$=$ $\sup_{z\in X}|\mathcal{F}(z)|\textstyle\left(\sigma_{0}(q+1)^{\omega}\left(1-(1-\alpha)^{q+1}\right)+\sum_{i=q+1}^{t-1}\sigma_{i}\right)$
$\leq$ $\sup_{z\in X}|\mathcal{F}(z)|\textstyle\left(\varsigma(q+2)^{-p}(q+1)^{\omega}(1-(1-\alpha)^{q+1})+\varsigma\frac{t^{1-p}-(q+1)^{1-p}}{1-p}\right)$
$\leq$ $\varsigma\sup_{z\in X}|\mathcal{F}(z)|\textstyle\left((q+1)^{\omega-p}+\frac{t^{1-p}-(q+1)^{1-p}}{1-p}\right),$
where $\alpha=\log(q+1)/(q+1)$, and the second and third inequalities follow from the geometric series bound and the definition of $\sigma_{t}$. Using this bound together with Lemma 15, we arrive at
$\mathcal{G}(x_{t})$ $\leq\left(\frac{\mathcal{G}(x_{0})}{(q+1)^{1-\omega}}+\left(2\varsigma(q+1)^{\omega-p}+\varsigma(q+2)^{-p}\right)\sup_{z\in X}|F(z)|\right)t^{-w}$
$\quad+\frac{(\sigma_{0}L_{f}+L_{g})D^{2}\log(q+1)}{2(q+1)^{1-w}}\left(1+16\sqrt{\log\left(\frac{8(q+2)^{2}}{\delta}\right)}\right)t^{-w}$
$\quad+\frac{8(\varsigma L_{f}+L_{g})D^{2}(t-1-q)^{1-w}t^{-w}}{1-w}\sqrt{\log\left(\frac{8(t+1)^{2}}{\delta}\right)}+\frac{(\varsigma L_{f}+L_{g})D^{2}}{1-\omega}t^{1-2\omega}\!\!\!$
$\quad+2\varsigma\sup_{z\in X}|F(z)|t^{-p}+\left(\varsigma(q+1)^{\omega-p}+\varsigma\frac{(t+1)^{1-p}-q^{1-p}}{1-p}\right)t^{-w}\sup_{z\in X}|\mathcal{F}(z)|.$
We next focus on the upper-level problem. Since $\mathcal{G}(x_{t+1})\geq 0$, using Lemma 15 and the definition of the sequence $\{\beta_{t}\}_{t\geq 0}$, we obtain
$\sum_{i=0}^{t}\beta_{i}\mathcal{F}(x_{i})$ $\leq\frac{1}{(q+1)^{1-\omega}}\mathcal{G}(x_{0})+\left(2\varsigma(q+1)^{\omega-p}\sup_{z\in X}+\varsigma(q+2)^{-p}\right)\sup_{z\in X}|F(z)|$
$\quad+\frac{(\sigma_{0}L_{f}+L_{g})D^{2}\log(q+1)}{2(q+1)^{1-w}}\left(1+16\sqrt{\log\left(\frac{8(q+2)^{2}}{\delta}\right)}\right)$
$\quad+\frac{8(\sigma_{0}L_{f}+L_{g})D(t-q)^{1-w}}{1-w}\sqrt{\log\left(\frac{8(t+2)^{2}}{\delta}\right)}$
$\quad+2\varsigma\sup_{z\in X}|F(z)|(t+1)^{\omega-p}+\frac{(\varsigma L_{f}+L_{g})D^{2}}{1-\omega}(t+1)^{1-\omega}.$
Observe next that
$\sum_{i=0}^{t}\beta_{i}$ $\geq\varsigma(q+2)^{-p}(q+1)^{\omega}(1-(1-\alpha)^{q+1})+\varsigma\frac{t^{1-p}-(q+1)^{1-p}}{1-p}$
$\geq\varsigma\frac{(q+2)^{w-p}}{3}+\varsigma\frac{t^{1-p}-(q+1)^{1-p}}{1-p}:=a+bt^{1-p}.$
where $a$ and $b$ are constants defined to make the equality hold. Combining these bounds, we arrive at
$\frac{\sum_{i=0}^{t}\beta_{i}\mathcal{F}(x_{i})}{\sum_{i=0}^{t}\beta_{i}}$ $\leq\left(\frac{\mathcal{G}(x_{0})}{(q+1)^{1-\omega}}+\left(2\varsigma(q+1)^{\omega-p}+\varsigma(q+2)^{-p}\right)\sup_{z\in X}|F(z)|\right)\left(a+bt^{1-p}\right)^{-1}$
$\quad+\frac{(\sigma_{0}L_{f}+L_{g})D^{2}\log(q+1)}{2(q+1)^{1-w}}\left(1+16\sqrt{\log\left(\frac{8(q+2)^{2}}{\delta}\right)}\right)\left(a+bt^{1-p}\right)^{-1}$
$\quad+\frac{8(\sigma_{0}L_{f}+L_{g})D(t-q)^{1-w}}{1-w}\sqrt{\log\left(\frac{8(t+2)^{2}}{\delta}\right)}\left(a+bt^{1-p}\right)^{-1}$
$\quad+2\varsigma\sup_{z\in X}|F(z)|(t+1)^{\omega-p}\left(a+bt^{1-p}\right)^{-1}$
$\quad+\frac{(\varsigma L_{f}+L_{g})D^{2}}{1-\omega}(t+1)^{1-\omega}\left(a+bt^{1-p}\right)^{-1}.$
Both bounds involving $\mathcal{G}$ and $\mathcal{F}$ hold with probability $\geq 1-\delta$. We want the right hand side of both bounds $\to 0$. In order to guarantee this, we choose $p,\omega$ to minimize the slowest rate in terms of $t$:
$\min_{p,w:0<p\leq\omega}\max\left\{t^{p-1},t^{w-1},t^{p-w},t^{-w},t^{-p},t^{1-2w},t^{1-w-p}\right\}=-\frac{1}{4},$
which is realized by setting $p=1/2$, $\omega=3/4$ as required. ∎
B.3 Additional numerical experiments
We present an additional numerical for a dictionary learning problem to illustrate the performance of Algorithms 1 and 2 in the nonconvex setting. We highlight that, similar to Theorems 3 and 4, one can show from Theorems 5 and 6 that there exists a subsequence of the iterates $\{x_{t}\}_{t\geq 0}$ generated by Algorithms 1 and 2 that converges almost surely to a stationary point of Eq. 1. This follows from the fact that the auxiliary function $\mathcal{F}$ can be shown to be Lipschitz continuous. Moreover, using the same argument as in the proof of [4, Theorem 4.5 and 4.7], it is possible to show that there exists a point $x_{t^{\star}}$ in a subsequence of $\{x_{t}\}_{t\geq 0}$ that satisfies a stationarity guarantee similar to that in Theorems 5 and 6. Since this is merely an existence result and the proof is long, we omit the details for brevity. Nonetheless, we present the convergence behavior of the last iterate of the sequence $\{x_{t}\}_{t\geq 0}$ in a dictionary learning problem, which is also studied in [4, 28].
In the dictionary learning problem, the goal is to learn a compact representation of the input data $A=\{a_{1},\dots,a_{n}\}\subseteq{\mathbf{R}}^{m}$. Formally, we aim to find a dictionary $D=[d_{1}\cdots d_{p}]\in{\mathbf{R}}^{m\times p}$ such that each data point $a_{i}$ can be approximated by a linear combination of the basis vectors in $D$. This leads to the following optimization problem
$\min_{D,X}&\quad\frac{1}{2n}\sum_{i\in[n]}\|a_{i}-Dx_{i}\|_{2}^{2}\\ \text{s.t.}~{}&\quad D\in{\mathbb{R}}^{m\times p},\ X=[x_{1}\cdots x_{n}]\in{\mathbb{R}}^{p\times n}\\ &\quad\|d_{j}\|_{2}\leq 1,\quad\forall j\in[p]\\ &\quad\|x_{i}\|_{1}\leq\delta,\quad\forall i\in[n],\end{aligned}$
where we refer to $X$ as the coefficient matrix.
In practice, data points usually arrive sequentially and the representation evolves gradually. Hence, the dictionary must be updated sequentially as well. Assume that we already have learned a dictionary $\hat{D}\in{\mathbf{R}}^{m\times p}$ and the corresponding coefficient matrix $\hat{X}\in{\mathbf{R}}^{p\times n}$ for some data set $A$. As a new dataset $A^{\prime}=\{a_{1}^{\prime},\dots,a_{n^{\prime}}^{\prime}\}\subseteq{\mathbf{R}}^{m}$ arrives, we intend to enrich our dictionary by learning $\tilde{D}\in{\mathbf{R}}^{m\times q}$ ($q>p$) and the coefficient matrix $\tilde{X}\in{\mathbf{R}}^{q\times n^{\prime}}$ for the new dataset while maintaining good performance of $\tilde{D}$ on the old dataset $A$ as well as the learned coefficient matrix $\hat{X}$ This leads to the following finite-sum bilevel problem:
$\min_{\tilde{D},\tilde{X}}$ $\quad F(\tilde{D},\tilde{X}):=\frac{1}{2n^{\prime}}\sum_{i\in[n^{\prime}]}\|a_{i}^{\prime}-\tilde{D}\tilde{x}_{i}\|_{2}^{2}$ (32)
s.t. $\quad\tilde{D}\in{\mathbb{R}}^{m\times q},\ \tilde{X}=[\tilde{x}_{1}\cdots\tilde{x}_{n}]\in{\mathbb{R}}^{q\times n^{\prime}}$
$\quad\|\tilde{x}_{i}\|_{1}\leq\delta,\quad\forall i\in[n^{\prime}],$
$\quad\tilde{D}\in\operatorname*{arg\,min}_{\|z_{j}\|_{2}\leq 1,\forall j\in[q]}g(Z):=\frac{1}{2n}\sum_{i\in[n]}\|a_{i}-Z\hat{x}_{i}\|_{2}^{2},$
where we denote $\hat{x}_{i}$ as the prolonged vector in ${\mathbf{R}}^{q}$ by appending zeros at the end.
In our experiment, we obtain $A,A^{\prime},\hat{X}$ from the code provided by Cao et al. [4] with $n=n^{\prime}=250$ and $p=40,q=50$ In addition, we also choose $\delta=3$. The experiment is repeated $10$ times, and we report the average performance within a $1$-minute execution time limit. For IR-SCG, presented in Algorithm 1, we set $\sigma_{t}=\varsigma(t+1)^{-2/7}$ and $\alpha_{t}=(t+1)^{-6/7}$, where $\varsigma=0.1$. For IR-FSCG, presented in Algorithm 2, we set $S=q=\lfloor\sqrt{n}\rfloor$, $\sigma_{t}=\varsigma(\max\{t,q+1\}+1)^{-1/2}$, where $\varsigma=0.1$, along with $\alpha_{t}=(t+1)^{-3/4}$ for every $t\geq q+1$ and $\alpha_{t}=\log(q+1)/(q+1)$ for every $t\leq q$. In both experiments, the linear optimization oracle admits an analytical solution; see [25, Section 4.1].
For performance comparison, we implement four algorithms: SBCGI [4, Algorithm 1] with stepsize $\gamma_{t}=0.1(t+1)^{-2/3}$, SBCGF [4, Algorithm 2] with constant stepsize $\gamma_{t}=10^{-3}$ (both using $K_{t}=0.01(t+1)^{-1/3},S=q=\lfloor\sqrt{n}\rfloor$), aR-IP-SeG [26] with long ($\gamma_{t}=10^{-2}/(t+1)^{3/4}$) and short stepsizes ($\gamma_{t}=10^{-4}/(t+1)^{3/4}$, $\rho_{t}=(t+1)^{1/4}$, $r=0.5$), and the stochastic variant of the dynamic barrier gradient descent (SDBGD) [20] with long ($\gamma_{t}=10^{-2}$) and short ($\gamma_{t}=5\times 10^{-3}$) stepsizes. We initialized all algorithms with randomized starting points. For SBCGI and SBCGF, we generated the required initial point $x^{\prime}_{0}$ by running the SPIDER-FW algorithm [57, Algorithm 2] on the inner-level problem in (6), using a stepsize of $\gamma_{t}=0.1/(t+1)$. The initialization process terminated either after $10^{5}$ stochastic oracle queries, with the resulting point serving as $x_{0}$ for both SBCGI and SBCGF. The implementation of these algorithms rely on certain linear optimization or projection oracles. For SBCGI and SBCGF, a linear optimization oracle over the $\ell_{2}$-norm ball intersecting with a half-space is required, which we employ CVXPY [10] to solve the problems, similar to the implementation in [4, Appendix F.1]. To compute the projection onto the feasible set required for aR-IP-SeG, we used [12, Algorithm 1] to compute projection on $\ell_{1}$ norm ball and projection onto $\ell_{2}$ norm ball admits an analytical solution. To approximate inner-level optimal value $G_{\operatorname{opt}}$, we again employ CVXPY to solve the inner-level problem.
Figure 3 shows the convergence behavior of different algorithms for the inner-level (left) and outer-level (right) problems in terms of execution time. Our proposed stochastic methods outperform existing approaches. In particular, IR-FSCG achieves better solutions for the outer-level objective while maintaining strong inner-level feasibility, surpassing its mini-batch counterpart, SBCGI. Similarly, IR-SCG outperforms its single-sample counterpart, SBCGI. Finally, the table below reports the average number of iterations and stochastic oracle calls over the 4-minute time limit for each algorithm. Our projection-free method completes more iterations within the fixed time budget, benefiting from a simpler linear minimization oracle compared to the projection-free approach in [4]. We also observe that SDBGD and IR-SCG perform less number of iterations compared to our methods.
Method Average # of iterations Average # of oracle calls
IR-SCG $32736.1$ $65472.2$
IR-FSCG $30532.6$ $1872851.0$
SBCGI $466.3$ $7440.6$
SBCGF $361.84$ $6147916.0$
aR-IP-SeG (short step) $6290.0$ $25160.0$
aR-IP-SeG (long step) $6278.1$ $25112.4$
SDBGD (short step) $11441.3$ $22882.6$
SDBGD (long step) $12239.3$ $24478.6$
Figure 3: The inner-level optimality gap (left) and the outer-level objective function (right) over time for different algorithms on the dictionary learning problem (32).
B.4 Finiteness of constants
The following lemma establishes that the quantities introduced in (15) are finite.
Lemma 16.
If 1 and 1 to 3 hold, then $C,\bar{C}_{\delta},C_{\delta}$ defined in (15) are finite. Furthermore, if $L$ defined in 2 is strictly less than $1$, then $V$ defined in (15d) is finite.
Proof of Lemma 16.
We first show that under 1 and 3, when the limit in 2 exists, the parameter $L$ satisfies $0\leq L\leq 1$. While it is trivial to see $L\geq 0$ thanks to 1, $L\leq 1$ needs some justifications. For the sake of contradiction, suppose $L>1$. Then for some sufficiently large $t$, we have
$t\left(\frac{\sigma_{t}}{\sigma_{t+1}}-1\right)>1\iff\frac{\sigma_{t}}{\sigma_{t+1}}>\frac{t+1}{t}\iff(t+1)\sigma_{t+1}<t\sigma_{t}.$
This further implies
$(t+1)\sigma_{t+1}^{2}\leq(t+1)\sigma_{t+1}\sigma_{t}\leq t\sigma_{t}^{2},$
which always contradicts 3. Thus, we have $0\leq L\leq 1$ under our blanket assumptions.
We now show that $C$ is finite. Observe that
$\lim_{t\to\infty}\frac{(t+1)t(\sigma_{t-1}-\sigma_{t})}{(t+1)t\sigma_{t}-t(t-1)\sigma_{t-1}}=\lim_{t\to\infty}\frac{(1+\frac{2}{t-1})(t-1)(\frac{\sigma_{t-1}}{\sigma_{t}}-1)}{2-(t-1)(\frac{\sigma_{t-1}}{\sigma_{t}}-1)}=\frac{L}{2-L}$
where the first equality is obtained by diving both the numerator and the denominator by $t\sigma_{t}$ and the second equality follows from 2. Moreover, we have
$\lim_{t\to\infty}\frac{(t+1)t(\sigma_{t-1}-\sigma_{t})}{(t+1)t\sigma_{t}-t(t-1)\sigma_{t-1}}\!$ $=\!\lim_{t\to\infty}\frac{\sum_{i\in[t]}(i+1)i(\sigma_{i-1}-\sigma_{i})-\sum_{i\in[t-1]}(i+1)i(\sigma_{i-1}-\sigma_{i})}{(t+1)t\sigma_{t}-t(t-1)\sigma_{t-1}}$
$=\lim_{t\to\infty}\frac{\sum_{i\in[t]}(i+1)i(\sigma_{i-1}-\sigma_{i})}{(t+1)t\sigma_{t}}$
where the second equality follows from the Stolz–Cesàro theorem. This yields
$\lim_{t\to\infty}\frac{\sum_{i\in[t]}(i+1)i(\sigma_{i-1}-\sigma_{i})}{(t+1)t\sigma_{t}}=\frac{L}{2-L}\implies\sup_{t\geq 1}\frac{\sum_{i\in[t]}(i+1)i(\sigma_{i-1}-\sigma_{i})}{(t+1)t\sigma_{t}}\in(0,\infty).$ (33)
Using 2, we have $\min_{t\geq 0}(t+1)\sigma_{t}>0$. Thus, $C$ is finite. It is also straightforward to verify that $\bar{C}_{\delta}$ is finite by 3. Consequently, $C_{\delta}$ is finite as well.
To conclude, we show $V$ is finite. From 3, $\{t\sigma_{t}\}_{t\geq 0}$ and $\{(t+1)t\sigma_{t}^{2}\}_{t\geq 0}$ are increasing and diverge to $\infty$. Note that
$\lim_{t\to\infty}\frac{\sum_{i\in[t]}(i+1)i(\sigma_{i-1}-\sigma_{i})\sigma_{i}}{(t+1)t\sigma_{t}^{2}}$ $=\lim_{t\to\infty}\frac{(t+2)(t+1)(\sigma_{t}-\sigma_{t+1})\sigma_{t+1}}{(t+2)(t+1)\sigma_{t+1}^{2}-(t+1)t\sigma_{t}^{2}}$
$=\lim_{t\to\infty}\frac{\left(\frac{t+2}{t}\right)t\left(\frac{\sigma_{t}}{\sigma_{t+1}}-1\right)}{\left(2-t\left(\frac{\sigma_{t}}{\sigma_{t+1}}-1\right)\right)\left(\frac{\sigma_{t}}{\sigma_{t+1}}+1\right)-2\frac{\sigma_{t}}{\sigma_{t+1}}}=\frac{L}{2(1-L)}$
where the first equality is implied by the Stolz-Cesàro theorem, the second equality is obtained from dividing both the denominator and the numerator by $(t+1)\sigma_{t+1}^{2}$, and the third equality comes from 3 and
$\lim_{t\to\infty}\frac{\sigma_{t}}{\sigma_{t+1}}=1+\lim_{t\to\infty}\frac{1}{t}\cdot t\left(\frac{\sigma_{t}}{\sigma_{t+1}}-1\right)=1.$
Note that $L/2(1-L)\in[0,\infty)$ since $L\in[0,1)$. Thus, $V$ is finite. This completes the proof. ∎
The following lemma establishes that the quantities introduced in (23) are finite.
Lemma 17.
If 1, 1, 2 and 4 hold, then $C_{q},\bar{C}_{\delta,q}$ defined in (23f), and (23g) are finite. Furthermore, if $L$ defined in 2 is less than $1$, then $V_{q}$ defined in (23h) is finite.
Proof of Lemma 17.
The proof is similar to the proof of Lemma 16. Details are omitted for brevity. ∎

### v2

@@
-Efficient Conditional Gradient Methods for Solving Stochastic Convex Bilevel Optimization Problems
+Conditional Gradient Methods with Standard LMO for Stochastic Simple Bilevel Optimization
 Abstract
-We propose efficient methods for solving stochastic convex bilevel optimization problems, where the goal is to minimize an outer stochastic objective function subject to the solution set of an inner stochastic optimization problem.
-Existing methods often rely on costly projection or linear optimization oracles over complex sets, which limits scalability.
-To overcome this, we propose an iteratively regularized conditional gradient framework that leverages efficient linear optimization oracles exclusively over the base feasible set.
+We propose efficient methods for solving stochastic simple bilevel optimization problems with convex inner levels, where the goal is to minimize an outer stochastic objective function subject to the solution set of an inner stochastic optimization problem.
+Existing methods often rely on costly projection or linear optimization oracles over complex sets, limiting their scalability.
+To overcome this, we propose an iteratively regularized conditional gradient approach that leverages linear optimization oracles exclusively over the base feasible set.
 Our proposed methods employ a vanishing regularization sequence that progressively emphasizes the inner problem while biasing towards desirable minimal outer objective solutions.
-Under standard convexity assumptions, we establish non-asymptotic convergence rates of $O(t^{-({1}/{2}-p)})$ for the outer objective and $O(t^{-p})$ for the inner objective, where $p\in(0,1/2)$ controls the regularization decay, in the one-sample stochastic setting, and $O(t^{-(1-p)})$ and $O(t^{-p})$ in the finite-sum setting using a mini-batch scheme, where $p\in(0,1)$.
-Experimental results on over-parametrized regression and $\ell_{1}$-constrained logistics regression tasks demonstrate the practical advantages of our approach over existing methods, confirming our theoretical findings.
+In the one-sample stochastic setting and under standard convexity assumptions, we establish non-asymptotic convergence rates of $\tilde{O}(t^{-1/4})$ for both the outer and inner objectives.
+In the finite-sum setting with a mini-batch scheme, the corresponding rates become $\tilde{O}(t^{-1/2})$.
+When the outer objective is nonconvex, we prove non-asymptotic convergence rates of $\tilde{O}(t^{-1/7})$ for both the outer and inner objectives in the one-sample stochastic setting, and $\tilde{O}(t^{-1/4})$ in the finite-sum setting.
+Experimental results on over-parametrized regression and dictionary learning tasks demonstrate the practical advantages of our approach over existing methods, confirming our theoretical findings.
 1 Introduction
-We consider stochastic convex bilevel optimization problems, where the goal is to minimize an outer stochastic convex objective function subject to the solution set of an inner stochastic convex optimization problem.
+We consider stochastic simple bilevel optimization problems, where the goal is to minimize an outer stochastic objective function subject to the solution set of an inner stochastic convex optimization problem.
 Formally, the problem is defined as
-$F_{\operatorname{opt}}:=\min_{x\in{\mathbb{R}}^{d}}\left\{F(x):={\mathbb{E}}[f(x,\theta)]:x\in X_{\operatorname{opt}}\right\},\ \text{where }X_{\operatorname{opt}}:=\operatorname*{arg\,min}_{z\in X}G(z):={\mathbb{E}}[g(z,\xi)].$ (1)
-Here $X\subset\mathbb{R}^{d}$ is a convex and compact set, and both $F$ and $G$ are smooth and convex, making the optimization problem (1) convex.
+$F_{\operatorname{opt}}:=\min_{x\in{\mathbf{R}}^{d}}\left\{F(x):={\mathbb{E}}[f(x,\theta)]:x\in X_{\operatorname{opt}}\right\},\ \text{where }X_{\operatorname{opt}}:=\operatorname*{arg\,min}_{z\in X}\left\{G(z):={\mathbb{E}}[g(z,\xi)]\right\}.$ (1)
+Here $X\subset\mathbb{R}^{d}$ is a compact convex set, and both $F$ and $G$ are smooth with $G$ is additionally convex, making the optimization problem (1) a possibly nonconvex problem with convex domain.
 This framework has broad applicability, including hyper-parameter optimization [16, 17, 39], meta-learning [17, 42], reinforcement learning [32] and game theory [50].
 In particular, problem (1) arises in learning applications where $G$ is the prediction error of a model $x$ on data, and $F$ is an auxiliary objective (e.g., regularization, or a validation data set).
-The dependence on data is through the expectation operators and the variables $\theta,\xi$.
+The dependence on data is through the expectation operators and the random variables $\theta,\xi$.
 The bilevel formulation provides a tuning-free alternative to the usual regularization approach where we optimize $\sigma F(x)+G(x)$ for a carefully tuned parameter $\sigma$.
 Despite its wide applications in machine learning there are three primary challenges in solving (1).
-First, we do not have an explicit representation of the optimal set $X_{\operatorname{opt}}$ in general, which prevents us from using some common operations in optimization such as projection onto or linear optimization over $X_{\operatorname{opt}}$.
+First, we do not have an explicit representation of the optimal set $X_{\operatorname{opt}}$ in general, preventing us from using some common operations in optimization such as projection onto or linear optimization over $X_{\operatorname{opt}}$.
 An alternative is to reformulate (1) using its value function formulation
@@
 1.1 Related works
+Simple bilevel optimization problems with nonconvex outer objectives
+This setting has been studied by Jiang et al.
+[29] in the deterministic case and by Cao et al.
+[4] in the stochastic case.
+These works propose conditional gradient methods that require performing linear optimization over intersections of the base domain $X$ with a halfspace.
+In contrast, our approach removes this requirement by relying solely on linear optimization over $X$.
+We note that we do not address the more general case with both nonconvex inner and outer objectives considered in [21, 24, 5].
+In the remainder of this section, we will focus on simple bilevel problems with convex inner and outer objectives.
 Iterative regularization methods.
@@
 Friedlander and Tseng [18] established fundamental conditions ensuring solution existence for $\sigma>0$, forming the basis for iterative regularization, where $\sigma$ is gradually reduced.
-This idea was first implemented by Cabot [3] in the unconstrained case $X={\mathbb{R}}^{d}$ using a decreasing sequence $\sigma_{t}$, and later extended to general convex constraints by Dutta and Pandit [13].
+This idea was first implemented by Cabot [3] in the unconstrained case $X={\mathbf{R}}^{d}$ using a positive decreasing sequence $\sigma_{t}$, and later extended to general convex constraints by Dutta and Pandit [14].
 While these methods use proximal steps, which can be computationally expensive, Solodov [49] proposed a more efficient gradient-based alternative.
@@
 More recently, Merchav et al.
-[36] proposed an accelerated proximal scheme with a $O(\log(t+1)/t)$ rate under smoothness assumption and $O(t^{-2})$ / $O(1/t^{2-\gamma})$ under a Hölderian error bound condition.
+[37] proposed an accelerated proximal scheme with a $O(\log(t+1)/t)$ rate under smoothness assumption and $O(t^{-2})$ and $O(1/t^{2-\gamma})$ for inner and outer objectives, respectively, under a Hölderian error bound condition with modulus $\gamma\in(1,2)$.
 Our work also follows an iterative regularization approach, but avoids expensive projections or proximal operations by leveraging linear optimization oracles.
@@
 [29] approximated $X_{\operatorname{opt}}$ by replacing the constraint $G(x)\leq G_{\operatorname{opt}}$ with a linear approximation, which removed the need for Slater’s condition, and applying the conditional gradient method, refining the approximation at each step.
-They established $O(t^{-1})$ convergence rates for both objectives under smoothness and compactness assumptions.
+They established $O(t^{-1})$ convergence rates for both objectives under smoothness and compactness assumptions in case $f$ is convex.
+Under non-convex setting, they claimed the convergence rates of $O(t^{-1/2})$ to the stationary point of the problem.
 However, their method requires a pre-specified tolerance parameter $\epsilon_{g}$, and only guarantees $(\epsilon_{g}/2)$-infeasibility.
@@
 [29] to the stochastic setting.
-Their methods achieve $\tilde{O}(t^{-1})$ convergence rate when the noise distributions have finite support, and $O(t^{-1/2})$ convergence rates under sub-Gaussian noise.
-However, their method requires linear optimization over intersections of $X$ with halfspaces.
+Their methods achieve $\tilde{O}(t^{-1})$ convergence rate when the noise distributions have finite support, and $\tilde{O}(t^{-1/2})$ convergence rates under sub-Gaussian noise.
+However, their method requires linear optimization over intersections of $X$ with a halfspace.
 In contrast, our work relaxes this requirement by relying solely on linear optimization over the base set $X$.
@@
 Several alternative approaches exist for simple bilevel optimization, including sequential averaging schemes [44, 46, 36], sublevel set-based methods [2, 6, 58, 21], accelerated algorithms [45, 54, 7, 6] and primal-dual strategies [47].
-These methods are less closely related to our framework and do not address the stochastic setting.
+These methods are less related to our framework and do not address the stochastic setting.
 More general bilevel problems have seen considerable attention, particularly in Stackelberg game formulations where the upper and lower levels model leader-follower dynamics [19, 8, 31, 28, 10, 33, 56, 34].
 Recent extensions include contextual bilevel optimization with exogenous variables [55, 25, 51] and problems with functional constraints at the lower level [30, 40].
-In contrast, our focus is restricted to the stochastic convex formulation of simple bilevel problems.
+Our focus in this paper, however, is restricted to the stochastic simple bilevel problems.
 1.2 Contributions
+We present two iterative regularization methods for solving problem (1) in both stochastic and finite-sum settings.
+The key idea is to introduce a decreasing regularization sequence $\{\sigma_{t}\}_{t\geq 1}$, and at each iteration $t$, perform a gradient-based update on the composite objective $\sigma_{t}F+G$.
+As $\sigma_{t}$ decreases, the algorithm gradually shifts focus from $F$ to $G$, thereby steering $x_{t}$ toward points in $X_{\operatorname{opt}}$.
+At the same time the $\sigma_{t}F$ term encourages $x_{t}$ towards a point in $X_{\operatorname{opt}}$ that minimizes $F$.
+In the stochastic formulation (1), computing exact gradients of the upper and lower objectives is computationally expensive.
+To address this, we employ sample-based gradient estimators.
+Specifically, we use the STOchastic Recursive Momentum (STORM) estimator [9] in the one-sample stochastic setting, and the Stochastic Path-Integrated Differential EstimatoR (SPIDER[38, 15] in the finite-sum setting.
 The key contributions of this paper are summarized below.
   • $\diamond$
-    One-sample stochastic setting: We develop a projection-free method for stochastic bilevel optimization that achieves high-probability convergence rates of $O(t^{-(1/2-p)}\sqrt{\log(dt^{2}/\delta)})$ for the outer-level problem and $O(t^{-p}\sqrt{\log(d/\delta)})$ for the inner-level problem, with probability at least $1-\delta$ for any $p\in(0,1/2)$.
+    One-sample stochastic setting: We develop a projection-free method for stochastic simple bilevel optimization that under convexity of $F$ on the base set $X$ achieves high-probability convergence rates of $O(t^{-1/4}\sqrt{\log(dt^{2}/\delta)})$ for the outer-level problem and $O(t^{-1/4}\sqrt{\log(d/\delta)})$ for the inner-level problem, with probability at least $1-\delta$.
+Without convexity of $F$, the method enjoys the non-asymptotic convergence rates of $O(t^{-\frac{1}{7}}\log({dt^{2}}/{\delta}))$ for both the outer- and inner-level objectives.
 Compared to the state-of-the-art approach by Cao et al.
-[4, Algorithm 1], which achieves the $O(t^{-1/2}\sqrt{\log(Td/\delta)})$ rate for both the outer and inner problems, our method offers several key advantages.
-First, it eliminates the need for optimization over intersections of the base feasible set with halfspaces, requiring only linear optimization over the original constraint set.
+[4, Algorithm 1], which achieves the $O(t^{-1/2}\sqrt{\log(td/\delta)})$ rate under convex settings and $O(t^{-1/3}\sqrt{\log(td/\delta)})$ under nonconvex settings for both the outer and inner problems, our method offers several key advantages.
+First, it eliminates the need for optimization over intersections of the base domain with a halfspace, requiring only linear optimization over the original constraint set.
 Second, it removes the dependency on their pre-specified tolerance parameter $\epsilon_{g}$, and the associated computationally expensive initialization procedure to find an $\epsilon_{g}$-optimal starting point.
-Thus, our algorithm enjoys anytime guarantees.
+Third, our algorithm enjoys anytime guarantees.
+Overall, the proposed algorithm is a simple, single-loop procedure that is significantly easier to implement than that of [4].
   • $\diamond$
-    Finite-sum setting: For problems with $n$ component functions, we establish convergence rates of $O(t^{-(1-p)}\sqrt{\log(t^{2}/\delta)})$ (outer) and $O(t^{-p}\sqrt{\log(1/\delta)})$ (inner) with probability $1-\delta$, for any $p\in(0,1)$.
+    Finite-sum setting: For problems with $n$ component functions, we establish convergence rates of $O(t^{-1/2}\sqrt{\log(t^{2}/\delta)})$ (outer) and $O(t^{-1/2}\sqrt{\log(1/\delta)})$ (inner) with probability $1-\delta$.
+Without convexity of $F$, the method enjoys the non-asymptotic convergence rates of $O(t^{-\frac{1}{4}}\log({dt^{2}}/{\delta}))$ for both the outer- and inner-level objectives.
 Although Cao et al.
-[4, Algorithm 2] achieves improved $O(t^{-1}\log(T)\sqrt{\log(T/\delta)})$ rates for both objectives, it again relies on more computationally demanding halfspace-intersection oracles, requires an additional initialization step, and depends on a fixed stepsize.
+[4, Algorithm 2] achieves improved $O(t^{-1}\log(t)\sqrt{\log(t/\delta)})$ rates under convex settings and $O(t^{-1/2}\sqrt{\log(t/\delta)})$ rates under nonconvex settings for both objectives, it again relies on more computationally demanding halfspace-intersection oracles, requires an additional initialization step, depends on a fixed stepsize, and does not offer anytime guarantees.
   • $\diamond$
-    Numerical Validation: We demonstrate the practical efficiency of our methods on over-parametrized regression and $\ell_{1}$-constrained logistic regression tasks, showing significant speedups over existing projection-based and projection-free approaches.
+    Numerical validation: We demonstrate the practical efficiency of our methods on over-parametrized regression and dictionary learning tasks, showing significant speedups over existing projection-based and projection-free approaches.
 The experiments validate both the convergence rates and computational advantages of our framework.
-Proof of these results are provided in Appendix A. Additional results are presented in Appendix B.
-2 Algorithms and convergence guarantees
-In this section, we present two iterative regularization approaches for solving problem (1) in both stochastic and finite-sum settings.
-The core idea is to introduce a decreasing regularization sequence $\{\sigma_{t}\}_{t\geq 1}$, and at each iteration $t$, perform a gradient-based update on the composite objective $\sigma_{t}F+G$.
-As $\sigma_{t}$ decreases, the algorithm gradually shifts focus from $F$ to $G$, thereby steering $x_{t}$ toward points in $X_{\operatorname{opt}}$.
-At the same time the $\sigma_{t}F$ term encourages $x_{t}$ towards a point in $X_{\operatorname{opt}}$ that minimizes $F$.
-Giang-Tran et al.
-[19] applied this idea to design a projection-free algorithm for the deterministic setting, establishing convergence rates of $O(t^{-p})$ for the upper level and $O(t^{-(1-p)})$ for the lower level when using $\sigma_{t}=\varsigma(t+1)^{-p}$ with $p\in(0,1)$.
-In the stochastic setting Eq.
-1, however, computing exact gradients of the upper and lower objective functions is computationally expensive.
-To address this challenge, we rely on sample-based gradient estimators.
-Specifically, given mini-batch sample sets ${\cal S}_{f}:=\{\theta_{1},\ldots,\theta_{|{\cal S}_{f}|}\}$ and ${\cal S}_{g}:=\{\xi_{1},\ldots,\xi_{|{\cal S}_{g}|}\}$ drawn independently at each iteration, we define the empirical gradients as
-$\nabla f_{{\cal S}_{f}}(x):=\frac{1}{|{\cal S}_{f}|}\sum_{\theta\in{\cal S}_{f}}\nabla f(x,\theta)\quad\&\quad\nabla g_{{\cal S}_{g}}(x):=\frac{1}{|{\cal S}_{g}|}\sum_{\xi\in{\cal S}_{g}}\nabla g(x,\xi),$
-which provide unbiased estimates of $\nabla F(x)$ and $\nabla G(x)$, respectively.
-A naïve approach is to use these sample gradients directly, that is, $\widehat{\nabla F}(x)=\nabla f_{\mathcal{S}_{f}}(x)$ and $\widehat{\nabla G}(x)=\nabla g_{\mathcal{S}_{g}}(x)$.
-However, these estimators maintain constant variance over optimization time, which prevents achieving optimal convergence rates.
-To address this limitation, we implement two distinct variance-reduced gradient estimation schemes tailored to different problem settings.
-  • $\diamond$
-    One-sample stochastic setting: When we access single samples of $\theta$ and $\xi$ per iteration, we employ the STOchastic Recursive Momentum (STORM) estimator [8], which progressively reduces variance through momentum-based updates under sub-Gaussian noise assumptions.
-The resulting method, called the Iteratively Regularized Stochastic Conditional Gradient (IR-SCG) algorithm, is summarized in Algorithm 1.
-We note that the use of the STORM estimator in projection-free methods was first introduced in [60].
-  • $\diamond$
-    Finite-sum setting: When the expectations in (1) are finite sums of size $n$, we use the the Stochastic Path-Integrated Differential EstimatoR (SPIDER[37, 14] with $\sqrt{n}$-sized mini-batches, achieving variance reduction through periodic gradient recomputation.
-The resulting method, called the Iteratively Regularized Finite-Sum Conditional Gradient (IR-FSCG) algorithm, is summarized in Algorithm 2.
-Here, for any $n\in{\mathbb{N}}$, $[n]:=\{1,\dots,n\}$.
-We note that the SPIDER estimator in projection-free methods has been previously used in [48, 57, 59, 22].
-While a mini-batch scheme can be applied even when expectations are not finite sums, we emphasize that this approach only improves the constants in the convergence rates and enhances algorithmic stability without affecting the asymptotic rate.
-Additionally, if the upper- and lower-level problems have different numbers of component functions ($n_{f}\neq n_{g}$), Algorithm 2 can be easily modified by introducing $S_{f}$ and $q_{f}$ for the upper level, and $S_{g}$ and $q_{g}$ for the lower level.
-Data: Parameters $\{\alpha_{t}\}_{t\geq 0}\subseteq[0,1]$,$\{\sigma_{t}\}_{t\geq 0}\subseteq\mathbb{R}_{++}$.
+The supplementary material is organized in the appendices as follows.
+Appendix˜A establishes the preparatory lemmas used throughout our analysis.
+Appendices B and C contain the detailed proofs for the results presented in Section˜2 and Section˜3, respectively.
+Appendix˜D provides additional material, including analyses of the convergence rate coefficients as well as details of the numerical implementations in Section˜4.
+2 Algorithm and convergence results for the one-sample stochastic setting
+In this section, we introduce the Iteratively Regularized Stochastic Conditional Gradient (IR-SCG) method, summarized in Algorithm˜1.
+This approach is a conditional gradient method that leverages the STORM estimator to progressively reduce noise variance via momentum-based updates, assuming sub-Gaussian noise.
+While the use of the STORM estimator in projection-free methods was first explored in [60], our IR-SCG algorithm offers several key advantages.
+Notably, unlike the method of Cao et al.
+[4], IR-SCG is an anytime algorithm that does not require careful initialization, preset step sizes, or knowledge of the total number of iterations $T$ in advance.
+Data: Parameters $\{\alpha_{t}\}_{t\geq 0}\subseteq[0,1]$$\{\sigma_{t}\}_{t\geq 0}\subseteq\mathbb{R}_{++}$.
 Result: sequences $\{x_{t},z_{t}\}_{t\geq 1}$.
@@
   Compute
-$v_{t}$ $\in\operatorname*{arg\,min}_{v\in X}\left\{(\sigma_{t}\widehat{\nabla F}_{t}+\widehat{\nabla G}_{t})^{\top}v\right\}$
+$v_{t}$ $\in\operatorname*{arg\,min}_{v\in X}\left\{\left(\sigma_{t}\widehat{\nabla F}_{t}+\widehat{\nabla G}_{t}\right)^{\top}v\right\}$
 $x_{t+1}$ $:=x_{t}+\alpha_{t}(v_{t}-x_{t})$
@@
 $z_{t+1}$ $:=\frac{(t+2)(t+1)\sigma_{t+1}x_{t+1}+\sum_{i\in[t+1]}(i+1)i(\sigma_{i-1}-\sigma_{i})x_{i}}{S_{t+1}}$
+  -10pt
 Algorithm 1 Iteratively Regularized Stochastic Conditional Gradient (IR-SCG) Algorithm.
-Data: Parameters $\{\alpha_{t}\}_{t\geq 0}\subseteq[0,1]$, $\{\sigma_{t}\}_{t\geq 0}\subseteq\mathbb{R}_{++}$, $S,q\in[n]$.
-Result: sequences $\{x_{t},z_{t}\}_{t\geq 1}$.
-Initialize $x_{0}\in X$;
-for $t=0,1,2,\ldots$ do
-       if $t=0\mod q$ then
-            Compute
-$\widehat{\nabla F}_{t}:=\nabla F(x_{t}),\quad\widehat{\nabla G}_{t}:=\nabla G(x_{t}).$
-      else
-            Draw $S$ new i.i.d. samples ${\cal S}_{f}=\{\theta_{1},\ldots,\theta_{S}\}$, ${\cal S}_{g}=\{\xi_{1},\ldots,\xi_{S}\}$;
-             Compute
-$\widehat{\nabla F}_{t}:=\widehat{\nabla F}_{t-1}+\nabla f_{\mathcal{S}_{f}}(x_{t})-\nabla f_{\mathcal{S}_{f}}(x_{t-1})$
-$\widehat{\nabla G}_{t}:=\widehat{\nabla G}_{t-1}+\nabla g_{\mathcal{S}_{g}}(x_{t})-\nabla g_{\mathcal{S}_{g}}(x_{t-1}).$
-      Compute
-$v_{t}$ $\in\operatorname*{arg\,min}_{v\in X}\left\{(\sigma_{t}\widehat{\nabla F}_{t}+\widehat{\nabla G}_{t})^{\top}v\right\}$
-$x_{t+1}$ $:=x_{t}+\alpha_{t}(v_{t}-x_{t})$
-if $t\geq q$ then
-            Compute
-$S_{t+1}$ $:=(t+2)(t+1)\sigma_{t+1}+\sum_{i\in[t+1]\setminus[q]}(i+1)i(\sigma_{i-1}-\sigma_{i})$
-$z_{t+1}$ $:=\frac{(t+2)(t+1)\sigma_{t+1}x_{t+1}+\sum_{i\in[t+1]\setminus[q]}(i+1)i(\sigma_{i-1}-\sigma_{i})x_{i}}{S_{t+1}}.$
-Algorithm 2 Iteratively Regularized Finite-Sum Conditional Gradient (IR-FSCG) Algorithm.
-We emphasize that, unlike [4], both Algorithms 1 and 2 are anytime methods that do not require careful initialization step, preset stepsizes, or a priori choice of the number of total iterations $T$ based on target accuracies.
-This anytime property enables us to establish asymptotic convergence guarantees to the optimal solution, as formally stated in Theorems 3 and 4 in Appendix B.
-To establish convergence rates for these algorithms, we impose the following standard assumptions on the bilevel problem (1).
-In the following, all norms refer to the Euclidean norm.
+To establish convergence rates for Algorithm˜1, we impose the following standard assumptions on the simple bilevel problem (1). In the following, all norms refer to the Euclidean norm.
 Assumption 1.
@@
     ${\mathbb{E}}\left[\exp\left(\|\nabla_{x}g(x,\xi)-\nabla G(x)\|^{2}/\sigma_{g}^{2}\right)\right]$ $\leq\exp(1).$
-Our convergence rate guarantees are detailed in Theorems 2 and 1.
-Theorem 1 (One-sample stochastic setting).
-Let $\{z_{t}\}_{t\geq 1}$ be the iterates generated by Algorithm 1 with $\alpha_{t}=2/(t+2)$ for any $t\geq 0$, and regularization parameters $\{\sigma_{t}\}_{t\geq 0}$ given by $\sigma_{t}:=\varsigma(t+1)^{-p}$ for some chosen $\varsigma>0$, $p\in(0,1/2)$.
-Under 1, for any $t\geq 1$, with probability at least $1-\delta$, we have
-$F(z_{t})-F_{\operatorname{opt}}\leq O\left(t^{-(1/2-p)}\sqrt{\log\left(\frac{dt^{2}}{\delta}\right)}\right)\quad\&\quad G(z_{t})-G_{\operatorname{opt}}$ $\leq O\left(t^{-p}\sqrt{\log\left(\frac{d}{\delta}\right)}\right).$
-Theorem 2 (Finite sum setting).
-Suppose the expectations defining $F$ and $G$ in Eq.
-1 are from uniform distributions over finite sets of size $[n]$.
-Let $\{z_{t}\}_{t\geq 1}$ be the iterates generated by Algorithm 2 with $\alpha_{t}=\log(q)/q$ for $0\leq t<q$, $\alpha_{t}=2/(t+2)$ for any $t\geq q$, and $\sigma_{t}:=\varsigma(\max\{t,q\}+1)^{-p}$ for some chosen $\varsigma>0$, $p\in(0,1)$ and $S=q$.
-Under 1, with probability at least $1-\delta$, for any $t\geq q$, it holds that
-$F(z_{t})-F_{\operatorname{opt}}$ $\leq O\left(\max\left\{1,\frac{q\log(q)}{t}\right\}t^{-(1-p)}\sqrt{\log\left(\frac{t^{2}}{\delta}\right)}\right),$
-$G(z_{t})-G_{\operatorname{opt}}$ $\leq O\left(t^{-p}\sqrt{\log\left(\frac{1}{\delta}\right)}\right).$
-We present a proof sketch here; complete proofs are provided in Appendix A. Both results share a common structure, with key differences highlighted below.
-Proof sketch of Theorems 1 and 2.
-We employ the notation $\Phi_{t}(x):=\sigma_{t}F(x)+G(x)$.
-Let $\widehat{\nabla\Phi}_{t}=\sigma_{t}\widehat{\nabla F}_{t}+\widehat{\nabla G}_{t}$ be the gradient estimate at step $t$.
-Using standard analysis for conditional gradient methods with the stepsize $\alpha_{t}=2/(t+2)$, Lemmas 2 and 1 in Appendix A establish the inequality
-$t(t+1)\left(\Phi_{t}(x_{t})-\Phi_{t}(x_{\operatorname{opt}})\right)$ $+\textstyle\sum_{i\in[t]\setminus[k]}(i+1)i(\sigma_{i-1}-\sigma_{i})(F(x_{i})-F_{\operatorname{opt}})$
-$\textstyle\leq(k+1)k\left(\Phi_{k}(x_{k})-\Phi_{k}(x_{\operatorname{opt}})\right)+2(\sigma_{0}L_{f}+L_{g})D^{2}(t-k)$ (3)
-$\quad\textstyle+\sum_{i\in[t]\setminus[k]}4iD\|\widehat{\nabla\Phi}_{i-1}-\nabla\Phi_{i-1}(x_{i-1})\|$
-for every $t\geq k$.
-Proposition 1 in Appendix A uses (3) to provide bounds for both the upper- and lower-level problems.
-Note that in the proof of Algorithm 1, we set $k=0$, while in in the proof of Algorithm 2, we set $k=q$.
-Nevertheless, the underlying idea remains essentially the same.
-The key is to bound the last term $\sum_{i\in[t]}4iD\|\widehat{\nabla\Phi}_{i-1}-\nabla\Phi_{i-1}(x_{i-1})\|$, which is the error term from using noisy gradient estimates.
-To establish this bound with high probability, we leverage the concentration properties of the STORM estimator employed in Algorithm 1.
-Thanks to [4, Lemma 4.1], we obtain the following key probabilistic guarantee.
-For any iteration $i\in[t]$ and confidence level $\eta_{i}\in(0,1)$, with probability at least $1-\eta_{i}$, we (jointly) have
-$\left\|\widehat{\nabla F}_{i}-\nabla F(x_{i})\right\|$ $\leq O\left(\frac{\sqrt{\log(4d/\eta_{i})}}{(i+1)^{1/2}}\right)\quad\&\quad\left\|\widehat{\nabla G}_{i}-\nabla G(x_{i})\right\|\leq O\left(\frac{\sqrt{\log(4d/\eta_{i})}}{(i+1)^{1/2}}\right),$
-where the notation $O(\cdot)$ hides problem-specific constants $L_{f},L_{g},\sigma_{f},\sigma_{g},D$.
-Proposition 2 combines these inequalities for all $i\in[t]$ with $\eta_{i}=\delta/(i(i+1))$ and applies a union bound to conclude that with probability at least $1-\delta$, for any $t\geq 1$, we have
-$\sum_{i\in[t]}i\left\|\widehat{\nabla\Phi}_{i-1}-\nabla\Phi_{i-1}(x_{i-1})\right\|\leq O\left((t+1)^{3/2}\sqrt{\log\left({8dt^{2}}/{\delta}\right)}\right).$
-Substituting this bound into (3), and using the convexity of $F$ as well as the feasibility of $x_{t}\in X$, implying the inequality $G(x_{t})-G_{\operatorname{opt}}\geq 0$, we obtain
-$F(z_{t})-F_{\operatorname{opt}}\leq O\left(\frac{\sqrt{\log(8dt^{2}/\delta)}}{(t+1)^{1/2-p}}\right).$
-This in turn yields the bounds on $G(z_{t})-G_{\operatorname{opt}}$. Details are presented in Lemmas 6, 4 and 5.
-Algorithm 2 uses the SPIDER estimator, resulting in slightly different probabilistic bounds that depend on the frequency $q$ in which we compute the full gradient and the mini-batch size $S$.
-Thanks to [41, Theorem 3.5], we show in Lemma 7 that with probability at least $1-\eta_{i}$, we (jointly) have
-$\left\|\widehat{\nabla F}_{i}\!-\!\nabla F(x_{i})\right\|\leq O\left(\frac{\log(q)}{\sqrt{qS}}\sqrt{\log\left(\frac{8}{\eta_{i}}\right)}\right)~{}~{}\&~{}~{}\left\|\widehat{\nabla G}_{i}\!-\!\nabla G(x_{i})\right\|\leq O\left(\frac{\log(q)}{\sqrt{qS}}\sqrt{\log\left(\frac{8}{\eta_{i}}\right)}\right)$
-when $1\leq i\leq q$, and
-$\left\|\widehat{\nabla F}_{i}-\nabla F(x_{i})\right\|$ $\leq O\left(\sqrt{\frac{(i-s_{i})}{S(i+1)(s_{i}+1)}}\sqrt{\log\left(\frac{8}{\eta_{i}}\right)}\right)$
-$\left\|\widehat{\nabla G}_{i}-\nabla G(x_{i})\right\|$ $\leq O\left(\sqrt{\frac{(i-s_{i})}{S(i+1)(s_{i}+1)}}\sqrt{\log\left(\frac{8}{\eta_{i}}\right)}\right),$
-when $i>q$, where $s_{i}:=q\lfloor i/q\rfloor$ is the closest multiple of $q$ below $i$.
-Since the numerator of the fraction term $i-s_{i}$ is always bounded by $q$, the same aggregation trick as in Proposition 3, except ignoring the first $q$ terms (which are handled in Lemma 8), leads to an improved dependence.
-Specifically, we can show that, with probability $1-\delta$, we have
-$\sum_{i\in[t]\setminus[q]}i\left\|\widehat{\nabla\Phi}_{i-1}-\nabla\Phi_{i-1}(x_{i-1})\right\|\leq O\left((8\sqrt{2}t-4q)\sqrt{\log\left({16t^{2}}/{\delta}\right)}\right).$
-This improvement carries through to get the improved dependence on $t$ in Theorem 2; full details are given in Lemmas 8, 9, 10 and 11. ∎
-Remark 1 (Sample complexity).
-To attain an $\epsilon$-level optimality gap for both the outer- and inner-level problems, Algorithm 1 with $p=1/4$ requires $O(1/\epsilon^{4})$ stochastic oracle calls under the assumptions of Theorem 1.
-Likewise, under the assumptions of Theorem 2 and with $S=q=\lfloor\sqrt{n}\rfloor$ and $p=1/2$, Algorithm 2 achieves the same accuracy with $O(\sqrt{n}/\epsilon^{2})$ stochastic oracle calls.
+First, we present the result for the case that $F$ is convex.
+Theorem 1.
+Let $\{z_{t}\}_{t\geq 1}$ be the iterates generated by Algorithm˜1 with $\alpha_{t}=2/(t+2)$ for any $t\geq 0$, and regularization parameters $\sigma_{t}:=\varsigma(t+1)^{-p}$ for some chosen $\varsigma>0$, $p\in(0,1/2)$.
+Under ˜1, for any $t\geq 1$, with probability at least $1-\delta$, it (jointly) holds that
+$F(z_{t})-F_{\operatorname{opt}}\leq O\left(t^{-(1/2-p)}\sqrt{\log\left({dt^{2}}/{\delta}\right)}\right),\quad G(z_{t})-G_{\operatorname{opt}}$ $\leq O\left(t^{-p}\sqrt{\log\left({d}/{\delta}\right)}\right).$
+Moreover, with probability $1$, we have
+$\lim_{t\to\infty}F(x_{t})=F_{\operatorname{opt}},\quad\lim_{t\to\infty}G(x_{t})=G_{\operatorname{opt}}.$
+Theorem˜1 provides the formal convergence guarantees for the IR-SCG algorithm in the convex setting.
+If we set $p=1/4$, Algorithm˜1 achieves an $\epsilon$-level optimality gap for both the outer and inner problems with a sample complexity of $O(1/\epsilon^{4})$.
+We next extend our analysis to the more general case where $F$ is not necessarily convex.
+In this setting, we require a different measure for stationarity.
+We define our stationarity measure using the Frank-Wolfe gap, which for a convex function $\mathcal{F}(x)=0$ if and only if $x$ is optimal.
+This provides a natural generalization for the non-convex case.
+Accordingly, for all $x\in X$, we define the functions
+$\mathcal{F}(x):=\max_{v\in X_{\operatorname{opt}}}\nabla F(x)^{\top}(x-v),\quad\mathcal{G}(x):=G(x)-G_{\operatorname{opt}}.$ (3)
+To avoid clutter, we present the result using the optimally tuned parameters of Algorithm˜1.
+Theorem 2.
+Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm˜1 with stepsizes $\alpha_{t}=(t\!+\!1)^{-\omega}$ and regularization parameters $\sigma_{t}=\varsigma(t\!+\!1)^{-p}$ for any $t\!\geq\!0$, with $p\!=\!\frac{2}{7}$ and $\omega\!=\!\frac{6}{7}$.
+If Assumption 1 holds with the exception that $F$ may be nonconvex, then with probability at least $1-\delta$, for every $t\geq 0$, it (jointly) holds that
+$\frac{\sum_{i=0}^{t}\sigma_{i}\mathcal{F}(x_{i})}{\sum_{i=0}^{t}\sigma_{i}}\leq O\left(t^{-\frac{1}{7}}\log\left({dt^{2}}/{\delta}\right)\right),\quad\mathcal{G}(x_{t})\leq O\left(t^{-\frac{1}{7}}\log\left({dt^{2}}/{\delta}\right)\right).$
+Moreover, with probability $1$, we have
+$\liminf_{t\to\infty}\mathcal{F}(x_{t})=0,\quad\lim_{t\to\infty}{\cal G}(x_{t})=0.$
+3 Algorithm and convergence results for the finite-sum setting
+In this section, we address the case where the expectations in (1) are finite sums over $n$ components.
+We introduce the Iteratively Regularized Finite-Sum Conditional Gradient (IR-FSCG) method, summarized in Algorithm˜2, where $n]:=\{1,\dots,n\}$.
+This approach is a conditional gradient method that leverages the SPIDER estimator, achieving variance reduction through periodic gradient recomputations with mini-batches of size $q$.
+While the SPIDER estimator has been previously integrated into projection-free methods [48, 57, 59, 22], here we show how it can be effectively adapted to our regularized setting.
+The primary advantage is that, unlike the method of Cao et al.
+[4], IR-FSCG preserves the desirable anytime properties of our approach while achieving a faster convergence rate that capitalizes on the finite-sum structure.
+Additionally, it does not require careful initialization, preset step sizes, or knowledge of the total number of iterations $T$ in advance.
+To establish its convergence, we again rely on ˜1, noting that condition ˜1(e) holds automatically in this finite-sum setting. First, we present the result when $F$ is convex.
+Data: Parameters $\{\alpha_{t}\}_{t\geq 0}\subseteq[0,1]$, $\{\sigma_{t}\}_{t\geq 0}\subseteq\mathbb{R}_{++}$, $S,q\in[n]$.
+Result: sequences $\{x_{t},z_{t}\}_{t\geq 1}$.
+Initialize $x_{0}\in X$;
+for $t=0,1,2,\ldots$ do
+ if $t=0\mod q$ then
+     Compute
+$\widehat{\nabla F}_{t}:=\nabla F(x_{t}),\quad\widehat{\nabla G}_{t}:=\nabla G(x_{t}).$
+ else
+     Draw $S$ new i.i.d. samples ${\cal S}_{f}=\{\theta_{1},\ldots,\theta_{S}\}$, ${\cal S}_{g}=\{\xi_{1},\ldots,\xi_{S}\}$;
+      Compute
+$\widehat{\nabla F}_{t}:=\widehat{\nabla F}_{t-1}+\nabla f_{\mathcal{S}_{f}}(x_{t})-\nabla f_{\mathcal{S}_{f}}(x_{t-1})$
+$\widehat{\nabla G}_{t}:=\widehat{\nabla G}_{t-1}+\nabla g_{\mathcal{S}_{g}}(x_{t})-\nabla g_{\mathcal{S}_{g}}(x_{t-1}).$
+  Compute
+$v_{t}$ $\in\operatorname*{arg\,min}_{v\in X}\left\{\left(\sigma_{t}\widehat{\nabla F}_{t}+\widehat{\nabla G}_{t}\right)^{\top}v\right\}$
+$x_{t+1}$ $:=x_{t}+\alpha_{t}(v_{t}-x_{t})$
+if $t\geq q$ then
+     Compute
+$S_{t+1}$ $:=(t+2)(t+1)\sigma_{t+1}+\sum_{i\in[t+1]\setminus[q]}(i+1)i(\sigma_{i-1}-\sigma_{i})$
+$z_{t+1}$ $:=\frac{(t+2)(t+1)\sigma_{t+1}x_{t+1}+\sum_{i\in[t+1]\setminus[q]}(i+1)i(\sigma_{i-1}-\sigma_{i})x_{i}}{S_{t+1}}.$
+Algorithm 2 Iteratively Regularized Finite-Sum Conditional Gradient (IR-FSCG) Algorithm.
+Theorem 3.
+Suppose $F$ and $G$ in Eq.˜1 are expectations over uniform distributions on $[n]$.
+Let $\{z_{t}\}_{t\geq 1}$ be the iterates generated by Algorithm˜2 with $\alpha_{t}=\log(q)/q$ for $0\leq t<q$, $\alpha_{t}=2/(t+2)$ for any $t\geq q$, and $\sigma_{t}:=\varsigma(\max\{t,q\}+1)^{-p}$ for some chosen $\varsigma>0$, $p\in(0,1)$ and $S=q$.
+Under ˜1, with probability at least $1-\delta$, for any $t\geq q$, it holds that
+$F(z_{t})-F_{\operatorname{opt}}\leq O\left(c(t,q)\,t^{-(1-p)}\sqrt{\log\left({t^{2}}/{\delta}\right)}\right),\quad G(z_{t})-G_{\operatorname{opt}}\leq O\left(t^{-p}\sqrt{\log\left({1}/{\delta}\right)}\right),$
+where $c(t,q)=\max\left\{1,{q\log(q)}/{t}\right\}$. Moreover, with probability $1$, we have
+$\lim_{t\to\infty}F(x_{t})=F_{\operatorname{opt}},\quad\lim_{t\to\infty}G(x_{t})=G_{\operatorname{opt}}.$
+Theorem˜3 provides the formal convergence guarantees for the IR-FSCG algorithm in the convex setting.
+Moreover, when we set $S=q=\lfloor\sqrt{n}\rfloor$ and $p=1/2$, Algorithm˜2 achieves an $\epsilon$-level optimality gap for both the outer and inner problems with a sample complexity of $O(\sqrt{n}/\epsilon^{2})$.
+Besides, Algorithm˜2 can be readily extended when the outer- and inner-level problems involve different numbers of component functions ($n_{f}\neq n_{g}$) by introducing $(S_{f},q_{f})$ for the outer level and $(S_{g},q_{g})$ for the inner level.
+Details are omitted for brevity.
+We next extend our analysis to the more general case where $F$ is not necessarily convex, using the auxiliary function (3) to assess stationarity.
+To avoid clutter, we present the result using the optimally tuned parameters of Algorithm˜2.
+Theorem 4.
+Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm˜2 with stepsizes $\alpha_{t}=\log(q+1)/(q+1)$ for any $0\leq t\leq q$, $\alpha_{t}=1/(t+1)^{\omega}$ for any $t>q$ and $S=q$, and regularization parameters $\sigma_{t}=\varsigma(\max\{t,q+1\}+1)^{-p}$ with $p=\frac{1}{2}$ and $w=\frac{3}{4}$.
+If ˜1 holds with the exception that $F$ may be nonconvex, then with probability at least $1-\delta$, for every $t\geq q$, it (jointly) holds that
+$\frac{\sum_{i=0}^{t}\beta_{i}\mathcal{F}(x_{i})}{\sum_{i=0}^{t}\beta_{i}}\leq O\left(t^{-\frac{1}{4}}\log\left({dt^{2}}/{\delta}\right)\right),\quad\mathcal{G}(x_{t})\leq O\left(t^{-\frac{1}{4}}\log\left({dt^{2}}/{\delta}\right)\right),$
+where
+$\beta_{i}=\begin{cases}\displaystyle\frac{\sigma_{0}(q+1)^{\omega}\log(q+1)}{q+1}\left(1-\frac{\log(q+1)}{q+1}\right)^{q-i}&0\leq i\leq q,\\ \sigma_{i}&i>q.\end{cases}$ (4)
+Moreover, with probability $1$, we have
+$\liminf_{t\to\infty}\mathcal{F}(x_{t})=0,\quad\lim_{t\to\infty}{\cal G}(x_{t})=0.$
 4 Numerical results
-We showcase the performance of our proposed algorithms in an over-parametrized regression and an $\ell_{1}$-constrained logistic regression problems.
-Each experiment is repeated $10$ times, and we report the average performance within a $4$-minute execution time limit.
-For IR-SCG, presented in Algorithm 1, we set $\sigma_{t}=\varsigma(t+1)^{-1/4}$, where $\varsigma=10$ in the first experiment and $\varsigma=0.01$ in the second experiment, along with $\alpha_{t}=2/(t+2)$.
-For IR-FSCG, presented in Algorithm 2, we set $S=q=\lfloor\sqrt{n}\rfloor$, $\sigma_{t}=\varsigma(\max\{t,q\}+1)^{-1/2}$, where $\varsigma=10$ in the first experiment and $\varsigma=0.01$ in the second experiment, along with $\alpha_{t}=2/(t+1)$ for every $t\geq q$ and $\alpha_{t}=\log(q)/q$ for every $t<q$.
-In both experiments, the linear optimization oracle admits an analytical solution; see [25, Section 4.1].
-For performance comparison, we implement four algorithms: SBCGI [4, Algorithm 1] with stepsize $\gamma_{t}=0.01/(t+1)$, SBCGF [4, Algorithm 2] with constant stepsize $\gamma_{t}=10^{-5}$ (both using $K_{t}=10^{-4}/\sqrt{t+1},S=q=\lfloor\sqrt{n}\rfloor$), aR-IP-SeG [26] with long ($\gamma_{t}=10^{-2}/(t+1)^{3/4}$) and short stepsizes ($\gamma_{t}=10^{-7}/(t+1)^{3/4}$, $\rho_{t}=10^{3}(t+1)^{1/4}$, $r=0.5$), and the stochastic variant of the dynamic barrier gradient descent (SDBGD) [20] with long ($\gamma_{t}=10^{-2}$) and short ($\gamma_{t}=10^{-6}$) stepsizes.
-We initialized all algorithms with randomized starting points.
-For SBCGI and SBCGF, we generated the required initial point $x^{\prime}_{0}$ by running the SPIDER-FW algorithm [57, Algorithm 2] on the inner-level problem in (6), using a stepsize of $\gamma_{t}=0.1/(t+1)$.
-The initialization process terminated either after $10^{5}$ stochastic oracle queries or when the computation time exceeded 100 seconds, with the resulting point serving as $x_{0}$ for both SBCGI and SBCGF.
-The implementation of these algorithms rely on certain linear optimization or projection oracles.
-For SBCGI and SBCGF, a linear optimization oracle over the $\ell_{1}$-norm ball intersecting with a half-space is required, which we employ CVXPY [10] to solve the problems, similar to the implementation in [4, Appendix F.1].
-To compute the projection onto the feasible set required for aR-IP-SeG, we used [12, Algorithm 1].
-To approximate the outer- and inner-level optimal values $F_{\operatorname{opt}},G_{\operatorname{opt}}$, we again employ CVXPY to solve the inner-level problem and the reformulation (2) of the bilevel problem.
-All implementation details follow those in the corresponding papers.
+We showcase the performance of our proposed algorithms in an over-parametrized regression and a dictionary learning problems.
+For performance comparison, we implement four algorithms: SBCGI [4, Algorithm 1], SBCGF [4, Algorithm 2], aR-IP-SeG [27], and the stochastic variant of the dynamic barrier gradient descent (SDBGD) [21].
+All implementation details follow those in the corresponding papers and available in Appendix˜D.
+To ensure reproducibility, all source codes are made available at https://github.com/brucegiang/CG-StoBilvl.
 4.1 Over-parameterized regression
-We study the following bilevel optimization problem
-$\begin{array}[]{cl}\displaystyle\min_{x\in{\mathbb{R}}^{d}}&\displaystyle F(x):=\frac{1}{|\mathcal{D}_{\text{val}}|}\|A_{\text{val}}x-b_{\text{val}}\|_{2}^{2}\\ \text{s.t.}&\displaystyle x\in\operatorname*{arg\,min}_{\|z\|_{1}\leq\delta}G(z):=\frac{1}{|\mathcal{D}_{\text{tr}}|}\|A_{\text{tr}}z-b_{\text{tr}}\|_{2}^{2},\end{array}$ (6)
+We first consider a simple bilevel optimization problem with a convex outer-level objective function
+$\begin{array}[]{cl}\displaystyle\min_{x\in{\mathbf{R}}^{d}}\displaystyle F(x):=\frac{1}{|\mathcal{D}_{\text{val}}|}\|A_{\text{val}}x-b_{\text{val}}\|_{2}^{2}\quad\text{s.t.}\quad\displaystyle x\in\operatorname*{arg\,min}_{\|z\|_{1}\leq\delta}G(z):=\frac{1}{|\mathcal{D}_{\text{tr}}|}\|A_{\text{tr}}z-b_{\text{tr}}\|_{2}^{2},\end{array}$ (6)
 where the goal is to minimize the validation loss by choosing among optimal solutions of the training loss constrained by the $\ell_{1}$-norm ball.
@@
 This dataset consists of $n=1068$ samples and $d=730$ features.
-Figure 1 shows the convergence behavior of different algorithms for the inner-level (left) and outer-level (right) optimality gaps in terms of execution time.
+We evaluate the performance of different algorithms across $10$ experiments with random initializations and report their average performance within a $4$-minute execution limit.
+Figure˜1 shows the convergence behavior of different algorithms for the inner-level (left) and outer-level (right) optimality gaps in terms of execution time.
 Our proposed stochastic methods outperform existing approaches.
-In particular, IR-FSCG achieves high-accuracy solutions for the outer-level objective while maintaining strong inner-level feasibility, surpassing its mini-batch counterpart, SBCGI.
+In particular, IR-FSCG achieves high-accuracy solutions for the outer-level objective while maintaining strong inner-level feasibility, surpassing its mini-batch counterpart, SBCGF.
 Similarly, IR-SCG outperforms its single-sample counterpart, SBCGI.
@@
 Furthermore, neither aR-IP-SeG nor SDBGD makes significant progress in optimizing the outer-level objective function.
-Finally, the table below reports the average number of iterations and stochastic oracle calls over the 4-minute time limit for each algorithm.
+The table below also reports the average number of iterations and stochastic oracle calls over the 4-minute time limit for each algorithm.
 Our projection-free method completes more iterations within the fixed time budget, benefiting from a simpler linear minimization oracle compared to the projection-free approach in [4].
@@
 SDBGD (long step) $1243720.1$ $2487440.2$
-Figure 1: The inner-level optimality gap (left) and the outer-level absolute optimality gap (right) over time for different algorithms on the over-parametrized regression problem Eq. 6.
-3.2 Minimal norm logistic regression
-Pedregosa et al. [39] studied the following $\ell_{1}$-constrained logistic regression problem
-$\operatorname*{arg\,min}_{\|x\|_{1}\leq\beta}\frac{1}{n}\sum_{i\in[n]}\psi(a_{i}^{\top}x,b_{i})+\frac{\lambda}{2}\|x\|_{2}^{2},$ (7)
-where $\psi$ is the logistic loss function.
-In this section we consider the bilevel variant of problem (7), in which we seek the minimal $\ell_{2}$-norm solution to $\ell_{1}$-constrained logistic regression problem, that is, we aim to solve the following bilevel optimization problem
-$\min_{x\in{\mathbf{R}}^{d}}$ $\quad F(x):=\frac{1}{2}\|x\|_{2}^{2}$ (8)
-s.t. $\quad x\in\operatorname*{arg\,min}_{\|z\|_{1}\leq\beta}G(z):=\frac{1}{n}\sum_{i\in[n]}\psi(a_{i}^{\top}z,b_{i}).$
-We note that here we treat the outer-level as deterministic.
-Therefore, the randomness in the implemented algorithms is inherited from the solely the inner-level problem.
-We evaluated the method on the Madelon dataset [21], which consists of a data matrix $A\in\mathbb{R}^{n\times d}$ with $n=2000$ samples and $d=500$ features, along with an output vector $b\in\mathbb{R}^{n}$.
-We set $\beta=20$ for all experiments.
-Figure 2 illustrates the convergence behavior of various algorithms in terms of execution time, showing the optimality gaps for the inner-level (left) and outer-level (right) problems.
-The table below reports the average number of iterations and stochastic oracle calls over the 4-minute time limit for each algorithm.
-The overall performance trends are consistent with the previous experiment.
-However, all baseline stochastic methods fail to make meaningful progress within the fixed time limit.
-In contrast, our proposed algorithms, in particular IR-SCG, demonstrate strong performance on both the lower- and upper-level problems.
+Figure 1: The inner-level optimality gap (left) and the outer-level absolute optimality gap (right) over time for different algorithms on the over-parametrized regression problem Eq.˜6.
+4.2 Dictionary learning
+In the dictionary learning problem, the goal is to learn a compact representation of the input data $A=\{a_{1},\dots,a_{n}\}\subseteq{\mathbf{R}}^{m}$.
+Formally, we aim to find a dictionary $D=[d_{1}\cdots d_{p}]\in{\mathbf{R}}^{m\times p}$ such that each data point $a_{i}$ can be approximated by a linear combination of the basis vectors in $D$.
+This leads to the following optimization problem
+$\begin{aligned} \min_{D,X}&\quad\frac{1}{2n}\sum_{i\in[n]}\|a_{i}-Dx_{i}\|_{2}^{2}\\ \text{s.t.}~&\quad D\in{\mathbf{R}}^{m\times p},\ X=[x_{1}\cdots x_{n}]\in{\mathbf{R}}^{p\times n}\\ &\quad\|d_{j}\|_{2}\leq 1,\quad\forall j\in[p],\quad\|x_{i}\|_{1}\leq\delta,\quad\forall i\in[n],\end{aligned}$
+where we refer to $X$ as the coefficient matrix.
+In practice, data points usually arrive sequentially and the representation evolves gradually.
+Hence, the dictionary must be updated sequentially as well.
+Assume that we already have learned a dictionary $\hat{D}\in{\mathbf{R}}^{m\times p}$ and the corresponding coefficient matrix $\hat{X}\in{\mathbf{R}}^{p\times n}$ for some data set $A$.
+As a new dataset $A^{\prime}=\{a_{1}^{\prime},\dots,a_{n^{\prime}}^{\prime}\}\subseteq{\mathbf{R}}^{m}$ arrives, we intend to enrich our dictionary by learning $\tilde{D}\in{\mathbf{R}}^{m\times q}$ ($q>p$) and the coefficient matrix $\tilde{X}\in{\mathbf{R}}^{q\times n^{\prime}}$ for the new dataset while maintaining good performance of $\tilde{D}$ on the old dataset $A$ as well as the learned coefficient matrix $\hat{X}$.
+This leads to the following simple bilevel problem with a nonconvex outer-level objective function
+$\min_{\tilde{D},\tilde{X}}$ $\quad F(\tilde{D},\tilde{X})=\frac{1}{2n^{\prime}}\sum_{i\in[n^{\prime}]}\|a_{i}^{\prime}-\tilde{D}\tilde{x}_{i}\|_{2}^{2}$ (7)
+s.t. $\quad\tilde{D}\in{\mathbf{R}}^{m\times q},\ \tilde{X}=[\tilde{x}_{1}\cdots\tilde{x}_{n}]\in{\mathbf{R}}^{q\times n^{\prime}},\quad\|\tilde{x}_{i}\|_{1}\leq\delta,\quad\forall i\in[n^{\prime}],$
+$\quad\tilde{D}\in\operatorname*{arg\,min}_{\|z_{j}\|_{2}\leq 1,\forall j\in[q]}g(Z)=\frac{1}{2n}\sum_{i\in[n]}\|a_{i}-Z\hat{x}_{i}\|_{2}^{2},$
+where we denote $\hat{x}_{i}$ as the prolonged vector in ${\mathbf{R}}^{q}$ by appending zeros at the end. We consider problem (7) on a synthetic dataset with a similar setup to [29, 4].
+We evaluate the performance of different algorithms across $10$ experiments with random initializations and report their average performance within a $1$-minute execution limit.
+Figure˜2 shows the convergence behavior of different algorithms for the inner-level (left) and outer-level (right) problems in terms of execution time.
+Our proposed stochastic methods outperform existing approaches.
+In particular, IR-FSCG achieves better solutions for the outer-level objective while maintaining strong inner-level feasibility, surpassing its mini-batch counterpart, SBCGI.
+Similarly, IR-SCG outperforms its single-sample counterpart, SBCGI.
+We again observe that both SBCGI and SBCGF exhibit degradation in their inner-level optimality gaps over time.
+The table below reports the average number of iterations and stochastic oracle calls over the $1$-minute time limit for each algorithm.
+Our projection-free method completes more iterations within the fixed time budget, benefiting from a simpler linear minimization oracle compared to the projection-free methods in [4].
+We also observe that SDBGD performs fewer iterations than IR-SCG, as it relies on a more costly projection oracle.
+Figure 2: The inner-level optimality gap (left) and the outer-level objective function (right) over time for different algorithms on the dictionary learning problem (7).
 Method Average # of iterations Average # of oracle calls
-IR-SCG $1726604.0$ $1726604.0$
-IR-FSCG $225263.8$ $19926327.2$
-SBCGI $6207.8$ $8852217.8$
-SBCGF $5986.5$ $9376410.0$
-aR-IP-SeG (short step) $1218831.0$ $2437662.0$
-aR-IP-SeG (long step) $973268.7$ $1946537.4$
-SDBGD (short step) $1623034.9$ $1623034.9$
-SDBGD (long step) $1618114.5$ $1618114.5$
-Figure 2: The inner-level optimality gap (left) and the outer-level absolute optimality gap (right) over time for different algorithms on the minimal norm logistic regression problem Eq. 8.
-Appendix A Convergence proofs
-This section provides the proofs of Theorems 1 and 2.
-We employ the notation $\Phi_{t}(x):=\sigma_{t}F(x)+G(x)$ and $\widehat{\nabla\Phi}_{t}:=\sigma_{t}\widehat{\nabla F}_{t}+\widehat{\nabla G}_{t}$ to denote the gradient estimate at step $t$.
+IR-SCG $32736.1$ $65472.2$
+IR-FSCG $30532.6$ $1872851.0$
+SBCGI $466.3$ $7440.6$
+SBCGF $361.84$ $6147916.0$
+aR-IP-SeG (short step) $6290.0$ $25160.0$
+aR-IP-SeG (long step) $6278.1$ $25112.4$
+SDBGD (short step) $11441.3$ $22882.6$
+SDBGD (long step) $12239.3$ $24478.6$
+Appendix A Preparatory Lemmas
+Throughout all appendices, we employ the notation $\Phi_{t}(x):=\sigma_{t}F(x)+G(x)$ and $\widehat{\nabla\Phi}_{t}:=\sigma_{t}\widehat{\nabla F}_{t}+\widehat{\nabla G}_{t}$ to denote the gradient estimate at step $t$.
+We also denote $x_{\operatorname{opt}}\in X_{\operatorname{opt}}$ as an optimal solution to (1).
 All norms refer to the Euclidean norm.
-A.1 Preparatory lemmas
+For any sequence $\{x_{t}\}_{t\geq 0}$ generated by either Algorithm˜1 or Algorithm˜2, we observe that by construction, $x_{0},v_{t}\in X$ .
+Hence, under convexity of $X$, $x_{t+1}\in X$ as it is a convex combination of $x_{t}$ and $v_{t}$, and $X$ is assumed to be convex.
+By a simple induction argument, one can easily show that the sequence $\{x_{t}\}_{t\geq 0}\in X$.
+A.1 Convex outer-level
 Lemma 1.
-Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm 1 or Algorithm 2 with some given stepsizes $\{\alpha_{t}\}_{t\geq 0}$. If 1 holds, then for every $t\geq k$, we have
-$\Phi_{t}(x_{t+1})-\Phi_{t}(x_{\operatorname{opt}})\!\leq\!(1-\alpha_{t})(\Phi_{t}(x_{t})-\Phi_{t}(x_{\operatorname{opt}}))+\frac{(L_{f}\sigma_{t}\!+\!L_{g})D^{2}\alpha_{t}^{2}}{2}+2\alpha_{t}D\left\|\nabla\Phi_{t}(x_{t})\!-\!\widehat{\nabla\Phi}_{t}\right\|.$
-Proof of Lemma 1.
+Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm˜1 or Algorithm˜2 with some given stepsizes $\{\alpha_{t}\}_{t\geq 0}$. If ˜1 holds, then for every $t\geq k$, we have
+$\Phi_{t}(x_{t+1})-\Phi_{t}(x_{\operatorname{opt}})\!\leq\!(1-\alpha_{t})(\Phi_{t}(x_{t})-\Phi_{t}(x_{\operatorname{opt}}))+\frac{(L_{f}\sigma_{t}\!+\!L_{g})D^{2}\alpha_{t}^{2}}{2}+2\alpha_{t}D\left\|\nabla\Phi_{t}(x_{t})\!-\!\widehat{\nabla\Phi}_{t}\right\|,$
+Proof of Lemma˜1.
 Recall that
-$x_{t+1}=x_{t}+\alpha_{t}(v_{t}-x_{t}),\quad\text{where}\quad v_{t}\in\operatorname*{arg\,min}_{v\in X}v^{\top}\widehat{\nabla\Phi}_{t}.$
+$x_{t+1}=x_{t}+\alpha_{t}(v_{t}-x_{t}),\quad\text{where}\quad v_{t}\in\operatorname*{arg\,min}_{v\in X}\left\{\left(\widehat{\nabla\Phi}_{t}\right)^{\top}v\right\}.$
 As both functions $F$ and $G$ are assumed to be convex and smooth, it is easy to show that $\Phi$ is convex and $\sigma_{t}L_{f}+L_{g}$-smooth.
@@
 $\Phi_{t}(x_{t+1})\!\leq\!\Phi_{t}(x_{t})\!+\!\alpha_{t}\nabla\Phi_{t}(x_{t})^{\top}(w_{t}\!-\!x_{t})\!+\!2D\alpha_{t}\|\widehat{\nabla\Phi}_{t}\!-\!\nabla\Phi_{t}(x_{t})\|+\frac{(\sigma_{t}L_{f}\!+\!L_{g})D^{2}\alpha_{t}^{2}}{2}.$ (8)
-Let $x_{\operatorname{opt}}\in X_{\operatorname{opt}}$, and set $w_{t}=x_{\operatorname{opt}}$.
+and set $w_{t}=x_{\operatorname{opt}}$.
 Thanks to the convexity of $\Phi_{t}$, one can easily show that $\nabla\Phi_{t}(x_{t})^{\top}(x_{\operatorname{opt}}-x_{t})\leq\Phi_{t}(x_{\operatorname{opt}})-\Phi_{t}(x_{t})$.
@@
 Lemma 2.
-Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm 1 or Algorithm 2 with stepsizes $\{\alpha_{t}\}_{t\geq 0}$ given by $\alpha_{t}=2/(t+2)$ for any $t\geq k$.
-If 1 holds, then for each $t\geq k$, we have
+Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm˜1 or Algorithm˜2 with stepsizes $\{\alpha_{t}\}_{t\geq 0}$ given by $\alpha_{t}=2/(t+2)$ for any $t\geq k$.
+If ˜1 holds, then for each $t\geq k$, we have
 $t(t+1)\left(\Phi_{t}(x_{t})-\Phi_{t}(x_{\operatorname{opt}})\right)$ $+\sum_{i\in[t]\setminus[k]}(i+1)i(\sigma_{i-1}-\sigma_{i})(F(x_{i})-F_{\operatorname{opt}})$ (9)
@@
 $\quad+\sum_{i\in[t]\setminus[k]}4iD\|\widehat{\nabla\Phi}_{i-1}-\nabla\Phi_{i-1}(x_{i-1})\|.$
-Proof of Lemma 2.
+Proof of Lemma˜2.
 By definition of $\Phi_{t}$, we have
@@
 $\eta_{t}:=\frac{2(L_{f}\sigma_{t}+L_{g})D^{2}}{(t+2)^{2}}+\frac{4D}{t+2}\|\widehat{\nabla\Phi}_{t}-\nabla\Phi_{t}(x_{t})\|-(\sigma_{t}-\sigma_{t+1})(F(x_{t+1})-F_{\operatorname{opt}}).$
-Applying Lemma 1 with stepsize $\alpha_{t}=2/(t+2)$ for all $t\geq k$ gives the recursion $h_{t+1}\leq\frac{t}{t+2}h_{t}+\eta_{t}$ for any $t\geq k$. This can be re-expressed as
+Applying Lemma˜1 with stepsize $\alpha_{t}=2/(t+2)$ for all $t\geq k$ gives the recursion $h_{t+1}\leq\frac{t}{t+2}h_{t}+\eta_{t}$ for any $t\geq k$. This can be re-expressed as
 $(t+1)th_{t}\leq t(t-1)h_{t-1}+t(t+1)\eta_{t-1}$
@@
 Proposition 1.
-Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm 1 or Algorithm 2 with stepsizes $\{\alpha_{t}\}_{t\geq 0}$ given by $\alpha_{t}=2/(t+2)$ for any $t\geq k$.
-If 1 holds, then for each $t\geq k$, we have
-[…]
+Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm˜1 or Algorithm˜2 with stepsizes $\{\alpha_{t}\}_{t\geq 0}$ given by $\alpha_{t}=2/(t+2)$ for any $t\geq k$.
+If ˜1 holds, then for each $t\geq k$, we have
+$\begin{aligned} G(x_{t})-G_{\operatorname{opt}}&\leq\frac{(k+1)k\left(\Phi_{k}(x_{k})-\Phi_{k}(x_{\operatorname{opt}})\right)}{t(t+1)}+\frac{2(\sigma_{0}L_{f}+L_{g})D^{2}(t-k)}{t(t+1)}\\[4.30554pt] &\quad+\sum_{i\in[t]\setminus[k]}\frac{4iD\|\widehat{\nabla\Phi}_{i-1}-\nabla\Phi_{i-1}(x_{i-1})\|}{t(t+1)}\\[4.30554pt] &\quad+\left(F_{\operatorname{opt}}-\min_{x\in X}F(x)\right)\left(\sigma_{t}+\frac{\sum_{i\in[t]}(i+1)i(\sigma_{i-1}-\sigma_{i})}{t(t+1)}\right),\end{aligned}$ (10)
 and
@@
 $\quad+\sum_{i\in[t]\setminus[k]}4iD\|\widehat{\nabla\Phi}_{i-1}-\nabla\Phi_{i-1}(x_{i-1})\|.$
-Proof of Proposition 1.
-By definition of the function $\Phi_{t}$, the first term in Eq. 10 satisfies
+Proof of Proposition˜1.
+By definition of the function $\Phi_{t}$, the first term in Eq.˜9 satisfies
 $t(t\!+\!1)\left(\Phi_{t}(x_{t})\!-\!\Phi_{t}(x_{\operatorname{opt}})\right)\!=\!t(t\!+\!1)\left(G(x_{t})\!-\!G(x_{\operatorname{opt}})\right)\!+\!t(t\!+\!1)\sigma_{t}\left(F(x_{t})\!-\!F(x_{\operatorname{opt}})\right).$ (12)
-By construction, $x_{0},v_{t}\in X$ Hence, $x_{t+1}\in X$ as it is a convex combination of $x_{t}$ and $v_{t}$, and $X$ is assumed to be convex.
-By a simple induction argument, one can easily show that the sequence $\{x_{t}\}_{t\geq 0}\in X$.
-Therefore, we have $\min_{x\in X}F(x)-F_{\operatorname{opt}}\leq F(x_{t})-F_{\operatorname{opt}}$ for all $t\geq 0$.
-Using this bound together with the identity in Eq.
-13, we can further lower bound the left-hand side of inequality (10), and arrive at the bound (11).
+Since $x_{t}\in X$, we have $\min_{x\in X}F(x)-F_{\operatorname{opt}}\leq F(x_{t})-F_{\operatorname{opt}}$ for all $t\geq 0$.
+Using this bound together with the identity in Eq.˜12, we can further lower bound the left-hand side of inequality (9), and arrive at the bound (10).
 This completes the proof of the first claim.
 For the second claim, since $\{x_{t}\}_{t\geq 0}\in X$ is feasible in the lower-level problem, it follows directly that $G(x_{t})\geq G_{\operatorname{opt}}$.
-Combining this inequality with (13) and  (10), we obtain Eq.
-12, which concludes the proof.
+Combining this inequality with (12) and  (9), we obtain Eq.˜11, which concludes the proof.
 ∎
-A.2 Proof of Theorem 1
-The proof proceeds by applying Proposition 1 with $k=0$.
-In this case, the main random component in Eq.
-11 and Eq.
-12 is the term $\sum_{i\in[t]}4iD\|\widehat{\nabla\Phi}_{i-1}-\nabla\Phi_{i-1}(x_{i-1})\|$.
-We first derive a probabilistic upper bound for this term, then carefully set the parameters in Algorithm 1 to establish the convergence rate for both the lower- and upper-level problems.
-We begin by bounding the gradient estimator used in Algorithm 1.
+A.2 Nonconvex outer-level
+When $F$ is possibly nonconvex, we start by establishing a bound, akin to Lemma 2.
 Lemma 3.
-Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm 1 with stepsize $\alpha_{t}=2/(t+2)$ for any $t\geq 0$.
-If 1 holds, then for any $t\geq 1$, given $\delta\in(0,1)$, with probability at least $1-\delta$, for some absolute constant $c$, it (jointly) holds that
+Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm˜1 or Algorithm˜2 with $\{\alpha_{t},\sigma_{t}\}_{t\geq 0}$ given by $\alpha_{t}=(t+1)^{-\omega}$ and $\sigma_{t}=\varsigma(t+1)^{-p}$ for $t\geq k$, with $0<p\leq\omega$.
+If ˜1 holds with the exception that $F$ may be nonconvex, then for every $t\geq k$, we have
+$(t+1)^{\omega}\mathcal{G}(x_{t+1})-k^{\omega}\mathcal{G}(x_{k})$ $\leq 2\varsigma\sup_{z\in X}|F(z)|(t+1)^{\omega-p}+\sigma_{k}F(x_{k})-\sum_{i=k}^{t}\sigma_{i}\mathcal{F}(x_{i})$
+$\quad+\sum_{i=k}^{t}\left(2D\left\|\widehat{\nabla\Phi}_{i}-\nabla\Phi_{i}(x_{i})\right\|+\frac{(\sigma_{i}L_{f}+L_{g})D^{2}}{2(i+1)^{\omega}}\right).$
+Proof of Lemma˜3.
+Using the definition of $\Phi_{t}=\sigma_{t}F+G$, for any $t\geq k$, we have
+$\sigma_{t}F(x_{t+1})+G(x_{t+1})$ $\leq\sigma_{t}F(x_{t})+G(x_{t})+\alpha_{t}\min_{v\in X}(\sigma_{t}\nabla F(x_{t})+\nabla G(x_{t}))^{\top}(v-x_{t})$
+$\quad+2D\alpha_{t}\|\widehat{\nabla\Phi}_{t}-\nabla\Phi_{t}(x_{t})\|+\frac{(\sigma_{t}L_{f}+L_{g})D^{2}\alpha_{t}^{2}}{2}$
+where the inequality follows directly from (8) and by minimizing over $w_{t}\in X$.
+We emphasize that (8) does not involve any convexity assumptions on $\Phi_{t}$ and relies solely on its smoothness.
+Furthermore, notice that for any $x\in X$,
+$\min_{v\in X}~(\sigma\nabla F(x)+\nabla G(x))^{\top}(v-x)$ $\leq\min_{v\in X_{\operatorname{opt}}}~(\sigma\nabla F(x)+\nabla G(x))^{\top}(v-x)$
+$\leq\min_{v\in X_{\operatorname{opt}}}~\left\{\sigma\nabla F(x)^{\top}(v-x)+G(v)-G(x)\right\}$
+$=-\sigma\mathcal{F}(x)-\mathcal{G}(x),$
+where the last equality follows from the definition of $\mathcal{F}$ and $\mathcal{G}$. Using the above bound, we may thus conclude that
+$\sigma_{t}F(x_{t+1})+G(x_{t+1})$ $\leq\sigma_{t}F(x_{t})+G(x_{t})-\alpha_{t}(\sigma_{t}\mathcal{F}(x_{t})+\mathcal{G}(x_{t}))$
+$\quad+2D\alpha_{t}\|\widehat{\nabla\Phi}_{t}-\nabla\Phi_{t}(x_{t})\|+\frac{(\sigma_{t}L_{f}+L_{g})D^{2}\alpha_{t}^{2}}{2}.$
+A simple re-arrangement and using the definition of $\mathcal{G}$ then yields the bound
+$\begin{aligned} \alpha_{t}(\sigma_{t}\mathcal{F}(x_{t})+\mathcal{G}(x_{t}))&\leq\sigma_{t}(F(x_{t})-F(x_{t+1}))+\mathcal{G}(x_{t})-\mathcal{G}(x_{t+1})\\ &\quad+2D\alpha_{t}\|\widehat{\nabla\Phi}_{t}-\nabla\Phi_{t}(x_{t})\|+\frac{(\sigma_{t}L_{f}+L_{g})D^{2}\alpha_{t}^{2}}{2}.\end{aligned}$ (13)
+Using the definition of $\alpha_{t}$ and the fact that $(t+1)^{\omega}\leq t^{\omega}-1$ for any $t\geq 0$, we arrive at
+$(t+1)^{\omega}\mathcal{G}(x_{t+1})$ $\leq t^{\omega}\mathcal{G}(x_{t})+(t+1)^{\omega}\sigma_{t}(F(x_{t})-F(x_{t+1}))-\sigma_{t}\mathcal{F}(x_{t})$
+$\quad+2D\|\widehat{\nabla\Phi}_{t}-\nabla\Phi_{t}(x_{t})\|+\frac{(\sigma_{t}L_{f}+L_{g})D^{2}}{2(t+1)^{\omega}}.$
+Unwinding this recursion back to $i=k$, we obtain
+$(t+1)^{\omega}\mathcal{G}(x_{t+1})$ $\leq k^{\omega}\mathcal{G}(x_{k})+\sum_{i=k}^{t}\left((i+1)^{\omega}\sigma_{i}(F(x_{i})-F(x_{i+1}))\right)-\sum_{i=k}^{t}\sigma_{i}\mathcal{F}(x_{i})$
+$\quad+\sum_{i=k}^{t}\left(2D\|\widehat{\nabla\Phi}_{i}-\nabla\Phi_{i}(x_{i})\|+\frac{(\sigma_{i}L_{f}+L_{g})D^{2}}{2(i+1)^{\omega}}\right).$
+Since $p\leq\omega$, we have $(i+1)^{\omega}\sigma_{i}-i^{\omega}\sigma_{i-1}\geq 0$. Therefore, we deduce that
+$\sum_{i=k}^{t}(i+1)^{\omega}\sigma_{i}(F(x_{i})-F(x_{i+1}))$
+$=\sum_{i=k}^{t}\Big(F(x_{i})((i+1)^{\omega}\sigma_{i}-i^{\omega}\sigma_{i-1})\Big)+\sigma_{k}F(x_{k})-(t+1)^{\omega}\sigma_{t}F(x_{t+1})$
+$\leq\sup_{z\in X}|F(z)|\sum_{i=k}^{t}\Big(((i+1)^{\omega}\sigma_{i}-i^{\omega}\sigma_{i-1})\Big)+\sigma_{k}F(x_{k})+(t+1)^{\omega}\sigma_{t}\sup_{z\in X}|F(z)|$
+$\leq 2\varsigma\sup_{z\in X}|F(z)|(t+1)^{\omega-p}+\sigma_{k}F(x_{k}).$
+This completes the proof. ∎
+[…]
+Lemma 4.
+If ˜1 holds, the function $\mathcal{F}$ as defined in (3) is $L_{f}D+\max_{z\in X}\|\nabla F(z)\|$-Lipschitz continuous over $X$.
+Proof of Lemma˜4.
+Given $x,y\in X$, we observe that
+$\mathcal{F}(y)$ $=\max_{v\in X_{\operatorname{opt}}}\left\{\nabla F(y)^{\top}(y-v)\right\}$
+$=\max_{v\in X_{\operatorname{opt}}}\{(\nabla F(y)-\nabla F(x))^{\top}(y-v)+\nabla F(x)^{\top}(y-x)+\nabla F(x)^{\top}(x-v)\}$
+$=\max_{v\in X_{\operatorname{opt}}}\{(\nabla F(y)-\nabla F(x))^{\top}(y-v)+\nabla F(x)^{\top}(y-x)+\nabla F(x)^{\top}(x-v)\}$
+$\leq\max_{v\in X_{\operatorname{opt}}}\{\|\nabla F(y)-\nabla F(x)\|\|y-v\|\!+\nabla F(x)^{\top}(y-x)+\nabla F(x)^{\top}(x-v)\}$
+$\leq\max_{v\in X_{\operatorname{opt}}}\{L_{f}\|x-y\|D+\nabla F(x)^{\top}(y-x)+\nabla F(x)^{\top}(x-v)\}$
+$=L_{f}D\|x-y\|+\nabla F(x)^{\top}(y-x)+\mathcal{F}(x)$
+$\leq\left(L_{f}D+\max_{z\in X}\|\nabla F(z)\|\right)\|y-x\|+\mathcal{F}(x),$
+where the first and third inequality follows from Cauchy-Schwartz inequality, while the second inequality follows from the $L_{f}$-smoothness of $F$.
+By interchanging the role of $y$ and $x$, we deduce that
+$|\mathcal{F}(y)\!-\!\mathcal{F}(x)|\leq\left(L_{f}D+\max_{z\in X}\|\nabla F(z)\|\right)\|y-x\|.$
+This concludes the proof. ∎
+Appendix B Proofs of Section 2
+B.1 Proof of Theorem˜1
+We restate Theorem˜1 with exact upper bounds on sub-optimality gaps.
+Theorem 1.
+Let $\{z_{t}\}_{t\geq 1}$ be the iterates generated by Algorithm˜1 with $\alpha_{t}=2/(t+2)$ for any $t\geq 0$, and regularization parameters $\{\sigma_{t}\}_{t\geq 0}$ given by $\sigma_{t}:=\varsigma(t+1)^{-p}$ for some chosen $\varsigma>0$, $p\in(0,1/2)$.
+Under ˜1, for any $t\geq 1$, with probability at least $1-\delta$, we have
+$F(z_{t})-F_{\text{opt}}\leq\frac{2(L_{f}+L_{g}/\varsigma)D^{2}}{(t+1)^{1-p}}+\frac{4c\left(4(L_{f}+L_{g}/\varsigma)D+3(\sigma_{f}+\sigma_{g}/\varsigma)\right)}{(t+1)^{1/2-p}}\sqrt{\log\left({8dt^{2}}/{\delta}\right)},$
+and
+$G(z_{t})-G_{\text{opt}}$
+$\leq\left(1+\frac{2p}{\min\{1,2(1-p)\}}\right)\left[(1+2p)\left(F_{\text{opt}}-\min_{x\in X}F(x)\right)\varsigma+2(\varsigma L_{f}+L_{g})D^{2}\right$.
+$\quad\left.+4c\left(8\left(L_{f}\varsigma+L_{g}\right)D^{2}+3\left(\sigma_{f}\varsigma+\sigma_{g}\right)D\right)\sqrt{\frac{2}{(1-2p)e}+\log\left({8d}/{\delta}\right)}\,\right](t+1)^{-p},$
+where $c$ is some absolute constant. Moreover, with probability $1$, we have
+$\lim_{t\to\infty}F(x_{t})=F_{\operatorname{opt}},\quad\lim_{t\to\infty}G(x_{t})=G_{\operatorname{opt}}.$
+The proof proceeds by applying Proposition˜1 with $k=0$.
+In this case, the main random component in Eq.˜10 and Eq.˜11 is the term $\sum_{i\in[t]}4iD\|\widehat{\nabla\Phi}_{i-1}-\nabla\Phi_{i-1}(x_{i-1})\|$.
+We first derive a probabilistic upper bound for this term, then carefully set the parameters in Algorithm˜1 to establish the convergence rate for both the lower- and upper-level problems.
+We begin by bounding the gradient estimator used in Algorithm˜1.
+Lemma 5.
+Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm˜1 with stepsize $\alpha_{t}=2/(t+2)$ for any $t\geq 0$.
+If ˜1 holds, then for any $t\geq 1$, given $\delta\in(0,1)$, with probability at least $1-\delta$, for some absolute constant $c$, it (jointly) holds that
 $\|\widehat{\nabla F}_{t}-\nabla F(x_{t})\|$ $\leq c\sqrt{2}\left(4L_{f}D+3\sigma_{f}\right)\frac{\sqrt{\log(4d/\delta)}}{(t+1)^{1/2}}$ (14)
 $\|\widehat{\nabla G}_{t}-\nabla G(x_{t})\|$ $\leq c\sqrt{2}\left(4L_{g}D+3\sigma_{g}\right)\frac{\sqrt{\log(4d/\delta)}}{(t+1)^{1/2}}.$
-Proof of Lemma 3.
+Proof of Lemma˜5.
 The proof closely follows the argument in [4, Lemma 4.1].
@@
 Proposition 2.
-Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm 1 with $\alpha_{t}=2/(t+2)$ for any $t\geq 0$.
-If 1 holds and the parameters $\{\sigma_{t}\}_{t\geq 0}$ are non-increasing and positive, then given $\delta\in(0,1)$, with probability at least $1-\delta$, for some absolute constant $c$ and any $t\geq 1$, we have
-$\sum_{i\in[t]}i\|\widehat{\nabla\Phi}_{i-1}-\nabla\Phi_{i-1}(x_{i-1})\|\leq c\left(4(\sigma_{0}L_{f}+L_{g})D+3(\sigma_{0}\sigma_{f}+\sigma_{g})\right)(t+1)^{3/2}\sqrt{\log\left({8dt^{2}}/{\delta}\right)}.$
-Proof of Proposition 2.
-Given $i\geq 1$ and $\delta\in(0,1)$, by Lemma 3, with probability at least $1-\delta/i(i+1)$, we have
+Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm˜1 with $\alpha_{t}=2/(t+2)$ for any $t\geq 0$.
+If ˜1 holds and the parameters $\{\sigma_{t}\}_{t\geq 0}$ are non-increasing and positive, then given $\delta\in(0,1)$, with probability at least $1-\delta$, for some absolute constant $c$ and any $t\geq 1$, we have
+$\sum_{i\in[t]}i\|\widehat{\nabla\Phi}_{i-1}-\nabla\Phi_{i-1}(x_{i-1})\|$
+$\leq c\left(4(\sigma_{0}L_{f}+L_{g})D+3(\sigma_{0}\sigma_{f}+\sigma_{g})\right)(t+1)^{3/2}\sqrt{\log\left({8dt^{2}}/{\delta}\right)}.$
+Proof of Proposition˜2.
+Given $i\geq 1$ and $\delta\in(0,1)$, by Lemma˜5, with probability at least $1-\delta/i(i+1)$, we have
 $i\|\widehat{\nabla\Phi}_{i-1}-\nabla\Phi_{i-1}(x_{i-1})\|$ $\leq i(\sigma_{i}\|\widehat{\nabla F}_{i-1}-\nabla F(x_{i-1})\|+\|\widehat{\nabla G}_{i-1}-\nabla G(x_{i-1})\|)$
@@
 where the second inequality follows from bounding $\sum_{i\in[t]}\sqrt{i}\leq\tfrac{2}{3}(t+1)^{3/2}$ via the Riemann sum approximation, and observing that $2\sqrt{2}\leq 3$. ∎
-To establish the desired convergence results for Algorithm 1, we need to impose certain conditions on the regularization parameters $\{\sigma_{t}\}_{t\geq 0}$, stated below.
+To establish the desired convergence results for Algorithm˜1, we need to impose certain conditions on the regularization parameters $\{\sigma_{t}\}_{t\geq 0}$, stated below.
 Condition 1.
-The regularization parameters $\{\sigma_{t}\}_{t\geq 0}$ are non-increasing, positive, and converge to 00.
+The regularization parameters $\{\sigma_{t}\}_{t\geq 0}$ are non-increasing, positive, and converge to 0.
 Condition 2.
@@
 $V$ $:=\sup_{t\geq 1}\left\{\frac{\sum_{i\in[t]}(i+1)i(\sigma_{i-1}-\sigma_{i})\sigma_{i}}{(t+1)t\sigma_{t}^{2}}\right\}.$ (15d)
-While it is not obvious a priori, Lemma 16 in Section B.4 guarantees that these quantities remain finite for any parameter choice satisfying 1 to 3.
-In particular, 1 to 3 hold when $\sigma_{t}=\varsigma(t+1)^{-p}$ for any $p\in(0,1)$, which is what is used in Theorems 1 and 2, and Lemma 6 provides explicit bounds for this case.
-We now analyze the sequence $\{x_{t}\}_{t\geq 0}$ in Algorithm 1 for the inner-level problem.
+While it is not obvious a priori, Lemma˜17 in Section˜D.1 guarantees that these quantities remain finite for any parameter choice satisfying 1 to 3.
+In particular, 1 to 3 hold when $\sigma_{t}=\varsigma(t+1)^{-p}$ for any $p\in(0,1)$, which is what is used in Theorems˜1 and 3, and Lemma˜8 provides explicit bounds for this case.
+We now analyze the sequence $\{x_{t}\}_{t\geq 0}$ in Algorithm˜1 for the inner-level problem.
 Lemma 6.
-Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm 1 with $\alpha_{t}=2/(t+2)$ for any $t\geq 0$, and let $C_{\delta}$ be defined as in (15a).
-If 1 and 1 to 3 hold, with probability at least $1-\delta$, it holds that
+Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm˜1 with $\alpha_{t}=2/(t+2)$ for any $t\geq 0$, and let $C_{\delta}$ be defined as in (15a).
+If ˜1 and 1 to 3 hold, with probability at least $1-\delta$, it holds that
 $G(x_{t})-G_{\operatorname{opt}}\leq C_{\delta}\sigma_{t},$ (16)
 for any $t\geq 0$.
-Proof of Lemma 4.
-Combining (11) from Proposition 1 with the probabilistic bound in Proposition 2, and after some straightforward calculations, with probability at least $1-\delta$, we have
+Proof of Lemma˜6.
+Combining (10) from Proposition˜1 with the probabilistic bound in Proposition˜2, and after some straightforward calculations, with probability at least $1-\delta$, we have
 $G(x_{t})-G_{\operatorname{opt}}$ $\leq\frac{8c\left(4(\sigma_{0}L_{f}+L_{g})D^{2}+3(\sigma_{0}\sigma_{f}+\sigma_{g})D\right)}{(t+1)^{1/2}}\sqrt{\log\left({8dt^{2}}/{\delta}\right)}$
@@
 By definition of $C_{\delta}$, the right-hand side is at most $C_{\delta}\sigma_{t}$ for any $t\geq 1$. This completes the proof. ∎
-We next analyze the convergence of the sequence $\{z_{t}\}_{t\geq 0}$ in Algorithm 1 for both outer- and inner-level problems.
+We next analyze the convergence of the sequence $\{z_{t}\}_{t\geq 0}$ in Algorithm˜1 for both outer- and inner-level problems.
 Lemma 7.
-Let $\{z_{t}\}_{t\geq 1}$ denote the iterates generated by Algorithm 1 with $\alpha_{t}=2/(t+2)$ for any $t\geq 0$, and let $C_{\delta},V$ be defined as in Eq.
-15.
-If 1 and 1 to 3 hold, then with probability at least $1-\delta$, it (jointly) holds that
+Let $\{z_{t}\}_{t\geq 1}$ denote the iterates generated by Algorithm˜1 with $\alpha_{t}=2/(t+2)$ for any $t\geq 0$, and let $C_{\delta},V$ be defined as in Eq.˜15.
+If ˜1 and 1 to 3 hold, then with probability at least $1-\delta$, it (jointly) holds that
 $F(z_{t})-F_{\operatorname{opt}}$ $\leq\frac{2(\sigma_{0}L_{f}+L_{g})D^{2}}{(t+1)\sigma_{t}}+\frac{4c\left(4(\sigma_{0}L_{f}+L_{g})D+3(\sigma_{0}\sigma_{f}+\sigma_{g})\right)}{(t+1)^{1/2}\sigma_{t}}\sqrt{\log\left({8dt^{2}}/{\delta}\right)},$
@@
 for any $t\geq 1$.
-Proof of Lemma 5.
-Recall from Algorithm 1 that we have defined
+Proof of Lemma˜7.
+Recall from Algorithm˜1 that we have defined
 $\begin{cases}S_{t+1}:=(t+2)(t+1)\sigma_{t+1}+\sum_{i\in[t+1]}(i+1)i(\sigma_{i-1}-\sigma_{i})\\[8.61108pt] z_{t+1}:=\frac{1}{S_{t+1}}\left((t+2)(t+1)\sigma_{t+1}x_{t+1}+\sum_{i\in[t+1]}(i+1)i(\sigma_{i-1}-\sigma_{i})x_{i}\right),\end{cases}$
@@
 $F(z_{t})-F_{\operatorname{opt}}$ $\leq\frac{2(\sigma_{0}L_{f}+L_{g})D^{2}t}{S_{t}}+\sum_{i\in[t]}\frac{4iD\|\widehat{\nabla\Phi}_{i-1}-\nabla\Phi_{i-1}(x_{i-1})\|}{S_{t}}.$
-Using the inequality $S_{t}\geq(t+1)t\sigma_{t}$, which holds thanks to 1, and applying Proposition 2, with probability at least $1-\delta$, we have
+Using the inequality $S_{t}\geq(t+1)t\sigma_{t}$, which holds thanks to ˜1, and applying Proposition˜2, with probability at least $1-\delta$, we have
 $F(z_{t})-F_{\operatorname{opt}}$ $\leq\frac{2(\sigma_{0}L_{f}+L_{g})D^{2}}{(t+1)\sigma_{t}}+\frac{8c\left(4(\sigma_{0}L_{f}+L_{g})D+3(\sigma_{0}\sigma_{f}+\sigma_{g})\right)}{(t+1)^{1/2}\sigma_{t}}\sqrt{\log\left({8dt^{2}}/{\delta}\right)}.$
@@
 The proof concludes by using the definition of $V$. ∎
-Lemma 5 establishes a convergence result in terms of the regularization parameters $\{\sigma_{t}\}_{t\geq 0}$.
-The next lemma specifies an update rule for these parameters and provides bounds on the quantities introduced in Eq.
-15.
+Lemma˜7 establishes a convergence result in terms of the regularization parameters $\{\sigma_{t}\}_{t\geq 0}$.
+The next lemma specifies an update rule for these parameters and provides bounds on the quantities introduced in Eq.˜15.
 Lemma 8.
-Consider the sequence $\sigma_{t}:=\varsigma(t+1)^{-p}$ for $t\geq 0$ and the quantities defined in Eq. 15.
+Consider the sequence $\sigma_{t}:=\varsigma(t+1)^{-p}$ for $t\geq 0$ and the quantities defined in Eq.˜15.
   • (i)
-    If $p\in(0,1)$, then $\{\sigma_{t}\}_{t\geq 0}$ satisfies 1 and 2 with $L=p$. Furthermore,
+    If $p\in(0,1)$, then $\{\sigma_{t}\}_{t\geq 0}$ satisfies ˜1 and ˜2 with $L=p$. Furthermore,
     $C\leq(1+2p)\left(F_{\operatorname{opt}}-\min_{x\in X}F(x)\right)+\frac{2(\varsigma L_{f}+L_{g})D^{2}}{\varsigma},\quad V\leq\frac{2p}{\min\{1,2(1-p)\}}.$
   • (ii)
-    If $p\in(0,1/2)$, then $\{\sigma_{t}\}_{t\geq 0}$ satisfies 3, and we have
-    $\bar{C}_{\delta}\leq 4c\left(8\left(L_{f}+\frac{L_{g}}{\varsigma}\right)D^{2}+3\left(\sigma_{f}+\frac{\sigma_{g}}{\varsigma}\right)D\right)\sqrt{\frac{2}{(1-2p)e}+\log\left(\frac{8d}{\delta}\right)}.$
-Proof of Lemma 6.
-As for assertion (i), it is trivial to see that the sequence $\{\sigma_{t}\}_{t\geq 0}$ satisfies Conditions 1. To validate 2, observe that
+    If $p\in(0,1/2)$, then $\{\sigma_{t}\}_{t\geq 0}$ satisfies ˜3, and we have
+    $\bar{C}_{\delta}\leq 4c\left(8\left(L_{f}+\frac{L_{g}}{\varsigma}\right)D^{2}+3\left(\sigma_{f}+\frac{\sigma_{g}}{\varsigma}\right)D\right)\sqrt{\frac{2}{(1-2p)e}+\log\left({8d}/{\delta}\right)}.$
+Proof of Lemma˜8.
+As for assertion (i), it is trivial to see that the sequence $\{\sigma_{t}\}_{t\geq 0}$ satisfies Conditions 1. To validate ˜2, observe that
 $\lim_{t\to\infty}t\left(\frac{\sigma_{t}}{\sigma_{t+1}}-1\right)=\lim_{t\to\infty}t\left(\left(1+\frac{1}{t+1}\right)^{p}-1\right)=\lim_{t\to\infty}\frac{t}{t+1}\cdot\lim_{\delta\to 0}\frac{(1+\delta)^{p}-1}{\delta}=p,$
@@
 $\sum_{i\in[t]}(i+1)i(\sigma_{i-1}-\sigma_{i})\sigma_{i}\leq 2p\varsigma^{2}t^{2(1-p)}.$
-Dividing both sides by $t(t+1)\sigma_{t}^{2}$, we deduce that $V\leq 2p$. When $01-2p<$ 0, we have
+Dividing both sides by $t(t+1)\sigma_{t}^{2}$, we deduce that $V\leq 2p$. When 12p<01-2p<0, we have
 $\textstyle\sum_{i\in[t]}(i+1)i(\sigma_{i-1}-\sigma_{i})\sigma_{i}\leq 2p\varsigma^{2}\left(1+\frac{1}{2-2p}\left(t^{2-2p}-1\right)\right)\leq\varsigma^{2}\frac{p}{1-p}t^{2(1-p)},$
@@
 Dividing both sides by $t(t+1)\sigma_{t}^{2}$, we deduce that $V\leq{p}/({1-p})$, thus the claim follows.
-As for assertion (ii), it is trivial to see that 3 holds.
+As for assertion (ii), it is trivial to see that ˜3 holds.
 We thus focus on bounding $\bar{C}_{\delta}$.
@@
 $\frac{8c\left(4(\sigma_{0}L_{f}+L_{g})D^{2}+3(\sigma_{0}\sigma_{f}+\sigma_{g})D\right)}{(t+1)^{1/2}\sigma_{t}}\sqrt{\log\left({8dt^{2}}/{\delta}\right)}$
-$=8c\left(4\left(L_{f}+\frac{L_{g}}{\varsigma}\right)D^{2}+3\left(\sigma_{f}+\frac{\sigma_{g}}{\varsigma}\right)D\right)\sqrt{\frac{1}{(t+1)^{1-2p}}\left(\frac{2}{1-2p}\log\left(t^{1-2p}\right)+\log\left(\frac{8d}{\delta}\right)\right)}$
-$\leq 8c\left(4\left(L_{f}+\frac{L_{g}}{\varsigma}\right)D^{2}+3\left(\sigma_{f}+\frac{\sigma_{g}}{\varsigma}\right)D\right)\sqrt{\frac{2}{(1-2p)e}+\log\left(\frac{8d}{\delta}\right)}.$
+$=8c\left(4\left(L_{f}+\frac{L_{g}}{\varsigma}\right)D^{2}\!+\!3\left(\sigma_{f}\!+\!\frac{\sigma_{g}}{\varsigma}\right)D\right)\sqrt{\frac{1}{(t+1)^{1-2p}}\left(\frac{2}{1-2p}\log\left(t^{1-2p}\right)\!+\!\log\left({8d}/{\delta}\right)\right)}$
+$\leq 8c\left(4\left(L_{f}+\frac{L_{g}}{\varsigma}\right)D^{2}+3\left(\sigma_{f}+\frac{\sigma_{g}}{\varsigma}\right)D\right)\sqrt{\frac{2}{(1-2p)e}+\log\left({8d}/{\delta}\right)}.$
 The claim then follows from the definition of $\bar{C}_{\delta}$. ∎
-Proof of Theorem 1.
-This is an immediate consequence of Lemmas 5 and 6. ∎
-A.3 Proof of Theorem 2
-The proof proceeds by applying Proposition 1 with $k=q$.
-Unlike the proof of Theorem 3, here we have two main random components in Eq.
-11 and Eq.
-12: $(q+1)q\left(\Phi_{q}(x_{q})-\Phi_{q}(x_{\operatorname{opt}})\right)$ and $\sum_{i\in[t]\setminus[q]}4iD\|\widehat{\nabla\Phi}_{i-1}-\nabla\Phi_{i-1}(x_{i-1})\|$.
-We first derive probabilistic upper bounds for these terms, then carefully set the parameters in Algorithm 2 to establish the convergence rate for both lower- and upper-level problems.
+Proof of Theorem˜1.
+The first claim on the non-asymptotic convergence guarantee follows directly from Lemmas˜7 and 8.
+As for the second claim on the asymptotic convergence, we prove a more general result: even if $\sigma_{t}$ is not set as $\varsigma(t+1)^{-p}$, the iterates of Algorithm˜1 still converge almost surely, provided the regularization sequence satisfies the conditions in Lemma˜6.
+Lemma˜6 implies that $\lim_{t\to\infty}G(x_{t})=G_{\operatorname{opt}}$ with probability at least $1-\delta$.
+As $\delta$ can be arbitrarily small, we may conclude that $\lim_{t\to\infty}G(x_{t})=G_{\operatorname{opt}}$, almost surely.
+This implies that any limit point of $\{x_{t}\}_{t\geq 0}$ is in $X_{\operatorname{opt}}$ almost surely.
+Since $F$ is convex, hence lower semi-continuous over $X$, and by definition of $F_{\operatorname{opt}}$, we have $\liminf_{t\to\infty}F(x_{t})\geq F_{\operatorname{opt}}$ almost surely.
+Besides, by combining Propositions˜1 and 2, we deduce that for any $\delta\in(0,1)$, with probability at least $1-\delta$, we have, for all $t\geq 1$,
+$F(x_{t})-F_{\operatorname{opt}}$ $\leq\frac{4c\left(4(\sigma_{0}L_{f}+L_{g})D^{2}+3(\sigma_{0}\sigma_{f}+\sigma_{g})D\right)}{(t+1)^{1/2}\sigma_{t}}\sqrt{\log\left({8dt^{2}}/{\delta}\right)}$
+$\quad+\frac{2(\sigma_{0}L_{f}+L_{g})D^{2}}{(t+1)\sigma_{t}}+\frac{\sum_{i\in[t]}(i+1)i(\sigma_{i-1}-\sigma_{i})(F_{\operatorname{opt}}-F(x_{i}))}{t(t+1)\sigma_{t}}$
+$\leq\frac{4c\left(4(\sigma_{0}L_{f}+L_{g})D^{2}+3(\sigma_{0}\sigma_{f}+\sigma_{g})D\right)}{(t+1)^{1/2}\sigma_{t}}\sqrt{\log\left({8dt^{2}}/{\delta}\right)}$
+$\quad+\frac{2(\sigma_{0}L_{f}+L_{g})D^{2}}{(t+1)\sigma_{t}}+\frac{\sum_{i\in[t]}(i+1)i(\sigma_{i-1}-\sigma_{i})\max\{F_{\operatorname{opt}}-F(x_{i}),0\}}{t(t+1)\sigma_{t}},$
+where the second inequality holds due to ˜1. We claim that all terms on the right hand side converge to 0 as $t\to\infty$. This is obvious except for the last term.
+Recall that almost surely $\liminf_{t\to\infty}F(x_{t})\geq F_{\operatorname{opt}}$.
+Therefore we have $\limsup_{t\to\infty}(F_{\operatorname{opt}}-F(x_{t}))\leq 0$ and hence $\lim_{t\to\infty}\max\{F_{\operatorname{opt}}-F(x_{t}),0\}=0$ almost surely.
+Applying the Stolz–Cesàro theorem and (34) leads to
+$\lim_{t\to\infty}\frac{\sum_{i\in[t]}(i+1)i(\sigma_{i-1}-\sigma_{i})\max\{F_{\operatorname{opt}}-F(x_{i}),0\}}{t(t+1)\sigma_{t}}=0,$
+almost surely. Then each term on the right hand side of the inequality converges to 0 as $t\to\infty$. It follows that with probability at least $1-\delta$,
+$\limsup_{t\to\infty}(F(x_{t})-F_{\operatorname{opt}})$ $\leq 0.$
+Thus, since this holds for any $\delta\in(0,1)$, it holds almost surely.
+In conclusion, we have shown that $\limsup_{t\to\infty}(F(x_{t})-F_{\operatorname{opt}})\leq 0$ and $\liminf_{t\to\infty}(F(x_{t})-F_{\operatorname{opt}})\geq 0$, which implies that $\lim_{t\to\infty}F(x_{t})=F_{\operatorname{opt}}$ almost surely.
+This completes the proof.
+∎
+B.2 Proof of Theorem˜2
+We restate Theorem˜2 with explicit upper bounds on the stationary gaps and for a more general choice of $p$ and $\omega$.
+Toward the end of the proof, we show that the optimal parameters are $p=2/7$ and $\omega=6/7$.
+Theorem 2.
+Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm˜1 with stepsizes $\alpha_{t}=(t+1)^{-\omega}$ and regularization parameters $\sigma_{t}=\varsigma(t+1)^{-p}$ for any $t\geq 0$.
+If Assumption 1 holds with the exception that $F$ may be nonconvex, then with probability at least $1-\delta$, for every $t\geq 0$, it (jointly) holds that
+$\frac{\sum_{i=0}^{t}(i+1)^{-p}\mathcal{F}(x_{i})}{\sum_{i=0}^{t}(i+1)^{-p}}$
+$\leq 4(1-p)\sup_{z\in X}|F(z)|(t+1)^{\omega-1}+2(1-p)F(x_{0})(t+1)^{p-1}+\tfrac{2(1-p)(\varsigma L_{f}+L_{g})D^{2}}{\varsigma(1-\omega)}(t+1)^{p-\omega}$
+$+\tfrac{4c(1-p)\sqrt{2}(2(\varsigma L_{f}+L_{g})D+\frac{3^{\omega}}{3^{\omega}-1}(\varsigma\sigma_{f}+\sigma_{g}))D}{\varsigma(1-\omega/2)}(t+1)^{p-\omega/2}\sqrt{\log\left({4d(t+2)^{2}}/{\delta}\right)},$
+and
+$\mathcal{G}(x_{t})$ $\leq\varsigma\sup_{z\in X}|F(z)|\left(2t^{-p}+\frac{t^{1-p-w}}{1-p}\right)+\varsigma F(x_{0})t^{-\omega}+\frac{(\varsigma L_{f}+L_{g})D^{2}}{1-\omega}t^{1-2\omega}$
+$\quad+\frac{2c\sqrt{2}(2(\varsigma L_{f}+L_{g})D+\frac{3^{\omega}}{3^{\omega}-1}(\varsigma\sigma_{f}+\sigma_{g}))D}{1-\omega/2}t^{1-3\omega/2}\sqrt{\log\left({4d(t+q)^{2}}/{\delta}\right)},$
+where $c$ is some absolute constant.
+Proof of Theorem˜2.
+Following the proof of [4, Lemma 4.1], for any $i\geq 0$, one can easily show that, with probability at least $1-\eta_{i}$, we have
+$\|\widehat{\nabla\Phi}_{i}-\nabla\Phi_{i}(x_{i})\|\leq c\sqrt{2}\left(2(\sigma_{i}L_{f}+L_{g})D+\frac{3^{\omega}}{3^{\omega}-1}(\sigma_{i}\sigma_{f}+\sigma_{g})\right)(i+1)^{-\omega/2}\sqrt{\log\left(\frac{4d}{\eta_{i}}\right)}.$
+Setting $\eta_{i}=\frac{\delta}{(i+1)(i+2)}$, applying a union bound, and using the fact that $\sigma_{i}\leq\sigma_{0}=\varsigma$ we have that with probability at least $1-\delta\sum_{i=0}^{t}\frac{1}{(i+1)(i+2)}\geq 1-\delta$, the following holds:
+$\sum_{i=0}^{t}\left(2D\|\widehat{\nabla\Phi}_{i}-\nabla\Phi_{i}(x_{i})\|+\frac{(\sigma_{i}L_{f}+L_{g})D^{2}}{2(i+1)^{\omega}}\right)$
+$\leq\sum_{i=0}^{t}\left(\frac{2\sqrt{2}c(2(\varsigma L_{f}+L_{g})D+\frac{3^{\omega}}{3^{\omega}-1}(\sigma_{i}\sigma_{f}+\sigma_{g}))D}{(i+1)^{\omega/2}}\sqrt{\log\left(\frac{4d(t+2)^{2}}{\delta}\right)}+\frac{(\varsigma L_{f}+L_{g})}{2(i+1)^{\omega}}D^{2}\right)$
+$\leq\frac{2\sqrt{2}c(2(\varsigma L_{f}+L_{g})D+\frac{3^{\omega}}{3^{\omega}-1}(\varsigma\sigma_{f}+\sigma_{g}))D}{1-\omega/2}(t+1)^{1-\omega/2}\sqrt{\log\left(\frac{4d(t+2)^{2}}{\delta}\right)}$
+$\quad+\frac{(\varsigma L_{f}+L_{g})D^{2}}{1-\omega}(t+1)^{1-\omega}.$
+Applying Lemma˜3 with $k=0$, with probability at least $1-\delta$, we thus have
+$\begin{aligned} \mathcal{G}(x_{t+1})&\!\leq\!2\varsigma\sup_{z\in X}|F(z)|(t\!+\!1)^{-p}\!+\!\varsigma F(x_{0})\!-\!\sum_{i=0}^{t}\sigma_{i}\mathcal{F}(x_{i})(t\!+\!1)^{-\omega}\!+\!\tfrac{(\varsigma L_{f}\!+\!L_{g})D^{2}}{1-\omega}(t\!+\!1)^{1\!-\!2\omega}\\ &\quad\!+\!\tfrac{2\sqrt{2}c(2(\varsigma L_{f}\!+\!L_{g})D\!+\!\frac{3^{\omega}}{3^{\omega}\!-\!1}(\varsigma\sigma_{f}\!+\!\sigma_{g}))D}{1\!-\!\omega/2}(t\!+\!1)^{1\!-\!3\omega/2}\sqrt{\log\left(\tfrac{4d(t\!+\!2)^{2}}{\delta}\right)}\end{aligned}$ (17)
+Note that the maximum in the definition of function $\mathcal{F}$ is with respect to $X_{\operatorname{opt}}$, not $X$.
+Therefore, $\mathcal{F}(x_{i})$ can take both positive and negative values since $x_{i}\in X$.
+However, we can derive the following bound
+$-\sum_{i=0}^{t}\sigma_{i}\mathcal{F}(x_{i})\leq\sup_{z\in X}|\mathcal{F}(z)|\sum_{i=0}^{t}\sigma_{i}\leq\varsigma\sup_{z\in X}|\mathcal{F}(z)|\frac{(t+1)^{1-p}}{1-p},$
+which further implies that
+$\mathcal{G}(x_{t})$ $\leq\varsigma\sup_{z\in X}|F(z)|\left(2t^{-p}+\frac{t^{1-p-w}}{1-p}\right)+\varsigma F(x_{0})t^{-\omega}+\frac{(\varsigma L_{f}+L_{g})D^{2}}{1-\omega}t^{1-2\omega}$
+$\quad+\frac{2c\sqrt{2}(2(\varsigma L_{f}+L_{g})D+\frac{3^{\omega}}{3^{\omega}-1}(\varsigma\sigma_{f}+\sigma_{g}))D}{1-\omega/2}t^{1-3\omega/2}\sqrt{\log\left(\frac{4d(t+q)^{2}}{\delta}\right)}.$ (18)
+We next focus on the upper-level problem. Since $\mathcal{G}(x_{t+1})\geq 0$, using Eq.˜17, we have
+$\sum_{i=0}^{t}\sigma_{i}\mathcal{F}(x_{i})$ $\leq 2\varsigma\sup_{z\in X}|F(z)|(t+1)^{\omega-p}+\varsigma F(x_{0})+\frac{(\varsigma L_{f}+L_{g})D^{2}}{1-\omega}(t+1)^{1-\omega}$
+$\quad+\frac{2c\sqrt{2}(2(\varsigma L_{f}+L_{g})D+\frac{3^{\omega}}{3^{\omega}-1}(\varsigma\sigma_{f}+\sigma_{g}))D}{1-\omega/2}(t+1)^{1-\omega/2}\sqrt{\log\left(\frac{4d(t+2)^{2}}{\delta}\right)}.$
+Dividing both sides by $\sum_{i=0}^{t}\sigma_{i}$, and furthermore exploiting the bounds
+$\sum_{i=0}^{t}\sigma_{i}\geq\varsigma\frac{(t+2)^{1-p}-1}{1-p}\geq\varsigma\frac{(t+1)^{1-p}}{2(1-p)},$
+we arrive at
+$\frac{\sum_{i=0}^{t}\sigma_{i}\mathcal{F}(x_{i})}{\sum_{i=0}^{t}\sigma_{i}}$ $\leq 4(1-p)\sup_{z\in X}|F(z)|(t+1)^{\omega-1}+2(1-p)F(x_{0})(t+1)^{p-1}$ (19)
+$\quad+\frac{2(1-p)(\varsigma L_{f}+L_{g})D^{2}}{\varsigma(1-\omega)}(t+1)^{p-\omega}$
+$\quad+\tfrac{4c(1-p)\sqrt{2}(2(\varsigma L_{f}+L_{g})D+\frac{3^{\omega}}{3^{\omega}-1}(\varsigma\sigma_{f}+\sigma_{g}))D}{\varsigma(1-\omega/2)}(t+1)^{p-\omega/2}\sqrt{\log\left(\frac{4d(t+2)^{2}}{\delta}\right)}.$
+Both bounds involving $\mathcal{G}$ and $\mathcal{F}$ hold with probability $\geq 1-\delta$.
+Hence, the first claim on the non-asymptotic convergence guarantee follows.
+Moreover, by optimizing over the parameters $p$ and $\omega$, we aim to ensure that the right-hand sides of both bounds converge to 0.
+To achieve this, we select $p$ and $\omega$ to minimize the slowest rate with respect to $t$:
+$\min_{p,w:0<p\leq\omega}\max\left\{-p,1-p-\omega,-\omega,1-2\omega,1-3\omega/2,\omega-1,p-1,p-\omega,p-\omega/2\right\}=-\frac{1}{7}$
+which is realized by setting $p=2/7$, $\omega=6/7$ as required. Substituting these values yields the bound presented in the main body of the paper.
+As for the second claim on asymptotic convergence, we prove a more general result: If
+$\max\left\{-p,1-p-\omega,-\omega,1-2\omega,1-3\omega/2,\omega-1,p-1,p-\omega,p-\omega/2\right\}<0,$
+then it holds that
+$\liminf_{t\to\infty}\mathcal{F}(x_{t})=\lim_{t\to\infty}\mathcal{G}(x_{t})=0.$
+Under this additional assumption and the fact that $\mathcal{G}(x_{t})\geq 0$, we deduce that $\lim_{t\to\infty}\mathcal{G}(x_{t})=0$ from taking the limit on both sides of (18).
+From the continuity of $G$, we deduce that any limit point of $\{x_{t}\}_{t\geq 0}$ is in $X_{\operatorname{opt}}$.
+the continuity of $\mathcal{F}$ from Lemma˜4 and the fact that $\mathcal{F}(x)\geq 0$ for any $x\in X_{\operatorname{opt}}$, we deduce that $\liminf_{t\to\infty}\mathcal{F}(x_{t})\geq 0$.
+Recall that
+$\liminf_{t\to\infty}\mathcal{F}(x_{t})=\lim_{t\to\infty}\inf_{k\geq t}\mathcal{F}(x_{k}),$
+and
+$\frac{\sum_{i=0}^{t}\sigma_{i}\mathcal{F}(x_{i})}{\sum_{i=0}^{t}\sigma_{i}}\geq\frac{\sum_{i=0}^{t}\sigma_{i}\inf_{k\geq i}\mathcal{F}(x_{k})}{\sum_{i=0}^{t}\sigma_{i}}.$
+Combining these observations and (19), we have
+$\frac{\sum_{i=0}^{t}\sigma_{i}\inf_{k\geq i}\mathcal{F}(x_{k})}{\sum_{i=0}^{t}\sigma_{i}}$
+$\leq 4(1-p)\sup_{z\in X}|F(z)|(t+1)^{\omega-1}+2(1-p)F(x_{0})(t+1)^{p-1}$
+$\quad+\frac{2(1-p)(\varsigma L_{f}+L_{g})D^{2}}{\varsigma(1-\omega)}(t+1)^{p-\omega}$
+$\quad+\tfrac{4c(1-p)\sqrt{2}(2(\varsigma L_{f}+L_{g})D+\frac{3^{\omega}}{3^{\omega}-1}(\varsigma\sigma_{f}+\sigma_{g}))D}{\varsigma(1-\omega/2)}(t+1)^{p-\omega/2}\sqrt{\log\left(\frac{4d(t+2)^{2}}{\delta}\right)}.$
+By the Stolz–Cesàro theorem (since $p\in(0,1)$, $\sum_{t\geq 0}\sigma_{t}=\infty$) and taking the limit on both sides of the above inequality, we obtain $\liminf_{t\to\infty}\mathcal{F}(x_{t})\leq 0$.
+Since this holds with probability at least $1-\delta$ for any $\delta\in(0,1)$, it should hold almost surely.
+Thus, we conclude the proof.
+∎
+Appendix C Proofs of Section 3
+C.1 Proof of Theorem˜3
+We restate Theorem˜3 with exact upper bounds on sub-optimality gaps.
+Theorem  3.
+Suppose the expectations defining $F$ and $G$ in Eq.˜1 are from uniform distributions over finite sets of size $[n]$.
+Let $\{z_{t}\}_{t\geq 1}$ be the iterates generated by Algorithm˜2 with $\alpha_{t}=\log(q)/q$ for $0\leq t<q$, $\alpha_{t}=2/(t+2)$ for any $t\geq q$, and $\sigma_{t}:=\varsigma(\max\{t,q\}+1)^{-p}$ for some chosen $\varsigma>0$, $p\in(0,1)$ and $S=q$.
+Under ˜1, with probability at least $1-\delta$, for any $t\geq q$, it holds that
+$F(z_{t})-F_{\text{opt}}$
+$\leq\frac{\max\left\{F(x_{0})-F_{\text{opt}},0\right\}+(G(x_{0})-G_{\text{opt}})/\varsigma+(L_{f}+L_{g}/\varsigma)D^{2}\left(\frac{1}{2}+16\sqrt{\log\left(\frac{16t^{2}}{\delta}\right)}\right)\log(q)}{(q+1)^{-1}(t+1)^{1-p}t}$
+$\quad+\frac{2(L_{f}+L_{g}/\varsigma)D^{2}(t-q)+\left(8\sqrt{2}t-4q\right)(L_{f}+L_{g}/\varsigma)D^{2}\sqrt{\log\left(\frac{16t^{2}}{\delta}\right)}}{(t+1)^{1-p}t},$
+$G(z_{t})-G_{\text{opt}}$
+$\leq\left((1+2p)\left(F_{\text{opt}}-\min_{x\in X}F(x)\right)\varsigma+2(\varsigma L_{f}+L_{g})D^{2}+\left(\max_{x\in X}F(x)-f(x_{0})\right)\varsigma\right$.
+$\quad\left.+\max_{x\in X}g(x)-g(x_{0})+\frac{2}{1-p}\left(L_{f}\varsigma+L_{g}\right)D^{2}\left(\frac{1}{2}+16\sqrt{\frac{2}{e(1-p)}+\log\left(16/\delta\right)}\right)\right$.
+$\quad\left.+8\sqrt{2}\left(L_{f}\varsigma+L_{g}\right)D^{2}\sqrt{\frac{1}{e(1-p)}+\log\left(16/\delta\right)}\right)\left(1+\frac{2p}{\min\{1,2(1-p)\}}\right)(t+1)^{-p}.$
+Moreover, with probability $1$, it holds that
+$\lim_{t\to\infty}F(x_{t})=F_{\operatorname{opt}},\quad\lim_{t\to\infty}G(x_{t})=G_{\operatorname{opt}}.$
+The proof proceeds by applying Proposition˜1 with $k=q$.
+Unlike the proof of Theorem˜1, here we have two main random components in Eq.˜10 and Eq.˜11: $(q+1)q\left(\Phi_{q}(x_{q})-\Phi_{q}(x_{\operatorname{opt}})\right)$ and $\sum_{i\in[t]\setminus[q]}4iD\|\widehat{\nabla\Phi}_{i-1}-\nabla\Phi_{i-1}(x_{i-1})\|$.
+We first derive probabilistic upper bounds for these terms, then carefully set the parameters in Algorithm˜2 to establish the convergence rate for both lower- and upper-level problems.
 In the following we use the notation
-$F(x)=\sum_{i\in[n]}f_{i}(x),\quad G(x)=\sum_{i\in[n]}g_{i}(x).$
+$F(x)=\frac{1}{n}\sum_{i\in[n]}f_{i}(x),\quad G(x)=\frac{1}{n}\sum_{i\in[n]}g_{i}(x).$
 For simplicity, we assume that $F$ and $G$ have the same size support, though our arguments are easily extended to the more general case. Also, given ${\cal S}\subseteq[n]$, we write
 $f_{{\cal S}}(x):=\frac{1}{|{\cal S}|}\sum_{i\in{\cal S}}f_{i}(x),\quad g_{{\cal S}}(x):=\frac{1}{|{\cal S}|}\sum_{i\in{\cal S}}f_{i}(x).$
-We begin with a probabilistic bound on the gradient estimator used in Algorithm 2.
+We begin with a probabilistic bound on the gradient estimator used in Algorithm˜2.
 Lemma 9.
-Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm 2 with $\alpha_{t}=\log(q)/q$ for any $0\leq t<q$ and $\alpha_{t}=2/(t+2)$ for any $t\geq q$.
-If 1 holds, then for any $t\geq 0$, given $\delta\in(0,1)$, with probability at least $1-\delta$, it (jointly) holds that
-$\left\|\widehat{\nabla F}_{t}-\nabla F(x_{t})\right\|$ $\leq 8L_{f}D\frac{\log(q)}{\sqrt{qS}}\sqrt{\log\left(\frac{8}{\delta}\right)},$ (17)
-$\left\|\widehat{\nabla G}_{t}-\nabla G(x_{t})\right\|$ $\leq 8L_{g}D\frac{\log(q)}{\sqrt{qS}}\sqrt{\log\left(\frac{8}{\delta}\right)},$
+Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm˜2 with $\alpha_{t}=\log(q)/q$ for any $0\leq t<q$ and $\alpha_{t}=2/(t+2)$ for any $t\geq q$.
+If ˜1 holds, then for any $t\geq 0$, given $\delta\in(0,1)$, with probability at least $1-\delta$, it (jointly) holds that
+$\left\|\widehat{\nabla F}_{t}-\nabla F(x_{t})\right\|$ $\leq 8L_{f}D\frac{\log(q)}{\sqrt{qS}}\sqrt{\log\left(8/\delta\right)},$ (20)
+$\left\|\widehat{\nabla G}_{t}-\nabla G(x_{t})\right\|$ $\leq 8L_{g}D\frac{\log(q)}{\sqrt{qS}}\sqrt{\log\left(8/\delta\right)},$
 provided that $0\leq t<q$, and
-$\left\|\widehat{\nabla F}_{t}-\nabla F(x_{t})\right\|$ $\leq 8L_{f}D\sqrt{\frac{(t-s_{t})}{S(t+1)(s_{t}+1)}}\sqrt{\log\left(\frac{8}{\delta}\right)}$ (18)
-$\left\|\widehat{\nabla G}_{t}-\nabla G(x_{t})\right\|$ $\leq 8L_{g}D\sqrt{\frac{(t-s_{t})}{S(t+1)(s_{t}+1)}}\sqrt{\log\left(\frac{8}{\delta}\right)},$
+$\left\|\widehat{\nabla F}_{t}-\nabla F(x_{t})\right\|$ $\leq 8L_{f}D\sqrt{\frac{(t-s_{t})}{S(t+1)(s_{t}+1)}}\sqrt{\log\left(8/\delta\right)}$ (21)
+$\left\|\widehat{\nabla G}_{t}-\nabla G(x_{t})\right\|$ $\leq 8L_{g}D\sqrt{\frac{(t-s_{t})}{S(t+1)(s_{t}+1)}}\sqrt{\log\left(8/\delta\right)},$
 where $s_{t}:=q\lfloor t/q\rfloor$, provided that $t>q$.
-Proof of Lemma 7.
+Proof of Lemma˜9.
 If $t=s_{t}$, we have $\widehat{\nabla F}_{t}=\nabla F(x_{t})$. Otherwise, let ${\cal S}_{t}\subset[n]$ be the index set of size $S$ chosen at iteration $t$, and for $i\in{\cal S}_{t}$ define
@@
 $P\left(\left\|\widehat{\nabla G}_{t}-\nabla G(x_{t})\right\|\geq\lambda\right)\leq 4\exp\left(-\frac{\lambda^{2}S(t+1)(s_{t}+1)}{64L_{g}^{2}D^{2}(t-s_{t})}\right),$
-if $t>q$. Given $\delta\in(0,1)$ and $1\leq t\leq q$, setting the right hand side $=\delta/2$ and solving for $\lambda$, we have
-$P\left(\left\|\widehat{\nabla F}_{t}-\nabla F(x_{t})\right\|\geq 8L_{f}D\frac{\log(q)}{\sqrt{qS}}\sqrt{\log\left(\frac{8}{\delta}\right)}\right)\leq\frac{\delta}{2}$
+if $t>q$. Given $\delta\in(0,1)$ and $1\leq t\leq q$, setting the right hand side $\delta/2$ and solving for $\lambda$, we have
+$P\left(\left\|\widehat{\nabla F}_{t}-\nabla F(x_{t})\right\|\geq 8L_{f}D\frac{\log(q)}{\sqrt{qS}}\sqrt{\log\left(8/\delta\right)}\right)\leq\frac{\delta}{2}$
 and
-$P\left(\left\|\widehat{\nabla G}_{t}-\nabla G(x_{t})\right\|\geq 8L_{g}D\frac{\log(q)}{\sqrt{qS}}\sqrt{\log\left(\frac{8}{\delta}\right)}\right)\leq\frac{\delta}{2}.$
+$P\left(\left\|\widehat{\nabla G}_{t}-\nabla G(x_{t})\right\|\geq 8L_{g}D\frac{\log(q)}{\sqrt{qS}}\sqrt{\log\left(8/\delta\right)}\right)\leq\frac{\delta}{2}.$
 Applying union bound, we deduce that with probability at least $1-\delta$, (20) holds. Arguing similarly for $t>q$, we also deduce that (21) holds with probability at least $1-\delta$. ∎
-To establish convergence results for Algorithm 2, we replace 3 with the following.
+To establish convergence results for Algorithm˜2, we replace ˜3 with the following.
 Condition 4.
 $\sigma_{t}=\sigma_{0}$ for $0\leq t\leq q$, $(t+1)\sigma_{t+1}>t\sigma_{t}$ for any $t\geq 0$, and $\log(t)=o(t\sigma_{t})$ as $t\to\infty$.
-Note that $\sigma_{t}$ is fixed during the initialization phase, and $\sigma_{t}^{2}$ in 3 is replaced by $\sigma_{t}^{2}$ in 4.
-Unlike the proof of Theorem 1, where we set $k=0$, the proof of Theorem 2 requires setting $k=q$.
-We therefore begin by analyzing the sequence $\{x_{t}\}_{t\geq 0}$ generated by Algorithm 2 over the first $q$ iterations.
+Note that $\sigma_{t}$ is fixed during the initialization phase, and $\sigma_{t}^{2}$ in ˜3 is replaced by $\sigma_{t}^{2}$ in ˜4.
+Unlike the proof of Theorem˜1, where we set $k=0$, the proof of Theorem˜3 requires setting $k=q$.
+We therefore begin by analyzing the sequence $\{x_{t}\}_{t\geq 0}$ generated by Algorithm˜2 over the first $q$ iterations.
 Lemma 10.
-Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm 2 with $\alpha_{t}=\log(q)/q$ for any $0\leq t<q$ and $S=q$.
-If 1 and 4 hold, we have then given $\delta\in(0,1)$, with probability at least $1-\delta\sum_{0\leq i<q}1/(i+1)(i+2)$, for any $t\geq q$, it holds that
+Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm˜2 with $\alpha_{t}=\log(q)/q$ for any $0\leq t<q$ and $S=q$.
+If ˜1 and ˜4 hold, we have then given $\delta\in(0,1)$, with probability at least $1-\delta\sum_{0\leq i<q}1/(i+1)(i+2)$, for any $t\geq q$, it holds that
 $\Phi_{q}(x_{q})-\Phi_{0}(x_{\operatorname{opt}})$ $\leq\left(1-\frac{\log(q)}{q}\right)^{q}\left(\Phi_{0}(x_{0})-\Phi_{0}(x_{\operatorname{opt}})\right)$
 $\quad+(L_{f}\sigma_{0}+L_{g})D^{2}\left(\frac{1}{2}+16\sqrt{\log\left(\frac{16t^{2}}{\delta}\right)}\right)\frac{\log(q)}{q}.$
-Proof of Lemma 8.
-By Lemma 1, for any $0\leq i<q$, we have
-[…]
-By 4, we have $\alpha_{0}=\dots=\alpha_{q-1}$ and
+Proof of Lemma˜10.
+By Lemma˜1, for any $0\leq i<q$, we have
+$\begin{aligned} \Phi_{i}(x_{i+1})-\Phi_{i}(x_{\operatorname{opt}})&\leq(1-\alpha_{i})(\Phi_{i}(x_{i})-\Phi_{i}(x_{\operatorname{opt}}))+\frac{(L_{f}\sigma_{i}+L_{g})D^{2}\alpha_{i}^{2}}{2}\\ &\quad+2D\alpha_{i}\|\nabla\Phi_{i}(x_{i})-\widehat{\nabla\Phi}_{i}\|.\end{aligned}$ (22)
+By ˜4, we have $\alpha_{0}=\dots=\alpha_{q-1}$ and
 $\sigma_{0}=\dots=\sigma_{q}\quad\implies\Phi_{0}=\dots=\Phi_{q},$
@@
 $\quad+\sum_{i=0}^{q-1}2D\alpha_{0}\|\nabla\Phi_{i}(x_{i})-\widehat{\nabla\Phi}_{i}\|(1-\alpha_{0})^{q-1-i},$
-where both inequalities are direct consequences of 4 and (19). Note that
+where both inequalities are direct consequences of ˜4 and (22). Note that
 $\sum_{i=0}^{q-1}(1-\alpha_{0})^{q-1-i}=\sum_{i=0}^{q-1}(1-\alpha_{0})^{i}\leq\sum_{i=0}^{\infty}(1-\alpha_{0})^{i}=\frac{1}{\alpha_{0}},$
 where the last equality follows from convergence of geometric series. We thus arrive at
-[…]
+$\begin{aligned} \Phi_{q}(x_{q})-\Phi_{q}(x_{\operatorname{opt}})&\leq(1-\alpha_{0})^{q}\left(\Phi_{0}(x_{0})-\Phi_{0}(x_{\operatorname{opt}})\right)+\sum_{i=0}^{q-1}\frac{(L_{f}\sigma_{0}+L_{g})D^{2}\alpha_{0}}{2}\\ &\quad+\sum_{i=0}^{q-1}2D\|\nabla\Phi_{i}(x_{i})-\widehat{\nabla\Phi}_{i}\|.\end{aligned}$ (23)
 Note that by the definition of $\Phi_{i}$ and the triangle inequality, we have
 $\|\nabla\Phi_{i}(x_{i})-\widehat{\nabla\Phi}_{i}\|\leq\sigma_{0}\|\widehat{\nabla F}_{i}-\nabla F(x_{i})\|+\|\widehat{\nabla G}_{i}-\nabla G(x_{i})\|.$
-Hence, applying Lemma 7 along with a union bound argument implies that, for any $\delta\in(0,1)$, with probability at least $1-\delta\sum_{0\leq i<q}1/(i+1)(i+2)$, we have
+Hence, applying Lemma˜9 along with a union bound argument implies that, for any $\delta\in(0,1)$, with probability at least $1-\delta\sum_{0\leq i<q}1/(i+1)(i+2)$, we have
 $\sum_{i=0}^{q-1}\|\nabla\Phi_{0}(x_{i})-\widehat{\nabla\Phi}_{i}\|\leq\sum_{i=0}^{q-1}8D(\sigma_{0}L_{f}+L_{g})\frac{\log(q)}{\sqrt{qS}}\sqrt{\log\left(\frac{8(i+1)(i+2)}{\delta}\right)}.$
-Plugging the above probabilistic bound into Eq. 20, noting that $t\geq q$ and $S=q$, and using the definition of $\alpha_{0}$ complete the proof. ∎
-We next analyze the sequence $\{x_{t}\}_{t\geq 0}$ generated by Algorithm 2 after the first $q$ iterations.
+Plugging the above probabilistic bound into Eq.˜23, noting that $t\geq q$ and $S=q$, and using the definition of $\alpha_{0}$ complete the proof. ∎
+We next analyze the sequence $\{x_{t}\}_{t\geq 0}$ generated by Algorithm˜2 after the first $q$ iterations.
 Proposition 3.
-Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm 2 with $\alpha_{t}=\log(q)/q$ for any $0\leq t<q$, $\alpha_{t}=2/(t+2)$ for any $t\geq q$ and $S=q$.
-If 1 holds and the parameters $\{\sigma_{t}\}_{t\geq 0}$ are non-increasing and positive, then given $\delta\in(0,1)$, with probability at least $1-\delta\sum_{i\in[t]\setminus[q]}1/i(i+1)$, for any $t>q$, it holds that
+Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm˜2 with $\alpha_{t}=\log(q)/q$ for any $0\leq t<q$, $\alpha_{t}=2/(t+2)$ for any $t\geq q$ and $S=q$.
+If ˜1 holds and the parameters $\{\sigma_{t}\}_{t\geq 0}$ are non-increasing and positive, then given $\delta\in(0,1)$, with probability at least $1-\delta\sum_{i\in[t]\setminus[q]}1/i(i+1)$, for any $t>q$, it holds that
 $\sum_{i\in[t]\setminus[q]}i\|\widehat{\nabla\Phi}_{i-1}-\nabla\Phi_{i-1}(x_{i-1})\|\leq\left(8\sqrt{2}t-4q\right)(\sigma_{0}L_{f}+L_{g})D\sqrt{\log\left(\frac{16t^{2}}{\delta}\right)}.$ (24)
-Proof of Proposition 3.
+Proof of Proposition˜3.
 For any $i\geq 1$, we have
@@
 $\leq(\sigma_{0}\|\widehat{\nabla F}_{i-1}-\nabla F(x_{i-1})\|+\|\widehat{\nabla G}_{i-1}-\nabla G(x_{i-1})\|).$ (25)
-Given $\delta\in(0,1)$ and $q<i\leq 2q$, by Lemma 7 and the inequality Eq. 22, with probability at least $1-\delta/i(i+1)$, we have
+Given $\delta\in(0,1)$ and $q<i\leq 2q$, by Lemma˜9 and the inequality Eq.˜25, with probability at least $1-\delta/i(i+1)$, we have
 $i\|\widehat{\nabla\Phi}_{i-1}-\nabla\Phi_{i-1}(x_{i-1})\|$ $\leq 8(\sigma_{0}L_{f}+L_{g})Di\sqrt{\frac{i-1-q}{qi(q+1)}}\sqrt{\log\left(\frac{8i(i+1)}{\delta}\right)}$
@@
 where the second inequality holds as $\frac{1}{q}\sum_{q<i\leq t}i\leq\sqrt{2}t-q/2$ for all $q<t\leq 2q$.
-Similarly, given $\delta\in(0,1)$ and $i>2q$, by Lemma 7 and the inequality Eq.
-22, with probability at least $1-\delta/i(i+1)$, we have
+Similarly, given $\delta\in(0,1)$ and $i>2q$, by Lemma˜9 and the inequality Eq.˜25, with probability at least $1-\delta/i(i+1)$, we have
 $i\|\widehat{\nabla\Phi}_{i-1}-\nabla\Phi_{i-1}(x_{i-1})\|$ $\leq\frac{8(\sigma_{0}L_{f}+L_{g})Di}{\sqrt{i(q\lfloor(i-1)/q\rfloor+1)}}\sqrt{\log\left(\frac{8i(i+1)}{\delta}\right)}$
@@
 $\leq 8\sqrt{2}(\sigma_{0}L_{f}+L_{g})D\sqrt{\log\left(\frac{8i(i+1)}{\delta}\right)},$
-where the second inequality follows since $q\lfloor(i-1)/q\rfloor+1\geq i-q$, and the third inequality follows since $i/(i-q)<2$ when $i>2q$.
+where the second inequality follows since $q\lfloor(i-1)/q\rfloor+1\geq i-q$, and the third inequality follows since i/(iq)<2i/(i-q)<2 when $i>2q$.
 Thus, if $t>2q$, with probability at least $1-\delta\sum_{2q<i\leq t}1/i(i+1)=1-\delta$, we have
@@
 $V_{q}$ $:=\sup_{t\geq q}\left\{\frac{\sum_{i\in[t]\setminus[q]}(i+1)i(\sigma_{i-1}-\sigma_{i})\sigma_{i}}{(t+1)t\sigma_{t}^{2}}\right\}.$ (26h)
-Lemma 16 in Section B.4 guarantees that these quantities remain finite for any parameter choice satisfying 1 to 2 and 4.
-In particular, these conditions hold when $\sigma_{t}=\varsigma(t+1)^{p}$ for any $p\in(0,1)$, which is what is used in Theorem 2.
-We now analyze the sequence $\{x_{t}\}_{t\geq 0}$ in Algorithm 2 for the inner-level problem.
+Lemma˜17 in Section˜D.1 guarantees that these quantities remain finite for any parameter choice satisfying 1 to 2 and 4.
+In particular, these conditions hold when $\sigma_{t}=\varsigma(t+1)^{p}$ for any $p\in(0,1)$, which is what is used in Theorem˜3.
+We now analyze the sequence $\{x_{t}\}_{t\geq 0}$ in Algorithm˜2 for the inner-level problem.
 Lemma 11.
-Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm 2 with $\alpha_{t}=\log(q)/q$ for $0\leq t<q$ $\alpha_{t}=2/(t+2)$ for any $t\geq q$, $S=q$, and let $C_{\delta,q}$ be defined as in (23g).
-If 1 and 1, 2 and 4 hold, then with probability at least $1-\delta$, it holds that
+Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm˜2 with $\alpha_{t}=\log(q)/q$ for $0\leq t<q$ $\alpha_{t}=2/(t+2)$ for any $t\geq q$, $S=q$, and let $C_{\delta,q}$ be defined as in (26g).
+If ˜1 and ˜1, 2 and 4 hold, then with probability at least $1-\delta$, it holds that
 $G(x_{t})-G_{\operatorname{opt}}\leq C_{\delta,q}\sigma_{t},$ (27)
 for any $t\geq q$.
-Proof of Lemma 9.
-Combining (11) from Proposition 1 with the probabilistic bounds in Lemma 8 and Proposition 3, and after some straightforward calculations, with probability at least $1-\delta$, we have
+Proof of Lemma˜11.
+Combining (10) from Proposition˜1 with the probabilistic bounds in Lemma˜10 and Proposition˜3, and after some straightforward calculations, with probability at least $1-\delta$, we have
 $G(x_{t})-G_{\operatorname{opt}}$
@@
 Here, the right-hand side is at most $C_{\delta,q}$, and therefore, $G(x_{t})-G_{\operatorname{opt}}\leq C_{\delta,q}\sigma_{t}$ for any $t\geq q$. ∎
-We next analyze the convergence of the sequence $\{z_{t}\}_{t\geq 0}$ in Algorithm 2 for both outer- and inner-level problems.
+We next analyze the convergence of the sequence $\{z_{t}\}_{t\geq 0}$ in Algorithm˜2 for both outer- and inner-level problems.
 Lemma 12.
-Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm 2 with $\alpha_{t}=\log(q)/q$ for $0\leq t<q$ $\alpha_{t}=2/(t+2)$ for any $t\geq q$, $S=q$, and let $C_{\delta,q}$ and $V_{q}$ defined as in Eq.
-23.
-If 1 and 1, 2 and 4 hold, then with probability at least $1-\delta$, it (jointly) holds that
+Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm˜2 with $\alpha_{t}=\log(q)/q$ for $0\leq t<q$ $\alpha_{t}=2/(t+2)$ for any $t\geq q$, $S=q$, and let $C_{\delta,q}$ and $V_{q}$ defined as in Eq.˜26.
+If ˜1 and ˜1, 2 and 4 hold, then with probability at least $1-\delta$, it (jointly) holds that
 $F(z_{t})-F_{\operatorname{opt}}$
@@
 for any $t\geq q$.
-Proof of Lemma 10.
-Recall from Algorithm 1 that we have defined
+Proof of Lemma˜12.
+Recall from Algorithm˜1 that we have defined
 $\begin{cases}S_{t+1}:=(t+2)(t+1)\sigma_{t+1}+\sum_{i\in[t+1]\backslash[q]}(i+1)i(\sigma_{i-1}-\sigma_{i})\\[8.61108pt] z_{t+1}:=\frac{1}{S_{t+1}}\left((t+2)(t+1)\sigma_{t+1}x_{t+1}+\sum_{i\in[t+1]\backslash[q]}(i+1)i(\sigma_{i-1}-\sigma_{i})x_{i}\right),\end{cases}$
@@
 $\quad+\sum_{i\in[t]}\frac{4iD\|\widehat{\nabla\Phi}_{i-1}-\nabla\Phi_{i-1}(x_{i-1})\|}{S_{t}}.$
-Using the inequality $S_{t}\geq(t+1)t\sigma_{t}$, which holds thanks to 1, and applying Lemma 8 and Proposition 2, with probability at least $1-\delta$, we have
+Using the inequality $S_{t}\geq(t+1)t\sigma_{t}$, which holds thanks to ˜1, and applying Lemma˜10 and Proposition˜2, with probability at least $1-\delta$, we have
 $F(z_{t})-F_{\operatorname{opt}}$
@@
 Lemma 13.
-Consider the sequence $\sigma_{t}:=\varsigma(t+1)^{-p}$ for $t\geq 0$ and the quantities defined in Eq.
-23.
-If $p\in(0,1)$, then the sequence $\{\sigma_{t}\}_{t\geq 0}$ satisfies 1, 2 and 4 with $L=p$.
+Consider the sequence $\sigma_{t}:=\varsigma(t+1)^{-p}$ for $t\geq 0$ and the quantities defined in Eq.˜26.
+If $p\in(0,1)$, then the sequence $\{\sigma_{t}\}_{t\geq 0}$ satisfies ˜1, 2 and 4 with $L=p$.
 Furthermore,
@@
 $\bar{C}_{\delta,q}$ $\leq\max_{x\in X}F(x)-f(x_{0})+\frac{\max_{x\in X}g(x)-g(x_{0})}{\varsigma}$
-$\quad+\frac{2}{1-p}\left(L_{f}+\frac{L_{g}}{\varsigma}\right)D^{2}\left(\frac{1}{2}+16\sqrt{\frac{2}{e(1-p)}+\log\left(\frac{16}{\delta}\right)}\right)$
-$\quad+8\sqrt{2}\left(L_{f}+\frac{L_{g}}{\varsigma}\right)D^{2}\sqrt{\frac{1}{e(1-p)}+\log\left(\frac{16}{\delta}\right)}.$
-Proof of Lemma 11.
-We observe that $C_{q},V_{q}$ can be bounded from above by $C,V$ defined in (15) under the conditions in Lemma 6.
+$\quad+\frac{2}{1-p}\left(L_{f}+\frac{L_{g}}{\varsigma}\right)D^{2}\left(\frac{1}{2}+16\sqrt{\frac{2}{e(1-p)}+\log\left(16/\delta\right)}\right)$
+$\quad+8\sqrt{2}\left(L_{f}+\frac{L_{g}}{\varsigma}\right)D^{2}\sqrt{\frac{1}{e(1-p)}+\log\left(16/\delta\right)}.$
+Proof of Lemma˜13.
+We observe that $C_{q},V_{q}$ can be bounded from above by $C,V$ defined in (15) under the conditions in Lemma˜8.
 Therefore, we focus on $\bar{C}_{\delta,q}$ for the remainder of the proof.
@@
 and
-$\frac{1}{t^{1-p}}\log\left(\frac{16t^{2}}{\delta}\right)=\frac{1}{t^{1-p}}\log\left(\frac{16}{\delta}\right)+\frac{2}{1-p}\frac{\log(t^{1-p})}{t^{1-p}}\leq\log\left(\frac{16}{\delta}\right)+\frac{2}{e(1-p)}.$
+$\frac{1}{t^{1-p}}\log\left(\frac{16t^{2}}{\delta}\right)=\frac{1}{t^{1-p}}\log\left(16/\delta\right)+\frac{2}{1-p}\frac{\log(t^{1-p})}{t^{1-p}}\leq\log\left(16/\delta\right)+\frac{2}{e(1-p)}.$
 Therefore,
 $\frac{(L_{f}\varsigma+L_{g})D^{2}\left(\frac{1}{2}+16\sqrt{\log\left(\frac{16t^{2}}{\delta}\right)}\right)\log(t)}{\varsigma t^{1-p}}$
-$\leq\frac{2}{1-p}\left(L_{f}+\frac{L_{g}}{\varsigma}\right)D^{2}\left(\frac{1}{2}+16\sqrt{\frac{2}{e(1-p)}+\log\left(\frac{16}{\delta}\right)}\right)$
+$\leq\frac{2}{1-p}\left(L_{f}+\frac{L_{g}}{\varsigma}\right)D^{2}\left(\frac{1}{2}+16\sqrt{\frac{2}{e(1-p)}+\log\left(16/\delta\right)}\right)$
 Similarly, we have
-$\frac{\left(8\sqrt{2}t-4q\right)(\sigma_{0}L_{f}+L_{g})D^{2}\sqrt{\log\left(\frac{16t^{2}}{\delta}\right)}}{(t+1)t\sigma_{t}}\leq 8\sqrt{2}\left(L_{f}+\frac{L_{g}}{\varsigma}\right)D^{2}\sqrt{\frac{1}{e(1-p)}+\log\left(\frac{16}{\delta}\right)}$
+$\frac{\left(8\sqrt{2}t-4q\right)(\sigma_{0}L_{f}+L_{g})D^{2}\sqrt{\log\left(\frac{16t^{2}}{\delta}\right)}}{(t+1)t\sigma_{t}}\leq 8\sqrt{2}\left(L_{f}+\frac{L_{g}}{\varsigma}\right)D^{2}\sqrt{\frac{1}{e(1-p)}+\log\left(16/\delta\right)}$
 This concludes the proof. ∎
-Proof of Theorem 2.
-This is an immediate consequence of Lemmas 10 and 11. ∎
-Appendix B Additional Results
-B.1 Asymptotic convergence
-We show that, even when $\sigma_{t}$ is not chosen as $\varsigma(t+1)^{-p}$ as in Theorems 1 and 2, we can still get almost sure convergence of our iterates for both Algorithms 1 and 2 (though without rates), as long as the regularization sequence satisfies the conditions in Lemma 4 or Lemma 9 respectively.
-Theorem 3.
-Let $\{x_{t}\}_{t\geq 0}$ be the iterates generated by either Algorithm 1 under the same conditions as Lemma 4. Then with probability $1$, we have
-$\lim_{t\to\infty}F(x_{t})=F_{\operatorname{opt}},\quad\lim_{t\to\infty}G(x_{t})=G_{\operatorname{opt}}.$
-Proof of Theorem 3.
-Lemma 4 implies that $\lim_{t\to\infty}G(x_{t})=G_{\operatorname{opt}}$ with probability at least $1-\delta$.
-As $\delta$ can be arbitrarily small, we may conclude that $\lim_{t\to\infty}G(x_{t})=G_{\operatorname{opt}}$, almost surely.
-This implies that any limit point of $\{x_{t}\}_{t\geq 0}$ is in $X_{\operatorname{opt}}$ almost surely.
-Since $F$ is convex, hence lower semi-continuous over $X$, and by definition of $F_{\operatorname{opt}}$, we have $\liminf_{t\to\infty}F(x_{t})\geq F_{\operatorname{opt}}$ almost surely.
-Besides, by combining Propositions 1 and 2, we deduce that for any $\delta\in(0,1)$, with probability at least $1-\delta$, we have, for all $t\geq 1$,
-$F(x_{t})-F_{\operatorname{opt}}$ $\leq\frac{4c\left(4(\sigma_{0}L_{f}+L_{g})D^{2}+3(\sigma_{0}\sigma_{f}+\sigma_{g})D\right)}{(t+1)^{1/2}\sigma_{t}}\sqrt{\log\left(\frac{8dt^{2}}{\delta}\right)}$
-$\quad+\frac{2(\sigma_{0}L_{f}+L_{g})D^{2}}{(t+1)\sigma_{t}}+\frac{\sum_{i\in[t]}(i+1)i(\sigma_{i-1}-\sigma_{i})(F_{\operatorname{opt}}-F(x_{i}))}{t(t+1)\sigma_{t}}$
-$\leq\frac{4c\left(4(\sigma_{0}L_{f}+L_{g})D^{2}+3(\sigma_{0}\sigma_{f}+\sigma_{g})D\right)}{(t+1)^{1/2}\sigma_{t}}\sqrt{\log\left(\frac{8dt^{2}}{\delta}\right)}$
-$\quad+\frac{2(\sigma_{0}L_{f}+L_{g})D^{2}}{(t+1)\sigma_{t}}+\frac{\sum_{i\in[t]}(i+1)i(\sigma_{i-1}-\sigma_{i})\max\{F_{\operatorname{opt}}-F(x_{i}),0\}}{t(t+1)\sigma_{t}},$
-where the second inequality holds due to 1. We claim that all terms on the right hand side converge to 00 as $t\to\infty$. This is obvious except for the last term.
-Recall that almost surely $\liminf_{t\to\infty}F(x_{t})\geq F_{\operatorname{opt}}$.
-Therefore we have $\limsup_{t\to\infty}(F_{\operatorname{opt}}-F(x_{t}))\leq 0$, hence $\lim_{t\to\infty}\max\{F_{\operatorname{opt}}-F(x_{t}),0\}=0$ almost surely.
-Applying the Stolz–Cesàro theorem and (33) leads to
-$\lim_{t\to\infty}\frac{\sum_{i\in[t]}(i+1)i(\sigma_{i-1}-\sigma_{i})\max\{F_{\operatorname{opt}}-F(x_{i}),0\}}{t(t+1)\sigma_{t}}=0,$
-almost surely. Then each term on the right hand side of the inequality converges to 00 as $t\to\infty$. It follows that with probability at least $1-\delta$,
-$\limsup_{t\to\infty}(F(x_{t})-F_{\operatorname{opt}})$ $\leq 0.$
-Thus, since this holds for any $\delta\in(0,1)$, it holds almost surely.
-In conclusion, we have shown that $\limsup_{t\to\infty}(F(x_{t})-F_{\operatorname{opt}})\leq 0$ and $\liminf_{t\to\infty}(F(x_{t})-F_{\operatorname{opt}})\geq 0$, which implies that $\lim_{t\to\infty}F(x_{t})=F_{\operatorname{opt}}$ almost surely.
-∎
-Theorem 4.
-Let $\{x_{t}\}_{t\geq 0}$ be the iterates generated by Algorithm 2 under the same conditions as Lemma 9. Then with probability $1$, we have
-$\lim_{t\to\infty}F(x_{t})=F_{\operatorname{opt}},\quad\lim_{t\to\infty}G(x_{t})=G_{\operatorname{opt}}.$
-Proof of Theorem 4.
-Lemma 9 implies that $\lim_{t\to\infty}G(x_{t})=G_{\operatorname{opt}}$ with probability at least $1-\delta$.
+Proof of Theorem˜3.
+The first part is an immediate consequence of Lemmas˜12 and 13.
+Thus, we devote this proof for the second part by proving that a slightly more general result that as long as 1 to 2 and ˜4 hold, the asymptotic convergence remains.
+Lemma˜11 implies that $\lim_{t\to\infty}G(x_{t})=G_{\operatorname{opt}}$ with probability at least $1-\delta$.
 As $\delta$ can be arbitrarily small, we conclude that $\lim_{t\to\infty}G(x_{t})=G_{\operatorname{opt}}$, almost surely.
@@
 Since $F$ is convex, hence lower semi-continuous over $X$, and by definition of $F_{\operatorname{opt}}$, we have $\liminf_{t\to\infty}F(x_{t})\geq F_{\operatorname{opt}}$ almost surely.
-Besides, by combining Proposition 1, Lemma 8 and Proposition 3, with probability at least $1-\delta$, we have
+Besides, by combining Proposition˜1, Lemma˜10 and Proposition˜3, with probability at least $1-\delta$, we have
 $F(x_{t})-F_{\operatorname{opt}}$
@@
 $\quad+\frac{\sum_{i\in[t]}(i+1)i(\sigma_{i-1}-\sigma_{i})\max\{F_{\operatorname{opt}}-F(x_{i}),0\}}{t(t+1)\sigma_{t}}.$
-Similar to the proof of Theorem 3, all terms on the right-hand side converge to 00 as $t\to\infty$, implying that $\limsup_{t\to\infty}(F(x_{t})-F_{\operatorname{opt}})\leq 0$.
+Similar to the proof of Theorem˜1, all terms on the right-hand side converge to 0 as $t\to\infty$, implying that $\limsup_{t\to\infty}(F(x_{t})-F_{\operatorname{opt}})\leq 0$ which holds with probability at least $1-\delta$ for any $\delta\in(0,1)$.
+Thus, $\limsup_{t\to\infty}(F(x_{t})-F_{\operatorname{opt}})\leq 0$ almost surely.
 Combined with $\liminf_{t\to\infty}(F(x_{t})-F_{\operatorname{opt}})\geq 0$, this yields $\lim_{t\to\infty}F(x_{t})=F_{\operatorname{opt}}$ almost surely, concluding the proof.
 ∎
-B.2 Nonconvex setting
-Define the auxiliary functions
-$\mathcal{F}(x):=\max_{v\in X_{\operatorname{opt}}}\nabla F(x)^{\top}(x-v)\quad\&\quad\mathcal{G}(x):=G(x)-G_{\operatorname{opt}}$
-for all $x\in X$.
-Recall that if $x\in X_{\operatorname{opt}}$ then $\mathcal{F}(x)=0$ is the necessary and sufficient optimality condition in convex case.
-Thus, $\mathcal{F}$ serves as a generalization of the stationarity condition $\nabla F(x)=0$ in the unconstrained setting.
-We start by establishing a bound, akin to Lemma 2, that holds even if $F$ is nonconvex.
-Lemma 12.
-Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm 1 or Algorithm 2 with $\{\alpha_{t},\sigma_{t}\}_{t\geq 0}$ given by $\alpha_{t}=(t+1)^{-\omega}$ and $\sigma_{t}=\varsigma(t+1)^{-p}$ for $t\geq k$, with $0<p\leq\omega$.
-If 1 holds with the exception that $F$ may be nonconvex, then for every $t\geq k$, we have
-$(t+1)^{\omega}\mathcal{G}(x_{t+1})-k^{\omega}\mathcal{G}(x_{k})$ $\leq 2\varsigma\sup_{z\in X}|F(z)|(t+1)^{\omega-p}+\sigma_{k}F(x_{k})-\sum_{i=k}^{t}\sigma_{i}\mathcal{F}(x_{i})$
-$\quad+\sum_{i=k}^{t}\left(2D\left\|\widehat{\nabla\Phi}_{i}-\nabla\Phi_{i}(x_{i})\right\|+\frac{(\sigma_{i}L_{f}+L_{g})D^{2}}{2(i+1)^{\omega}}\right).$
-Proof of Lemma 12.
-Using the definition of $\Phi_{t}=\sigma_{t}F+G$, for any $t\geq k$, we have
-$\sigma_{t}F(x_{t+1})+G(x_{t+1})$ $\leq\sigma_{t}F(x_{t})+G(x_{t})+\alpha_{t}\min_{v\in X}(\sigma_{t}\nabla F(x_{t})+\nabla G(x_{t}))^{\top}(v-x_{t})$
-$\quad+2D\alpha_{t}\|\widehat{\nabla\Phi}_{t}-\nabla\Phi_{t}(x_{t})\|+\frac{(\sigma_{t}L_{f}+L_{g})D^{2}\alpha_{t}^{2}}{2}$
-where the inequality follows directly from (9) and by minimizing over $w_{t}\in X$.
-We emphasize that (9) does not involve any convexity assumptions on $\Phi_{t}$ and relies solely on its smoothness.
-Furthermore, notice that
-$\min_{v\in X}~{}(\sigma\nabla F(x)+\nabla G(x))^{\top}(v-x)$ $\leq\min_{v\in X_{\operatorname{opt}}}~{}(\sigma\nabla F(x)+\nabla G(x))^{\top}(v-x)$
-$\leq\min_{v\in X_{\operatorname{opt}}}~{}\left\{\sigma\nabla F(x)^{\top}(v-x)+G(x)-G(v)\right\}$
-$=-\sigma\mathcal{F}(x)-\mathcal{G}(x),$
-where the last equality follows from the definition of $\mathcal{F}$ and $\mathcal{G}$. Using the above bound, we may thus conclude that
-$\sigma_{t}F(x_{t+1})+G(x_{t+1})$ $\leq\sigma_{t}F(x_{t})+G(x_{t})-\alpha_{t}(\sigma_{t}\mathcal{F}(x_{t})+\mathcal{G}(x_{t}))$
-$\quad+2D\alpha_{t}\|\widehat{\nabla\Phi}_{t}-\nabla\Phi_{t}(x_{t})\|+\frac{(\sigma_{t}L_{f}+L_{g})D^{2}\alpha_{t}^{2}}{2}.$
-A simple re-arrangement and using the definition of $\mathcal{G}$ then yields the bound
-[…]
-Using the definition of $\alpha_{t}$, we arrive at
-$(t+1)^{\omega}\mathcal{G}(x_{t+1})$ $\leq t^{\omega}\mathcal{G}(x_{t})+(t+1)^{\omega}\sigma_{t}(F(x_{t})-F(x_{t+1}))-\sigma_{t}\mathcal{F}(x_{t})$
-$\quad+2D\|\widehat{\nabla\Phi}_{t}-\nabla\Phi_{t}(x_{t})\|+\frac{(\sigma_{t}L_{f}+L_{g})D^{2}}{2(t+1)^{\omega}}.$
-Unwinding this recursion back to $i=k$, we obtain
-$(t+1)^{\omega}\mathcal{G}(x_{t+1})$ $\leq k^{\omega}\mathcal{G}(x_{k})+\sum_{i=k}^{t}\left((i+1)^{\omega}\sigma_{i}(F(x_{i})-F(x_{i+1}))\right)-\sum_{i=k}^{t}\sigma_{i}\mathcal{F}(x_{i})$
-$\quad+\sum_{i=k}^{t}\left(2D\|\widehat{\nabla\Phi}_{i}-\nabla\Phi_{i}(x_{i})\|+\frac{(\sigma_{i}L_{f}+L_{g})D^{2}}{2(i+1)^{\omega}}\right).$
-Since $p\leq\omega$, we have $(i+1)^{\omega}\sigma_{i}-i^{\omega}\sigma_{i-1}\geq 0$, therefore
-$\sum_{i=k}^{t}(i+1)^{\omega}\sigma_{i}(F(x_{i})-F(x_{i+1}))$
-$=\sum_{i=k}^{t}\Big{(}F(x_{i})((i+1)^{\omega}\sigma_{i}-i^{\omega}\sigma_{i-1})\Big{)}+\sigma_{k}F(x_{k})-(t+1)^{\omega}\sigma_{t}F(x_{t+1})$
-$\leq\sup_{z\in X}|F(z)|\sum_{i=k}^{t}\Big{(}((i+1)^{\omega}\sigma_{i}-i^{\omega}\sigma_{i-1})\Big{)}+\sigma_{k}F(x_{k})+(t+1)^{\omega}\sigma_{t}\sup_{z\in X}|F(z)|$
-$=2\varsigma\sup_{z\in X}|F(z)|(t+1)^{\omega-p}+\sigma_{k}F(x_{k}).$
-This completes the proof. ∎
-B.2.1 One-sample stochastic setting
-In this section we establish the convergence of Algorithm 1 in the nonconvex setting.
-Theorem 5.
-Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm 1 with stepsizes $\alpha_{t}=(t\!+\!1)^{-\omega}$ and regularization parameters $\sigma_{t}=\varsigma(t\!+\!1)^{-p}$ for any $t\!\geq\!0$, with $p\!=\!\frac{2}{7}$ and $\omega\!=\!\frac{6}{7}$.
-If Assumption 1 holds with the exception that $F$ may be nonconvex, then with probability at least $1-\delta$, for every $t\geq 0$, it (jointly) holds that
-$\frac{\sum_{i=0}^{t}\sigma_{i}\mathcal{F}(x_{i})}{\sum_{i=0}^{t}\sigma_{i}}\leq O\left(t^{-\frac{1}{7}}\log\left(\frac{dt^{2}}{\delta}\right)\right)\quad\&\quad\mathcal{G}(x_{t})\leq O\left(t^{-\frac{1}{7}}\log\left(\frac{dt^{2}}{\delta}\right)\right).$
-Proof of Theorem 5.
-Following the proof of [4, Lemma 4.1], for any $i\geq 0$, one can easily show that, with probability at least $1-\eta_{i}$, we have
-$\|\widehat{\nabla\Phi}_{i}-\nabla\Phi_{i}(x_{i})\|\leq c\sqrt{2}\left(2(\sigma_{i}L_{f}+L_{g})D+\frac{3^{\omega}}{3^{\omega}-1}(\sigma_{i}\sigma_{f}+\sigma_{g})\right)(i+1)^{-\omega/2}\sqrt{\log\left(\frac{4d}{\eta_{i}}\right)}.$
-Setting $\eta_{i}=\frac{\delta}{(i+1)(i+2)}$, applying a union bound, and using the fact that $\sigma_{i}\leq\sigma_{0}=\varsigma$ we have that with probability at least $1-\delta\sum_{i=0}^{t}\frac{1}{(i+1)(i+2)}\geq 1-\delta$, the following holds:
-$\sum_{i=0}^{t}\left(2D\|\widehat{\nabla\Phi}_{i}-\nabla\Phi_{i}(x_{i})\|+\frac{(\sigma_{i}L_{f}+L_{g})D^{2}}{2(i+1)^{\omega}}\right)$
-$\leq\sum_{i=0}^{t}\left(\frac{2\sqrt{2}c(2(\varsigma L_{f}+L_{g})D+\frac{3^{\omega}}{3^{\omega}-1}(\sigma_{i}\sigma_{f}+\sigma_{g}))D}{(i+1)^{\omega/2}}\sqrt{\log\left(\frac{4d(t+2)^{2}}{\delta}\right)}+\frac{(\varsigma L_{f}+L_{g})}{2(i+1)^{\omega}}D^{2}\right)$
-$\leq\frac{2\sqrt{2}c(2(\varsigma L_{f}+L_{g})D+\frac{3^{\omega}}{3^{\omega}-1}(\varsigma\sigma_{f}+\sigma_{g}))D}{1-\omega/2}(t+1)^{1-\omega/2}\sqrt{\log\left(\frac{4d(t+2)^{2}}{\delta}\right)}$
-$\quad+\frac{(\varsigma L_{f}+L_{g})D^{2}}{1-\omega}(t+1)^{1-\omega}.$
-Applying Lemma 12 with $k=0$, with probability at least $1-\delta$, we thus have
-[…]
-Note that the function $\mathcal{F}$ is defined with respect to $X_{\operatorname{opt}}$, not $X$.
-Therefore, $\mathcal{F}(x_{i})$ can take both positive and negative values since $x_{i}\in X$.
-However, we can derive the following bound
-$-\sum_{i=0}^{t}\sigma_{i}\mathcal{F}(x_{i})\leq\sup_{z\in X}|\mathcal{F}(z)|\sum_{i=0}^{t}\sigma_{i}\leq\varsigma\sup_{z\in X}|\mathcal{F}(z)|\frac{(t+1)^{1-p}}{1-p},$
-which further implies that
-$\mathcal{G}(x_{t})$ $\leq\varsigma\sup_{z\in X}|F(z)|\left(2t^{-p}+\frac{t^{1-p-w}}{1-p}\right)+\varsigma F(x_{0})t^{-\omega}+\frac{(\varsigma L_{f}+L_{g})D^{2}}{1-\omega}t^{1-2\omega}$
-$\quad+\frac{2c\sqrt{2}(2(\varsigma L_{f}+L_{g})D+\frac{3^{\omega}}{3^{\omega}-1}(\varsigma\sigma_{f}+\sigma_{g}))D}{1-\omega/2}t^{1-3\omega/2}\sqrt{\log\left(\frac{4d(t+q)^{2}}{\delta}\right)}.$
-We next focus on the upper-level problem. Since $\mathcal{G}(x_{t+1})\geq 0$, using Eq. 26, we have
-$\sum_{i=0}^{t}\sigma_{i}\mathcal{F}(x_{i})$ $\leq 2\varsigma\sup_{z\in X}|F(z)|(t+1)^{\omega-p}+\varsigma F(x_{0})+\frac{(\varsigma L_{f}+L_{g})D^{2}}{1-\omega}(t+1)^{1-\omega}$
-$+\frac{2c\sqrt{2}(2(\varsigma L_{f}+L_{g})D+\frac{3^{\omega}}{3^{\omega}-1}(\varsigma\sigma_{f}+\sigma_{g}))D}{1-\omega/2}(t+1)^{1-\omega/2}\sqrt{\log\left(\frac{4d(t+2)^{2}}{\delta}\right)}.$
-Dividing both sides by $\sum_{i=0}^{t}\sigma_{i}$, and furthermore exploiting the bounds
-$\sum_{i=0}^{t}\sigma_{i}\geq\varsigma\frac{(t+2)^{1-p}-1}{1-p}\geq\varsigma\frac{(t+1)^{1-p}}{2(1-p)},$
-we arrive at
-$\frac{\sum_{i=0}^{t}\sigma_{i}\mathcal{F}(x_{i})}{\sum_{i=0}^{t}\sigma_{i}}$ $\leq 4(1-p)\sup_{z\in X}|F(z)|(t+1)^{\omega-1}+2(1-p)F(x_{0})(t+1)^{p-1}$
-$\quad+\frac{2(1-p)(\varsigma L_{f}+L_{g})D^{2}}{\varsigma(1-\omega)}(t+1)^{p-\omega}$
-$+\tfrac{4c(1-p)\sqrt{2}(2(\varsigma L_{f}+L_{g})D+\frac{3^{\omega}}{3^{\omega}-1}(\varsigma\sigma_{f}+\sigma_{g}))D}{\varsigma(1-\omega/2)}(t+1)^{p-\omega/2}\sqrt{\log\left(\frac{4d(t+2)^{2}}{\delta}\right)}.$
-Both bounds involving $\mathcal{G}$ and $\mathcal{F}$ hold with probability $\geq 1-\delta$.
-We want the right hand side of both bounds $\to 0$.
-In order to guarantee this, we choose $p,\omega$ to minimize the slowest rate in terms of $t$:
-$\min_{p,w:0<p\leq\omega}\max\left\{-p,1-p-\omega,-\omega,1-2\omega,1-3\omega/2,\omega-1,p-1,p-\omega,p-\omega/2\right\}=-\frac{1}{7}$
-which is realized by setting $p=2/7$, $\omega=6/7$ as required. ∎
-B.2.2 Finite-sum setting
-In this section we establish the convergence of Algorithm 2 in the nonconvex setting. To that end, we introduce the sequence $\{\beta_{t}\}_{t\geq 0}$, defined as
-$\beta_{t}=\begin{cases}\displaystyle\frac{\sigma_{0}(q+1)^{\omega}\log(q+1)}{q+1}\left(1-\frac{\log(q+1)}{q+1}\right)^{q-t}&0\leq t\leq q,\\ \sigma_{t}&t>q.\end{cases}$ (27)
-Theorem 6.
-Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm 2 with stepsizes $\alpha_{t}=\log(q+1)/(q+1)$ for any $0\leq t\leq q$, $\alpha_{t}=1/(t+1)^{\omega}$ for any $t>q$ and $S=q$, and regularization parameters $\sigma_{t}=\varsigma(\max\{t,q+1\}+1)^{-p}$ with $p=\frac{1}{2}$ and $w=\frac{3}{4}$.
-If 1 holds with the exception that $F$ may be nonconvex and the sequence $\{\beta_{t}\}_{t\geq 0}$ is defined as in (27), then with probability at least $1-\delta$, for every $t\geq q$, it (jointly) holds that
-$\frac{\sum_{i=0}^{t}\beta_{i}\mathcal{F}(x_{i})}{\sum_{i=0}^{t}\beta_{i}}\leq O\left(t^{-\frac{1}{4}}\log\left(\frac{dt^{2}}{\delta}\right)\right)\quad\&\quad\mathcal{G}(x_{t})\leq O\left(t^{-\frac{1}{4}}\log\left(\frac{dt^{2}}{\delta}\right)\right).$
+C.2 Proof of Theorem˜4
+Here, we restate Theorem˜4 with exact upper bounds on stationary gaps.
+Theorem  4.
+Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm˜2 with stepsizes $\alpha_{t}=\log(q+1)/(q+1)$ for any $0\leq t\leq q$, $\alpha_{t}=1/(t+1)^{\omega}$ for any $t>q$ and $S=q$, and regularization parameters $\sigma_{t}=\varsigma(\max\{t,q+1\}+1)^{-p}$ with $p=\frac{1}{2}$ and $w=\frac{3}{4}$.
+If ˜1 holds with the exception that $F$ may be nonconvex and the sequence $\{\beta_{t}\}_{t\geq 0}$ is defined as in (4), then with probability at least $1-\delta$, for every $t\geq q$, it (jointly) holds that
+$\frac{\sum_{i=0}^{t}\beta_{i}\mathcal{F}(x_{i})}{\sum_{i=0}^{t}\beta_{i}}$ $\leq\left(\frac{G(x_{0})-G_{\text{opt}}}{(q+1)^{1-\omega}}+\left(2\varsigma(q+1)^{\omega-p}+\varsigma(q+2)^{-p}\right)\sup_{z\in X}|F(z)|\right)\left(a+bt^{1-p}\right)^{-1}$
+$\quad+\frac{(\sigma_{0}L_{f}+L_{g})D^{2}\log(q+1)}{2(q+1)^{1-w}}\left(1+16\sqrt{\log\left(\frac{8(q+2)^{2}}{\delta}\right)}\right)\left(a+bt^{1-p}\right)^{-1}$
+$\quad+\frac{8(\sigma_{0}L_{f}+L_{g})D(t-q)^{1-w}}{1-w}\sqrt{\log\left(\frac{8(t+2)^{2}}{\delta}\right)}\left(a+bt^{1-p}\right)^{-1}$
+$\quad+2\varsigma\sup_{z\in X}|F(z)|(t+1)^{\omega-p}\left(a+bt^{1-p}\right)^{-1}$
+$\quad+\frac{(\varsigma L_{f}+L_{g})D^{2}}{1-\omega}(t+1)^{1-\omega}\left(a+bt^{1-p}\right)^{-1},$
+$\mathcal{G}(x_{t})$ $\leq\left(\frac{G(x_{0})-G_{\text{opt}}}{(q+1)^{1-\omega}}+\left(2\varsigma(q+1)^{\omega-p}+\varsigma(q+2)^{-p}\right)\sup_{z\in X}|F(z)|\right)t^{-w}$
+$\quad+\frac{(\varsigma L_{f}+L_{g})D^{2}\log(q+1)}{2(q+1)^{1-w}}\left(1+16\sqrt{\log\left(\frac{8(q+2)^{2}}{\delta}\right)}\right)t^{-w}$
+$\quad+\frac{8(\varsigma L_{f}+L_{g})D^{2}(t-1-q)^{1-w}t^{-w}}{1-w}\sqrt{\log\left(\frac{8(t+1)^{2}}{\delta}\right)}+\frac{(\varsigma L_{f}+L_{g})D^{2}}{1-\omega}t^{1-2\omega}\!\!\!$
+$\quad+2\varsigma\sup_{z\in X}|F(z)|t^{-p}+\left(\varsigma(q+1)^{\omega-p}+\varsigma\frac{(t+1)^{1-p}-q^{1-p}}{1-p}\right)t^{-w}\sup_{z\in X}|\mathcal{F}(z)|,$
+where $\beta_{i}$’s are defined in (4) and $a,b$ are constants that satisfy
+$\varsigma\frac{(q+2)^{w-p}}{3}+\varsigma\frac{t^{1-p}-(q+1)^{1-p}}{1-p}=a+bt^{1-p}.$
 The proof relies on several intermediate results. We begin with the following probabilistic bound.
 Lemma 14.
-Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm 2 with $\alpha_{t}\!=\!\alpha$ for any $0\leq t\leq q$ and $\alpha_{t}=(t+1)^{-\omega}$ with $\omega>\frac{1}{2}$ for any $t>q$.
-If 1 holds with the exception that $F$ may be nonconvex, then for any $t\geq 0$, given $\delta\in(0,1)$, with probability at least $1-\delta$, it (jointly) holds that
-$\left\|\widehat{\nabla F}_{t}-\nabla F(x_{t})\right\|$ $\leq 8L_{f}D\sqrt{\frac{\alpha^{2}q}{S}}\sqrt{\log\left(\frac{8}{\delta}\right)}$ (28)
-$\left\|\widehat{\nabla G}_{t}-\nabla G(x_{t})\right\|$ $\leq 8L_{g}D\sqrt{\frac{\alpha^{2}q}{S}}\sqrt{\log\left(\frac{8}{\delta}\right)},$
+Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm˜2 with $\alpha_{t}\!=\!\alpha$ for any $0\leq t\leq q$ and $\alpha_{t}=(t+1)^{-\omega}$ with $\omega>\frac{1}{2}$ for any $t>q$.
+If ˜1 holds with the exception that $F$ may be nonconvex, then for any $t\geq 0$, given $\delta\in(0,1)$, with probability at least $1-\delta$, it (jointly) holds that
+$\left\|\widehat{\nabla F}_{t}-\nabla F(x_{t})\right\|$ $\leq 8L_{f}D\sqrt{\frac{\alpha^{2}q}{S}}\sqrt{\log\left(8/\delta\right)}$ (28)
+$\left\|\widehat{\nabla G}_{t}-\nabla G(x_{t})\right\|$ $\leq 8L_{g}D\sqrt{\frac{\alpha^{2}q}{S}}\sqrt{\log\left(8/\delta\right)},$
 provided that $0\leq t<q$, and
-$\left\|\widehat{\nabla F}_{t}-\nabla F(x_{t})\right\|$ $\leq 8L_{f}D\sqrt{\frac{(t^{1-2w}-s_{t}^{1-2w})}{S(1-2w)}}\sqrt{\log\left(\frac{8}{\delta}\right)}$ (29)
-$\left\|\widehat{\nabla G}_{t}-\nabla G(x_{t})\right\|$ $\leq 8L_{g}D\sqrt{\frac{(t^{1-2w}-s_{t}^{1-2w})}{S(1-2w)}}\sqrt{\log\left(\frac{8}{\delta}\right)},$
+$\left\|\widehat{\nabla F}_{t}-\nabla F(x_{t})\right\|$ $\leq 8L_{f}D\sqrt{\frac{(t^{1-2w}-s_{t}^{1-2w})}{S(1-2w)}}\sqrt{\log\left(8/\delta\right)}$ (29)
+$\left\|\widehat{\nabla G}_{t}-\nabla G(x_{t})\right\|$ $\leq 8L_{g}D\sqrt{\frac{(t^{1-2w}-s_{t}^{1-2w})}{S(1-2w)}}\sqrt{\log\left(8/\delta\right)},$
 where $s_{t}:=q\lfloor t/q\rfloor$, provided that $t>q$.
-Proof of Lemma 13.
-Recall from the proof of Lemma 7 that, employing the concentration inequality in [41, Theorem 3.5] yields
+Proof of Lemma˜14.
+Recall from the proof of Lemma˜9 that, employing the concentration inequality in [41, Theorem 3.5] yields
 $P\left(\left\|\widehat{\nabla F}_{t}-\nabla F(x_{t})\right\|\geq\lambda\right)$ $\leq 4\exp\left(-\frac{\lambda^{2}}{4\sum_{s=s_{t}+1}^{t}\frac{4L_{f}^{2}D^{2}}{S}\alpha_{s-1}^{2}}\right).$
@@
 $P\left(\left\|\widehat{\nabla G}_{t}-\nabla G(x_{t})\right\|\geq\lambda\right)\leq 4\exp\left(-\frac{\lambda^{2}S(1-2w)}{64L_{g}^{2}D^{2}(t^{1-2w}-s_{t}^{1-2w})}\right),$
-if $t>q$. Given $\delta\in(0,1)$ and $1\leq t\leq q$, setting the right hand side $=\delta/2$ and solving for $\lambda$ yields
-$P\left(\left\|\widehat{\nabla F}_{t}-\nabla F(x_{t})\right\|\geq 8L_{f}D\sqrt{\frac{\alpha^{2}q}{S}}\sqrt{\log\left(\frac{8}{\delta}\right)}\right)\leq\frac{\delta}{2}$
+if $t>q$. Given $\delta\in(0,1)$ and $1\leq t\leq q$, setting the right hand side $\delta/2$ and solving for $\lambda$ yields
+$P\left(\left\|\widehat{\nabla F}_{t}-\nabla F(x_{t})\right\|\geq 8L_{f}D\sqrt{\frac{\alpha^{2}q}{S}}\sqrt{\log\left(8/\delta\right)}\right)\leq\frac{\delta}{2}$
 and
-$P\left(\left\|\widehat{\nabla G}_{t}-\nabla G(x_{t})\right\|\geq 8L_{g}D\sqrt{\frac{\alpha^{2}q}{S}}\sqrt{\log\left(\frac{8}{\delta}\right)}\right)\leq\frac{\delta}{2}.$
+$P\left(\left\|\widehat{\nabla G}_{t}-\nabla G(x_{t})\right\|\geq 8L_{g}D\sqrt{\frac{\alpha^{2}q}{S}}\sqrt{\log\left(8/\delta\right)}\right)\leq\frac{\delta}{2}.$
 Applying union bound, we deduce that with probability at least $1-\delta$, (20) holds. Arguing similarly for $t>q$, we also deduce that (21) holds with probability at least $1-\delta$. ∎
-We next analyze the sequence $\{x_{t}\}_{t\geq 0}$ generated by Algorithm 2 over the first $q$ iterations.
+We next analyze the sequence $\{x_{t}\}_{t\geq 0}$ generated by Algorithm˜2 over the first $q$ iterations.
 Lemma 15.
-Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm 2 with $\alpha_{t}=\alpha$ for any $0\leq t\leq q$ and $S=q$.
-If 1 holds with the exception that $F$ may be nonconvex and 4 is satisfied, then given $\delta\in(0,1)$, with probability at least $1-\delta\sum_{0\leq i\leq q}1/(i+1)(i+2)$, for any $10\leq t\leq q+$ 1, it holds that
+Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm˜2 with $\alpha_{t}=\alpha$ for any $0\leq t\leq q$ and $S=q$.
+If ˜1 holds with the exception that $F$ may be nonconvex and ˜4 is satisfied, then given $\delta\in(0,1)$, with probability at least $1-\delta\sum_{0\leq i\leq q}1/(i+1)(i+2)$, for any $0\leq t\leq q+1$, it holds that
 $\mathcal{G}(x_{t})$ $\leq(1-\alpha)^{t}\mathcal{G}(x_{0})+2\sigma_{0}\sup_{z\in X}|F(z)|-\sigma_{0}\sum_{i=0}^{t-1}(1-\alpha)^{t-i}\alpha\mathcal{F}(x_{i})$
 $\quad+\frac{(\sigma_{0}L_{f}+L_{g})D^{2}\alpha}{2}\left(1+16\sqrt{\log\left(\frac{8(t+2)^{2}}{\delta}\right)}\right).$
-Proof of Lemma 14.
-Applying the bound (25) and after a straightforward re-arrangement, under 4, for any $0\leq i\leq t$, we obtain
+Proof of Lemma˜15.
+Applying the bound (13) and after a straightforward re-arrangement, under ˜4, for any $0\leq i\leq t$, we obtain
 $\mathcal{G}(x_{i+1})$ $\leq(1-\alpha)\mathcal{G}(x_{i})+\sigma_{0}(F(x_{i})-\alpha\mathcal{F}(x_{i})-F(x_{i+1}))$
 $\quad+2D\alpha\|\nabla\Phi_{0}(x_{i})-\widehat{\nabla\Phi}_{i}\|+\frac{(L_{f}\sigma_{0}+L_{g})D^{2}\alpha^{2}}{2}.$
-Given $\delta\in(0,1)$, we apply Lemma 13 to deduce that with probability at least $1-\delta/(i+1)(i+2)$, we have
+Given $\delta\in(0,1)$, we apply Lemma˜14 to deduce that with probability at least $1-\delta/(i+1)(i+2)$, we have
 $\mathcal{G}(x_{i+1})$ $\leq(1-\alpha)\mathcal{G}(x_{i})+\sigma_{0}(F(x_{i})-\alpha\mathcal{F}(x_{i})-F(x_{i+1}))$
@@
 conclude the proof. ∎
-We finally analyze the sequence $\{x_{t}\}_{t\geq 0}$ generated by Algorithm 2 after the first $q$ iterations.
+We finally analyze the sequence $\{x_{t}\}_{t\geq 0}$ generated by Algorithm˜2 after the first $q$ iterations.
 Proposition 4.
-Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm 2 with $\alpha_{t}=\alpha$ for any $0\leq t<q$, $\alpha_{t}=(t+1)^{-w}$ with $\omega>\frac{1}{2}$ for any $t\geq q$ and $S=q$.
-If 1 holds with the exception that $F$ may be nonconvex and the parameters $\{\sigma_{t}\}_{t\geq 0}$ are non-increasing and positive, then given $\delta\in(0,1)$, with probability at least $1-\delta\sum_{i\in[t]\setminus[q]}\frac{1}{(i+1)(i+2)}$, for any $t>q$, it holds that
+Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm˜2 with $\alpha_{t}=\alpha$ for any $0\leq t<q$, $\alpha_{t}=(t+1)^{-w}$ with $\omega>\frac{1}{2}$ for any $t\geq q$ and $S=q$.
+If ˜1 holds with the exception that $F$ may be nonconvex and the parameters $\{\sigma_{t}\}_{t\geq 0}$ are non-increasing and positive, then given $\delta\in(0,1)$, with probability at least $1-\delta\sum_{i\in[t]\setminus[q]}\frac{1}{(i+1)(i+2)}$, for any $t>q$, it holds that
 $\sum_{i\in[t]\setminus[q]}\left\|\widehat{\nabla\Phi}_{i}-\nabla\Phi_{i}(x_{i})\right\|\leq\frac{8(\sigma_{0}L_{f}+L_{g})D(t-q)^{1-w}}{1-w}\sqrt{\log\left(\frac{8(t+2)^{2}}{\delta}\right)}.$ (30)
-Proof of Proposition 4.
+Proof of Proposition˜4.
 For any $i\geq 0$, we have
@@
 $\frac{(i^{1-2w}-s_{i}^{1-2w})}{(1-2w)}=c_{i}^{-2w}(i-s_{i})\leq\frac{q}{(i-q)^{2w}}.$
-Given $\delta\in(0,1)$ and $i>q$, by Lemma 13 and the inequality Eq. 31, with probability at least $1-\delta/(i+1)(i+2)$, we have
+Given $\delta\in(0,1)$ and $i>q$, by Lemma˜14 and the inequality Eq.˜31, with probability at least $1-\delta/(i+1)(i+2)$, we have
 $\|\widehat{\nabla\Phi}_{i}-\nabla\Phi_{i}(x_{i})\|$ $\leq\frac{8(\sigma_{0}L_{f}+L_{g})D}{(i-q)^{w}}\sqrt{\log\left(\frac{8(i+1)(i+2)}{\delta}\right)}.$
@@
 where the last inequality follows from the Riemann sum approximation. ∎
-We now instantiate the regularization parameter and analyze Algorithm 2.
+We now instantiate the regularization parameter and analyze Algorithm˜2.
 Lemma 16.
-Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm 2 with $\alpha_{t}=\alpha:=\log(q+1)/(q+1)$ for any $0\leq t\leq q$, $\alpha_{t}=1/(t+1)^{\omega}$ for any $t>q$ and $S=q$.
-If 1 holds with the exception that $F$ may be nonconvex and 4 is satisfied, then given $\delta\in(0,1)$, with probability at least $1-\delta\sum_{0\leq i<t-1}1/(i+1)(i+2)$, for any $t>q+1$, it holds that
+Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm˜2 with $\alpha_{t}=\alpha:=\log(q+1)/(q+1)$ for any $0\leq t\leq q$, $\alpha_{t}=1/(t+1)^{\omega}$ for any $t>q$ and $S=q$.
+If ˜1 holds with the exception that $F$ may be nonconvex and ˜4 is satisfied, then given $\delta\in(0,1)$, with probability at least $1-\delta\sum_{0\leq i<t-1}1/(i+1)(i+2)$, for any $t>q+1$, it holds that
 $t^{\omega}\mathcal{G}(x_{t})$ $\leq\frac{\mathcal{G}(x_{0})}{(q+1)^{1-\omega}}+\left(2\varsigma(q+1)^{\omega-p}+\varsigma(q+1)^{-p}\right)\sup_{z\in X}|F(z)|$
@@
 $\quad+2\varsigma\sup_{z\in X}|F(z)|t^{\omega-p}-\sigma_{0}(q+1)^{\omega}\sum_{i=0}^{q}\alpha\mathcal{F}(x_{i})(1-\alpha)^{q-i}-\sum_{i=q+1}^{t-1}\sigma_{i}\mathcal{F}(x_{i}).$
-Proof of Lemma 15.
-By Lemma 12, for any $t\geq q+1$, we have
+Proof of Lemma˜16.
+By Lemma˜3, for any $t\geq q+1$, we have
 $(t+1)^{\omega}\mathcal{G}(x_{t+1})$ $\leq(q+1)^{\omega}\mathcal{G}(x_{q+1})+2\varsigma\sup_{z\in X}|F(z)|(t+1)^{\omega-p}+\sigma_{q+1}F(x_{q+1})$
 $\quad-\sum_{i=q+1}^{t}\sigma_{i}\mathcal{F}(x_{i})+\sum_{i=q+1}^{t}\left(2\|\widehat{\nabla\Phi}_{i}-\nabla\Phi_{i}(x_{i})\|D+\frac{(\sigma_{i}L_{f}+L_{g})D^{2}}{2(i+1)^{\omega}}\right).$
-Applying Lemma 14, with probability at least $1-\delta\sum_{0\leq i\leq q}1/(i+1)(i+2)$, we have
+Applying Lemma˜15, with probability at least $1-\delta\sum_{0\leq i\leq q}1/(i+1)(i+2)$, we have
 $\mathcal{G}(x_{q+1})$ $\leq\frac{1}{q+1}\mathcal{G}(x_{0})+2\sigma_{0}\sup_{z\in X}|F(z)|-\sigma_{0}\sum_{i=0}^{q}\alpha\mathcal{F}(x_{i})(1-\alpha)^{q-i}$
 $\quad+\frac{(\sigma_{0}L_{f}+L_{g})D^{2}\log(q+1)}{2(q+1)}\left(1+16\sqrt{\log\left(\frac{8(q+2)^{2}}{\delta}\right)}\right).$
-Combining these two bounds with Proposition 4 then yields
+Combining these two bounds with Proposition˜4 then yields
 $(t+1)^{\omega}\mathcal{G}(x_{t+1})$ $\leq\frac{1}{(q+1)^{1-\omega}}\mathcal{G}(x_{0})+2\varsigma(q+1)^{\omega-p}\sup_{z\in X}|F(z)|$
@@
 The proof concludes using the simple inequality $\sigma_{q+1}F(x_{q+1})\leq\varsigma(q+2)^{-p}\sup_{z\in X}|F(z)|$. ∎
-We are now well-equipped to prove Theorem 6.
-Proof of Theorem 6.
-We first focus on the terms involving $\mathcal{F}(x_{i})$ in Lemma 15.
+We are now well-equipped to prove Theorem˜4.
+Proof of Theorem˜4.
+We first focus on the terms involving $\mathcal{F}(x_{i})$ in Lemma˜16.
 Note that the function $\mathcal{F}$ is defined with respect to $X_{\operatorname{opt}}$, not $X$.
@@
 However, we can derive the following bound
-$\textstyle-\sigma_{0}(q+1)^{\omega}\sum_{i=0}^{q}\alpha\mathcal{F}(x_{i})(1-\alpha)^{q-i}-\sum_{i=q+1}^{t-1}\sigma_{i}\mathcal{F}(x_{i})$
+$-\sigma_{0}(q+1)^{\omega}\sum_{i=0}^{q}\alpha\mathcal{F}(x_{i})(1-\alpha)^{q-i}-\sum_{i=q+1}^{t-1}\sigma_{i}\mathcal{F}(x_{i})$
 $\leq$ $\sup_{z\in X}|\mathcal{F}(z)|\textstyle\left(\sigma_{0}(q+1)^{\omega}\sum_{i=0}^{q}\alpha(1-\alpha)^{q-i}+\sum_{i=q+1}^{t-1}\sigma_{i}\right)$
@@
 $\leq$ $\varsigma\sup_{z\in X}|\mathcal{F}(z)|\textstyle\left((q+1)^{\omega-p}+\frac{t^{1-p}-(q+1)^{1-p}}{1-p}\right),$
-where $\alpha=\log(q+1)/(q+1)$, and the second and third inequalities follow from the geometric series bound and the definition of $\sigma_{t}$. Using this bound together with Lemma 15, we arrive at
-$\mathcal{G}(x_{t})$ $\leq\left(\frac{\mathcal{G}(x_{0})}{(q+1)^{1-\omega}}+\left(2\varsigma(q+1)^{\omega-p}+\varsigma(q+2)^{-p}\right)\sup_{z\in X}|F(z)|\right)t^{-w}$
+where $\alpha=\log(q+1)/(q+1)$, and the second and third inequalities follow from the geometric series bound and the definition of $\sigma_{t}$. Using this bound together with Lemma˜16, we arrive at
+$\mathcal{G}(x_{t})$ $\leq\left(\frac{\mathcal{G}(x_{0})}{(q+1)^{1-\omega}}+\left(2\varsigma(q+1)^{\omega-p}+\varsigma(q+2)^{-p}\right)\sup_{z\in X}|F(z)|\right)t^{-w}$ (32)
 $\quad+\frac{(\sigma_{0}L_{f}+L_{g})D^{2}\log(q+1)}{2(q+1)^{1-w}}\left(1+16\sqrt{\log\left(\frac{8(q+2)^{2}}{\delta}\right)}\right)t^{-w}$
@@
 $\quad+2\varsigma\sup_{z\in X}|F(z)|t^{-p}+\left(\varsigma(q+1)^{\omega-p}+\varsigma\frac{(t+1)^{1-p}-q^{1-p}}{1-p}\right)t^{-w}\sup_{z\in X}|\mathcal{F}(z)|.$
-We next focus on the upper-level problem. Since $\mathcal{G}(x_{t+1})\geq 0$, using Lemma 15 and the definition of the sequence $\{\beta_{t}\}_{t\geq 0}$, we obtain
+We next focus on the upper-level problem. Since $\mathcal{G}(x_{t+1})\geq 0$, using Lemma˜16 and the definition of the sequence $\{\beta_{t}\}_{t\geq 0}$, we obtain
 $\sum_{i=0}^{t}\beta_{i}\mathcal{F}(x_{i})$ $\leq\frac{1}{(q+1)^{1-\omega}}\mathcal{G}(x_{0})+\left(2\varsigma(q+1)^{\omega-p}\sup_{z\in X}+\varsigma(q+2)^{-p}\right)\sup_{z\in X}|F(z)|$
@@
 where $a$ and $b$ are constants defined to make the equality hold. Combining these bounds, we arrive at
-$\frac{\sum_{i=0}^{t}\beta_{i}\mathcal{F}(x_{i})}{\sum_{i=0}^{t}\beta_{i}}$ $\leq\left(\frac{\mathcal{G}(x_{0})}{(q+1)^{1-\omega}}+\left(2\varsigma(q+1)^{\omega-p}+\varsigma(q+2)^{-p}\right)\sup_{z\in X}|F(z)|\right)\left(a+bt^{1-p}\right)^{-1}$
+$\frac{\sum_{i=0}^{t}\beta_{i}\mathcal{F}(x_{i})}{\sum_{i=0}^{t}\beta_{i}}$ $\leq\left(\frac{\mathcal{G}(x_{0})}{(q+1)^{1-\omega}}+\left(2\varsigma(q+1)^{\omega-p}+\varsigma(q+2)^{-p}\right)\sup_{z\in X}|F(z)|\right)\left(a+bt^{1-p}\right)^{-1}$ (33)
 $\quad+\frac{(\sigma_{0}L_{f}+L_{g})D^{2}\log(q+1)}{2(q+1)^{1-w}}\left(1+16\sqrt{\log\left(\frac{8(q+2)^{2}}{\delta}\right)}\right)\left(a+bt^{1-p}\right)^{-1}$
@@
 $\quad+\frac{(\varsigma L_{f}+L_{g})D^{2}}{1-\omega}(t+1)^{1-\omega}\left(a+bt^{1-p}\right)^{-1}.$
-Both bounds involving $\mathcal{G}$ and $\mathcal{F}$ hold with probability $\geq 1-\delta$.
+Both bounds involving $\mathcal{G}$ and $\mathcal{F}$ hold with probability at least $1-\delta$.
 We want the right hand side of both bounds $\to 0$.
 In order to guarantee this, we choose $p,\omega$ to minimize the slowest rate in terms of $t$:
-$\min_{p,w:0<p\leq\omega}\max\left\{t^{p-1},t^{w-1},t^{p-w},t^{-w},t^{-p},t^{1-2w},t^{1-w-p}\right\}=-\frac{1}{4},$
-which is realized by setting $p=1/2$, $\omega=3/4$ as required. ∎
-B.3 Additional numerical experiments
-We present an additional numerical for a dictionary learning problem to illustrate the performance of Algorithms 1 and 2 in the nonconvex setting.
-We highlight that, similar to Theorems 3 and 4, one can show from Theorems 5 and 6 that there exists a subsequence of the iterates $\{x_{t}\}_{t\geq 0}$ generated by Algorithms 1 and 2 that converges almost surely to a stationary point of Eq.
-1.
-This follows from the fact that the auxiliary function $\mathcal{F}$ can be shown to be Lipschitz continuous.
-Moreover, using the same argument as in the proof of [4, Theorem 4.5 and 4.7], it is possible to show that there exists a point $x_{t^{\star}}$ in a subsequence of $\{x_{t}\}_{t\geq 0}$ that satisfies a stationarity guarantee similar to that in Theorems 5 and 6.
-Since this is merely an existence result and the proof is long, we omit the details for brevity.
-Nonetheless, we present the convergence behavior of the last iterate of the sequence $\{x_{t}\}_{t\geq 0}$ in a dictionary learning problem, which is also studied in [4, 28].
-In the dictionary learning problem, the goal is to learn a compact representation of the input data $A=\{a_{1},\dots,a_{n}\}\subseteq{\mathbf{R}}^{m}$.
-Formally, we aim to find a dictionary $D=[d_{1}\cdots d_{p}]\in{\mathbf{R}}^{m\times p}$ such that each data point $a_{i}$ can be approximated by a linear combination of the basis vectors in $D$.
-This leads to the following optimization problem
-$\min_{D,X}&\quad\frac{1}{2n}\sum_{i\in[n]}\|a_{i}-Dx_{i}\|_{2}^{2}\\ \text{s.t.}~{}&\quad D\in{\mathbb{R}}^{m\times p},\ X=[x_{1}\cdots x_{n}]\in{\mathbb{R}}^{p\times n}\\ &\quad\|d_{j}\|_{2}\leq 1,\quad\forall j\in[p]\\ &\quad\|x_{i}\|_{1}\leq\delta,\quad\forall i\in[n],\end{aligned}$
-where we refer to $X$ as the coefficient matrix.
-In practice, data points usually arrive sequentially and the representation evolves gradually.
-Hence, the dictionary must be updated sequentially as well.
-Assume that we already have learned a dictionary $\hat{D}\in{\mathbf{R}}^{m\times p}$ and the corresponding coefficient matrix $\hat{X}\in{\mathbf{R}}^{p\times n}$ for some data set $A$.
-As a new dataset $A^{\prime}=\{a_{1}^{\prime},\dots,a_{n^{\prime}}^{\prime}\}\subseteq{\mathbf{R}}^{m}$ arrives, we intend to enrich our dictionary by learning $\tilde{D}\in{\mathbf{R}}^{m\times q}$ ($q>p$) and the coefficient matrix $\tilde{X}\in{\mathbf{R}}^{q\times n^{\prime}}$ for the new dataset while maintaining good performance of $\tilde{D}$ on the old dataset $A$ as well as the learned coefficient matrix $\hat{X}$ This leads to the following finite-sum bilevel problem:
-$\min_{\tilde{D},\tilde{X}}$ $\quad F(\tilde{D},\tilde{X}):=\frac{1}{2n^{\prime}}\sum_{i\in[n^{\prime}]}\|a_{i}^{\prime}-\tilde{D}\tilde{x}_{i}\|_{2}^{2}$ (32)
-s.t. $\quad\tilde{D}\in{\mathbb{R}}^{m\times q},\ \tilde{X}=[\tilde{x}_{1}\cdots\tilde{x}_{n}]\in{\mathbb{R}}^{q\times n^{\prime}}$
-$\quad\|\tilde{x}_{i}\|_{1}\leq\delta,\quad\forall i\in[n^{\prime}],$
-$\quad\tilde{D}\in\operatorname*{arg\,min}_{\|z_{j}\|_{2}\leq 1,\forall j\in[q]}g(Z):=\frac{1}{2n}\sum_{i\in[n]}\|a_{i}-Z\hat{x}_{i}\|_{2}^{2},$
-where we denote $\hat{x}_{i}$ as the prolonged vector in ${\mathbf{R}}^{q}$ by appending zeros at the end.
+$\min_{p,w:0<p\leq\omega}\max\left\{p-1,w-1,p-w,-w,-p,1-2w,1-w-p\right\}=-\frac{1}{4},$
+which is realized by setting $p=1/2$, $\omega=3/4$ as required.
+As for the second claim on asymptotic convergence, we prove a more general result: If
+$\max\left\{p-1,w-1,p-w,-w,-p,1-2w,1-w-p\right\}<0,$
+then the asymptotic convergence holds.
+Under this additional assumption and the fact that $\mathcal{G}(x_{t})\geq 0$, we deduce that $\lim_{t\to\infty}\mathcal{G}(x_{t})=0$ from taking the limit on both sides of (32).
+Recall that
+$\liminf_{t\to\infty}\mathcal{F}(x_{t})=\lim_{t\to\infty}\inf_{k\geq t}\mathcal{F}(x_{k}),$
+and
+$\frac{\sum_{i=0}^{t}\beta_{i}\mathcal{F}(x_{i})}{\sum_{i=0}^{t}\beta_{i}}\geq\frac{\sum_{i=0}^{t}\beta_{i}\inf_{k\geq i}\mathcal{F}(x_{k})}{\sum_{i=0}^{t}\beta_{i}}.$
+Combining these observations and (33), we have
+$\frac{\sum_{i=0}^{t}\beta_{i}\inf_{k\geq i}\mathcal{F}(x_{k})}{\sum_{i=0}^{t}\beta_{i}}$
+$\leq\left(\frac{\mathcal{G}(x_{0})}{(q+1)^{1-\omega}}+\left(2\varsigma(q+1)^{\omega-p}+\varsigma(q+2)^{-p}\right)\sup_{z\in X}|F(z)|\right)\left(a+bt^{1-p}\right)^{-1}$
+$\quad+\frac{(\sigma_{0}L_{f}+L_{g})D^{2}\log(q+1)}{2(q+1)^{1-w}}\left(1+16\sqrt{\log\left(\frac{8(q+2)^{2}}{\delta}\right)}\right)\left(a+bt^{1-p}\right)^{-1}$
+$\quad+\frac{8(\sigma_{0}L_{f}+L_{g})D(t-q)^{1-w}}{1-w}\sqrt{\log\left(\frac{8(t+2)^{2}}{\delta}\right)}\left(a+bt^{1-p}\right)^{-1}$
+$\quad+2\varsigma\sup_{z\in X}|F(z)|(t+1)^{\omega-p}\left(a+bt^{1-p}\right)^{-1}+\frac{(\varsigma L_{f}+L_{g})D^{2}}{1-\omega}(t+1)^{1-\omega}\left(a+bt^{1-p}\right)^{-1}.$
+By the Stolz–Cesàro theorem (since $p\in(0,1)$, $\sum_{t\geq 0}\beta_{t}=\infty$) and taking the limit on both sides of the above inequality, we obtain $\liminf_{t\to\infty}\mathcal{F}(x_{t})\leq 0$.
+Since this holds with probability at least $1-\delta$ for any $\delta\in(0,1)$, it also holds almost surely.
+Due to the fact that any limit point of $\{x_{t}\}_{t}$ is in $X_{\operatorname{opt}}$, the continuity of $\mathcal{F}$ from Lemma˜4 and the fact that $\mathcal{F}(x)\geq 0$ for any $x\in X_{\operatorname{opt}}$, we deduce that $\liminf_{t\to\infty}\mathcal{F}(x_{t})\geq 0$.
+Thus, we conclude the proof.
+∎
+Appendix D Additional Discussions
+D.1 Finiteness of constants
+The following lemma establishes that the quantities introduced in (15) are finite.
+Lemma 17.
+If ˜1 and 1 to 3 hold, then $C,\bar{C}_{\delta},C_{\delta}$ defined in (15) are finite. Furthermore, if $L$ defined in ˜2 is strictly less than 11, then $V$ defined in (15d) is finite.
+Proof of Lemma˜17.
+We first show that under ˜1 and ˜3, when the limit in ˜2 exists, the parameter $L$ satisfies $0\leq L\leq 1$.
+While it is trivial to see $L\geq 0$ thanks to ˜1, $L\leq 1$ needs some justifications.
+For the sake of contradiction, suppose $L>1$.
+Then for some sufficiently large $t$, we have
+$t\left(\frac{\sigma_{t}}{\sigma_{t+1}}-1\right)>1\iff\frac{\sigma_{t}}{\sigma_{t+1}}>\frac{t+1}{t}\iff(t+1)\sigma_{t+1}<t\sigma_{t}.$
+This further implies
+$(t+1)\sigma_{t+1}^{2}\leq(t+1)\sigma_{t+1}\sigma_{t}\leq t\sigma_{t}^{2},$
+which always contradicts ˜3. Thus, we have $0\leq L\leq 1$ under our blanket assumptions.
+We now show that $C$ is finite. Observe that
+$\lim_{t\to\infty}\frac{(t+1)t(\sigma_{t-1}-\sigma_{t})}{(t+1)t\sigma_{t}-t(t-1)\sigma_{t-1}}=\lim_{t\to\infty}\frac{(1+\frac{2}{t-1})(t-1)(\frac{\sigma_{t-1}}{\sigma_{t}}-1)}{2-(t-1)(\frac{\sigma_{t-1}}{\sigma_{t}}-1)}=\frac{L}{2-L}$
+where the first equality is obtained by diving both the numerator and the denominator by $t\sigma_{t}$ and the second equality follows from ˜2. Moreover, we have
+$\lim_{t\to\infty}\frac{(t+1)t(\sigma_{t-1}-\sigma_{t})}{(t+1)t\sigma_{t}-t(t-1)\sigma_{t-1}}\!$ $=\!\lim_{t\to\infty}\frac{\sum_{i\in[t]}(i+1)i(\sigma_{i-1}-\sigma_{i})-\sum_{i\in[t-1]}(i+1)i(\sigma_{i-1}-\sigma_{i})}{(t+1)t\sigma_{t}-t(t-1)\sigma_{t-1}}$
+$=\lim_{t\to\infty}\frac{\sum_{i\in[t]}(i+1)i(\sigma_{i-1}-\sigma_{i})}{(t+1)t\sigma_{t}}$
+where the second equality follows from the Stolz–Cesàro theorem. This yields
+$\lim_{t\to\infty}\frac{\sum_{i\in[t]}(i+1)i(\sigma_{i-1}-\sigma_{i})}{(t+1)t\sigma_{t}}=\frac{L}{2-L}\implies\sup_{t\geq 1}\frac{\sum_{i\in[t]}(i+1)i(\sigma_{i-1}-\sigma_{i})}{(t+1)t\sigma_{t}}\in(0,\infty).$ (34)
+Using ˜2, we have $\min_{t\geq 0}(t+1)\sigma_{t}>0$. Thus, $C$ is finite. It is also straightforward to verify that $\bar{C}_{\delta}$ is finite by ˜3. Consequently, $C_{\delta}$ is finite as well.
+To conclude, we show $V$ is finite. From ˜3, $\{t\sigma_{t}\}_{t\geq 0}$ and $\{(t+1)t\sigma_{t}^{2}\}_{t\geq 0}$ are increasing and diverge to $\infty$. Note that
+$\lim_{t\to\infty}\frac{\sum_{i\in[t]}(i+1)i(\sigma_{i-1}-\sigma_{i})\sigma_{i}}{(t+1)t\sigma_{t}^{2}}$ $=\lim_{t\to\infty}\frac{(t+2)(t+1)(\sigma_{t}-\sigma_{t+1})\sigma_{t+1}}{(t+2)(t+1)\sigma_{t+1}^{2}-(t+1)t\sigma_{t}^{2}}$
+$=\lim_{t\to\infty}\frac{\left(\frac{t+2}{t}\right)t\left(\frac{\sigma_{t}}{\sigma_{t+1}}-1\right)}{\left(2-t\left(\frac{\sigma_{t}}{\sigma_{t+1}}-1\right)\right)\left(\frac{\sigma_{t}}{\sigma_{t+1}}+1\right)-2\frac{\sigma_{t}}{\sigma_{t+1}}}=\frac{L}{2(1-L)}$
+where the first equality is implied by the Stolz-Cesàro theorem, the second equality is obtained from dividing both the denominator and the numerator by $(t+1)\sigma_{t+1}^{2}$, and the third equality comes from ˜3 and
+$\lim_{t\to\infty}\frac{\sigma_{t}}{\sigma_{t+1}}=1+\lim_{t\to\infty}\frac{1}{t}\cdot t\left(\frac{\sigma_{t}}{\sigma_{t+1}}-1\right)=1.$
+Note that $L/2(1-L)\in[0,\infty)$ since $L\in[0,1)$. Thus, $V$ is finite. This completes the proof. ∎
+The following lemma establishes that the quantities introduced in (26) are finite.
+Lemma 18.
+If ˜1, ˜1, 2 and 4 hold, then $C_{q},\bar{C}_{\delta,q}$ defined in (26f), and (26g) are finite. Furthermore, if $L$ defined in ˜2 is less than 11, then $V_{q}$ defined in (26h) is finite.
+Proof of Lemma˜18.
+The proof follows a similar argument to that of Lemma˜17. Details are omitted for brevity. ∎
+D.2 Over-parameterized regression: implementation details
+The implementation details are as follows.
+For IR-SCG, presented in Algorithm˜1, we set $\sigma_{t}=\varsigma(t+1)^{-1/4}$, where $\varsigma=10$, along with $\alpha_{t}=2/(t+2)$.
+For IR-FSCG, presented in Algorithm˜2, we set $S=q=\lfloor\sqrt{n}\rfloor$, $\sigma_{t}=\varsigma(\max\{t,q\}+1)^{-1/2}$, where $\varsigma=10$, along with $\alpha_{t}=2/(t+1)$ for every $t\geq q$ and $\alpha_{t}=\log(q)/q$ for every $t<q$.
+Our linear minimization oracle involves optimization over an $\ell_{1}$-norm ball, which admits an analytical solution; see [26, Section 4.1].
+We compare our proposed methods against SBCGI with stepsize $\gamma_{t}=0.01/(t+1)$, SBCGF with constant stepsize $\gamma_{t}=10^{-5}$ (both using $K_{t}=10^{-4}/\sqrt{t+1},S=q=\lfloor\sqrt{n}\rfloor$), aR-IP-SeG with long ($\gamma_{t}=10^{-2}/(t+1)^{3/4}$) and short stepsizes ($\gamma_{t}=10^{-7}/(t+1)^{3/4}$, $\rho_{t}=10^{3}(t+1)^{1/4}$, $r=0.5$), and SDBGD with long ($\gamma_{t}=10^{-2}$) and short ($\gamma_{t}=10^{-6}$) stepsizes.
+We initialized all algorithms with randomized starting points.
+For SBCGI and SBCGF, we generated the required initial point $x^{\prime}_{0}$ by running the SPIDER-FW algorithm [57, Algorithm 2] on the inner-level problem in (6), using a stepsize of $\gamma_{t}=0.1/(t+1)$.
+The initialization process terminated either after $10^{5}$ stochastic oracle queries or when the computation time exceeded 100 seconds, with the resulting point serving as $x_{0}$ for both SBCGI and SBCGF.
+The implementation of these algorithms rely on certain linear optimization or projection oracles.
+For SBCGI and SBCGF, a linear optimization oracle over the $\ell_{1}$-norm ball intersecting with a half-space is required, which we employ CVXPY [11] to solve the problems, similar to the implementation in [4, Appendix F.1].
+To compute the projection onto the feasible set required for aR-IP-SeG, we used [13, Algorithm 1].
+To approximate the outer- and inner-level optimal values $F_{\operatorname{opt}},G_{\operatorname{opt}}$, we again employ CVXPY to solve the inner-level problem and the reformulation (2) of the bilevel problem.
+D.3 Dictionary learning: implementation details
 In our experiment, we obtain $A,A^{\prime},\hat{X}$ from the code provided by Cao et al.
-[4] with $n=n^{\prime}=250$ and $p=40,q=50$ In addition, we also choose $\delta=3$.
-The experiment is repeated $10$ times, and we report the average performance within a $1$-minute execution time limit.
-For IR-SCG, presented in Algorithm 1, we set $\sigma_{t}=\varsigma(t+1)^{-2/7}$ and $\alpha_{t}=(t+1)^{-6/7}$, where $\varsigma=0.1$.
-For IR-FSCG, presented in Algorithm 2, we set $S=q=\lfloor\sqrt{n}\rfloor$, $\sigma_{t}=\varsigma(\max\{t,q+1\}+1)^{-1/2}$, where $\varsigma=0.1$, along with $\alpha_{t}=(t+1)^{-3/4}$ for every $t\geq q+1$ and $\alpha_{t}=\log(q+1)/(q+1)$ for every $t\leq q$.
-In both experiments, the linear optimization oracle admits an analytical solution; see [25, Section 4.1].
+[4] with $n=n^{\prime}=250$ and $p=40,q=50$ .
+In addition, we also choose $\delta=3$.
+For IR-SCG, presented in Algorithm˜1, we set $\sigma_{t}=\varsigma(t+1)^{-2/7}$ and $\alpha_{t}=(t+1)^{-6/7}$, where $\varsigma=0.1$.
+For IR-FSCG, presented in Algorithm˜2, we set $S=q=\lfloor\sqrt{n}\rfloor$, $\sigma_{t}=\varsigma(\max\{t,q+1\}+1)^{-1/2}$, where $\varsigma=0.1$, along with $\alpha_{t}=(t+1)^{-3/4}$ for every $t\geq q+1$ and $\alpha_{t}=\log(q+1)/(q+1)$ for every $t\leq q$.
+In both experiments, the linear optimization oracle over the $\ell_{2}$-norm ball admits an analytical solution; see [26, Section 4.1].
 For performance comparison, we implement four algorithms: SBCGI [4, Algorithm 1] with stepsize $\gamma_{t}=0.1(t+1)^{-2/3}$, SBCGF [4, Algorithm 2] with constant stepsize $\gamma_{t}=10^{-3}$ (both using $K_{t}=0.01(t+1)^{-1/3},S=q=\lfloor\sqrt{n}\rfloor$), aR-IP-SeG [27] with long ($\gamma_{t}=10^{-2}/(t+1)^{3/4}$) and short stepsizes ($\gamma_{t}=10^{-4}/(t+1)^{3/4}$, $\rho_{t}=(t+1)^{1/4}$, $r=0.5$), and the stochastic variant of the dynamic barrier gradient descent (SDBGD) [21] with long ($\gamma_{t}=10^{-2}$) and short ($\gamma_{t}=5\times 10^{-3}$) stepsizes.
@@
 For SBCGI and SBCGF, a linear optimization oracle over the $\ell_{2}$-norm ball intersecting with a half-space is required, which we employ CVXPY [11] to solve the problems, similar to the implementation in [4, Appendix F.1].
-To compute the projection onto the feasible set required for aR-IP-SeG, we used [12, Algorithm 1] to compute projection on $\ell_{1}$ norm ball and projection onto $\ell_{2}$ norm ball admits an analytical solution.
+To compute the projection onto the feasible set required for aR-IP-SeG, we used [13, Algorithm 1] to compute projection on $\ell_{1}$ norm ball and projection onto $\ell_{2}$ norm ball admits an analytical solution.
 To approximate inner-level optimal value $G_{\operatorname{opt}}$, we again employ CVXPY to solve the inner-level problem.
-Figure 3 shows the convergence behavior of different algorithms for the inner-level (left) and outer-level (right) problems in terms of execution time.
-Our proposed stochastic methods outperform existing approaches.
-In particular, IR-FSCG achieves better solutions for the outer-level objective while maintaining strong inner-level feasibility, surpassing its mini-batch counterpart, SBCGI.
-Similarly, IR-SCG outperforms its single-sample counterpart, SBCGI.
-Finally, the table below reports the average number of iterations and stochastic oracle calls over the 4-minute time limit for each algorithm.
-Our projection-free method completes more iterations within the fixed time budget, benefiting from a simpler linear minimization oracle compared to the projection-free approach in [4].
-We also observe that SDBGD and IR-SCG perform less number of iterations compared to our methods.
-Method Average # of iterations Average # of oracle calls
-IR-SCG $32736.1$ $65472.2$
-IR-FSCG $30532.6$ $1872851.0$
-SBCGI $466.3$ $7440.6$
-SBCGF $361.84$ $6147916.0$
-aR-IP-SeG (short step) $6290.0$ $25160.0$
-aR-IP-SeG (long step) $6278.1$ $25112.4$
-SDBGD (short step) $11441.3$ $22882.6$
-SDBGD (long step) $12239.3$ $24478.6$
-Figure 3: The inner-level optimality gap (left) and the outer-level objective function (right) over time for different algorithms on the dictionary learning problem (32).
-B.4 Finiteness of constants
-The following lemma establishes that the quantities introduced in (15) are finite.
-Lemma 16.
-If 1 and 1 to 3 hold, then $C,\bar{C}_{\delta},C_{\delta}$ defined in (15) are finite. Furthermore, if $L$ defined in 2 is strictly less than $1$, then $V$ defined in (15d) is finite.
-Proof of Lemma 16.
-We first show that under 1 and 3, when the limit in 2 exists, the parameter $L$ satisfies $0\leq L\leq 1$.
-While it is trivial to see $L\geq 0$ thanks to 1, $L\leq 1$ needs some justifications.
-For the sake of contradiction, suppose $L>1$.
-Then for some sufficiently large $t$, we have
-$t\left(\frac{\sigma_{t}}{\sigma_{t+1}}-1\right)>1\iff\frac{\sigma_{t}}{\sigma_{t+1}}>\frac{t+1}{t}\iff(t+1)\sigma_{t+1}<t\sigma_{t}.$
-This further implies
-$(t+1)\sigma_{t+1}^{2}\leq(t+1)\sigma_{t+1}\sigma_{t}\leq t\sigma_{t}^{2},$
-which always contradicts 3. Thus, we have $0\leq L\leq 1$ under our blanket assumptions.
-We now show that $C$ is finite. Observe that
-$\lim_{t\to\infty}\frac{(t+1)t(\sigma_{t-1}-\sigma_{t})}{(t+1)t\sigma_{t}-t(t-1)\sigma_{t-1}}=\lim_{t\to\infty}\frac{(1+\frac{2}{t-1})(t-1)(\frac{\sigma_{t-1}}{\sigma_{t}}-1)}{2-(t-1)(\frac{\sigma_{t-1}}{\sigma_{t}}-1)}=\frac{L}{2-L}$
-where the first equality is obtained by diving both the numerator and the denominator by $t\sigma_{t}$ and the second equality follows from 2. Moreover, we have
-$\lim_{t\to\infty}\frac{(t+1)t(\sigma_{t-1}-\sigma_{t})}{(t+1)t\sigma_{t}-t(t-1)\sigma_{t-1}}\!$ $=\!\lim_{t\to\infty}\frac{\sum_{i\in[t]}(i+1)i(\sigma_{i-1}-\sigma_{i})-\sum_{i\in[t-1]}(i+1)i(\sigma_{i-1}-\sigma_{i})}{(t+1)t\sigma_{t}-t(t-1)\sigma_{t-1}}$
-$=\lim_{t\to\infty}\frac{\sum_{i\in[t]}(i+1)i(\sigma_{i-1}-\sigma_{i})}{(t+1)t\sigma_{t}}$
-where the second equality follows from the Stolz–Cesàro theorem. This yields
-$\lim_{t\to\infty}\frac{\sum_{i\in[t]}(i+1)i(\sigma_{i-1}-\sigma_{i})}{(t+1)t\sigma_{t}}=\frac{L}{2-L}\implies\sup_{t\geq 1}\frac{\sum_{i\in[t]}(i+1)i(\sigma_{i-1}-\sigma_{i})}{(t+1)t\sigma_{t}}\in(0,\infty).$ (33)
-Using 2, we have $\min_{t\geq 0}(t+1)\sigma_{t}>0$. Thus, $C$ is finite. It is also straightforward to verify that $\bar{C}_{\delta}$ is finite by 3. Consequently, $C_{\delta}$ is finite as well.
-To conclude, we show $V$ is finite. From 3, $\{t\sigma_{t}\}_{t\geq 0}$ and $\{(t+1)t\sigma_{t}^{2}\}_{t\geq 0}$ are increasing and diverge to $\infty$. Note that
-$\lim_{t\to\infty}\frac{\sum_{i\in[t]}(i+1)i(\sigma_{i-1}-\sigma_{i})\sigma_{i}}{(t+1)t\sigma_{t}^{2}}$ $=\lim_{t\to\infty}\frac{(t+2)(t+1)(\sigma_{t}-\sigma_{t+1})\sigma_{t+1}}{(t+2)(t+1)\sigma_{t+1}^{2}-(t+1)t\sigma_{t}^{2}}$
-$=\lim_{t\to\infty}\frac{\left(\frac{t+2}{t}\right)t\left(\frac{\sigma_{t}}{\sigma_{t+1}}-1\right)}{\left(2-t\left(\frac{\sigma_{t}}{\sigma_{t+1}}-1\right)\right)\left(\frac{\sigma_{t}}{\sigma_{t+1}}+1\right)-2\frac{\sigma_{t}}{\sigma_{t+1}}}=\frac{L}{2(1-L)}$
-where the first equality is implied by the Stolz-Cesàro theorem, the second equality is obtained from dividing both the denominator and the numerator by $(t+1)\sigma_{t+1}^{2}$, and the third equality comes from 3 and
-$\lim_{t\to\infty}\frac{\sigma_{t}}{\sigma_{t+1}}=1+\lim_{t\to\infty}\frac{1}{t}\cdot t\left(\frac{\sigma_{t}}{\sigma_{t+1}}-1\right)=1.$
-Note that $L/2(1-L)\in[0,\infty)$ since $L\in[0,1)$. Thus, $V$ is finite. This completes the proof. ∎
-The following lemma establishes that the quantities introduced in (23) are finite.
-Lemma 17.
-If 1, 1, 2 and 4 hold, then $C_{q},\bar{C}_{\delta,q}$ defined in (23f), and (23g) are finite. Furthermore, if $L$ defined in 2 is less than $1$, then $V_{q}$ defined in (23h) is finite.
-Proof of Lemma 17.
-The proof is similar to the proof of Lemma 16. Details are omitted for brevity. ∎